A Complete Bibliography of Publications in IEEE Transactions on Software Engineering (2010–2019)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

20 May 2021 Version 1.17

Title word cross-reference

#ifdef [MRG^{+}18].

π [PMR11]. T [HPP^{+}14].

-Calculus [PMR11]. -Wise [HPP^{+}14].

1-out-of-2 [LP13a, LP13b].

Ability [RDT^{+}10]. Abstract [GHL^{+}13]. Abstracting [MSSF13]. Abstractions [dCBGU12]. Accelerating [UR18]. Acceptable [Sha10]. Acceptance [MHS^{+}13]. Accepting [WSC^{+}15]. Access [Ano12o, BBH17, CSJ^{+}16, MMPT19, TNTN12]. Accessibility [STST12]. Accuracy [AGA13, LRCL14, PLN10].

Accurate [XXCL19, XXCL21]. across [KPJ^{+}16]. Actions [dSR11]. Active [KMK^{+}13, KMD15]. Activities [LL14, MBM^{+}17]. Activity [SMWO11]. Actually [YZL^{+}15]. Adaptable [RRL^{+}13]. Adaptation [CCG^{+}12, CKMR13, DFG16, FTG16, LLNL16, MPS12, SBWM18, NB13]. Adapting [ByBEM15, PW12]. Adaptive [ABY16, CWG^{+}18, CB17, COLY15, DR14, EEM13, LYC14, SER^{+}10, SFH^{+}17, WNP^{+}20, YDCP14]. Adequacy [SYB18]. Admit [GRH17]. Admitted [dSMST17, YXS^{+}19]. Advanced [GW19]. advertisement [Ano10b, Ano10o, Ano11j, Ano11i, Ano11q, Ano11r, Ano11-32, Ano12o, Ano12v, Ano12w, Ano12x, Ano12-32]. Affordances [MOd19]. against [HPSW16, NR19]. Agent [MLS^{+}14, PZTM13]. Aggregate [ZHMZ17]. aggregation [DJ11]. Agile [HA16, HNM13]. Agreements [SRE10]. AIG [NZ16]. Aircraft [MLS^{+}14]. Alerts [Ano11z]. Algebra [MPS12, OF13, TDGH12, TGH12]. Algorithm [Gar10, LMW^{+}17, MSY14, MA13, WZ11, XJRL12]. Algorithm-Based [Gar10]. Algorithmic [WDS10]. Algorithms [AML11, BDR^{+}20, BBA^{+}15, BBL10, HT16, POPD15, SFH^{+}17]. Aligning [AGL^{+}15]. Alleviate [BKP^{+}18, LKTC14]. Allocation [EB10, PRT10]. Alloy [OF13]. Ambient [CD12]. Ambiguity [PLM15]. Amir [Nus10e]. Amorphous [ACH^{+}13]. Analogy [KMBK12]. Analogy-Based [KMBK12]. Analyses [LGN12]. Analysis [AB12, Arc12, BSGM15, BTS17, BDP16, BKML14, BDM18, BP16a, BWRR17, BBL10, BMB^{+}15, CCP13, CGdL17, DFG16, DLR13, FTG16, GFM^{+}15, GAB^{+}10, GRH13, HCHM^{+}13, HMC19, HBC13, HYH^{+}17, HJ14, HTG19b, IRBW16, JGP13, JH11, LMST10, LS13, LHWZ17, LLC^{+}16, MTHG15, MMPT19, MPP11, MSJ^{+}17, MWMK19, MSY14, MBEN15, NRDB17, OLJM16, OT12, PG13, RM13, SBGM17, SDMW14, SBL12, SSP18, SJMH^{+}19, SYX14, TLLV20, TSBB20, TGH12, UR18, WZX^{+}13, XCZ^{+}19, dlVBWM16]. Analyzing [AEP^{+}14, BKML14, BTEGF12, CXK^{+}19, DLvL14, NWCS20]. Android [BKML14, BSGM15, BYW^{+}18, BCBR18, BLVBC^{+}15, LRCL14, OLJM16, SBGM17, WLX^{+}16, YLQ^{+}18]. Annotated [DPS12]. Annual [Ano11a, Ano12a, Ano13a]. Anomaly [BBH17]. ANSI [CFMS12]. ANSI-C [CFMS12]. Ant [CZ13]. Antipatterns [PMG19]. API [ANN^{+}19, BLVBC^{+}15, MR13, RBK^{+}13, ZNKZ18, ZM19]. API-Misuse [ANN^{+}19]. APIs [SPTRC18]. App [BSGM15, LBK^{+}15, MSJ^{+}17]. Application [ABHPW10, BPH10, BBH17, CCS^{+}13, JJW^{+}15, LZ12, MN17, OLJM16, RDT^{+}10, RLMM15]. Application-Aware [LZ12].

Application-Level [MN17]. Applications [AKD^{+}10, ADTP12, BYW^{+}18, BCE^{+}14, CMCS12, CBCB12, CSJ^{+}16, CD12, DS11, JPAB19, KT10, LZX10, LXCL14, LMW^{+}17, MSCC13, MGP^{+}12, MC10, MvDR12, MM13, MMP15, MGB12, PCPS15, PHLW18, RRL^{+}13, SER^{+}10, SRZ15, TSBB20, YWHH19, vH19, NB13]. Applying [HA16, Lat14]. Approach [AGR19, BKP^{+}18, BRLM11, CRV13, CSF^{+}16, CCP13, DLGRCBC15, EKAYW16, FAHCC14, HC18, JGP13, KKS^{+}14, LH14, LH12, MPS17, MHZ^{+}12, MSK^{+}18, MBP^{+}17, NBKC15, OY16, PSA13, PRT10, STST12, STM13, TSBB20, VBP12, WSC^{+}15, YDCP14, ZKZ10, dCMS^{+}17]. Approaches [HKB17, HTG18, KKMY10, MJC17, RBH18, HTG19a]. Appropriating [DSB14]. Approximate [BPHB19, KRG^{+}13]. Approximation [Cai18, TFB14]. Apps [BCBR18, BLVBC^{+}15, MSK^{+}18, MBCC^{+}20, SBR^{+}19, WLX^{+}16, YLQ^{+}18]. Appstore [LLL^{+}18]. Appstore-Service [LLL^{+}18]. Architectural [KGM^{+}16, KKMY10, LM11, MMPT10, MLM^{+}13, MCH16]. Architecture [ABG^{+}13, BKBR12, CXK^{+}19, DLGRCBC15, LCG^{+}18, MMMR12, PRT10, SCF^{+}12, XXCL19, XXCL21]. Architecture-Based [BKBR12, PRT10]. Architectures [BGM11, BMB^{+}15]. Archival [Nus11b]. ARENA [MBP^{+}17]. Arguments [PG13]. Arising [SRvdH12]. Array [MYK19]. Arrays [DXCW17]. Art [KZDB18]. ASCENT [WDS10]. Aspect [BH10, KGRX12, MGB12]. Aspect-Oriented [BH10, KGRX12, MGB12]. Aspectizing [TNTN12]. Aspects [BCB^{+}17, EALK^{+}15, LH14]. Assembly [CGMM16]. Assertion [Cha10]. Assess [CPR16]. Assessing [AGI^{+}13, LB10, MHS^{+}13, MBLD11, TMK15]. Assessment [HC18, LYC14, SBGM17, ZXC16]. Assessor [LP13a]. Assets [ACDO12]. Assignment [BBL10, LLC^{+}16]. associate [Nus13b]. Associations [FRGCH20]. Assumptions [KMBK12, LAL^{+}14]. Assurance [CWG^{+}18, Li12, KSA^{+}13]. Asymmetric [NR19]. Asymptotic [LYC14, SFCR16]. Asynchronous [LW13]. Atomic [WGC^{+}16]. Atomicity [LPZ12, SHCX16]. Atomicity-Violation [LPZ12]. Attack [ANC^{+}10, MW11]. Attribute [HYJY14]. Attributes [AA18]. Auction [HYJY14]. Augmentation [YHW20]. Augmented [SL12]. Author [SHB18]. Authoring [BXXL19]. authors [Ano10v, Ano10-27, Ano10y, Ano10z, Ano10x, Ano10w, Ano11-27, Ano11-31, Ano11-30, Ano11-29, Ano11-28, Ano12-31]. Autofolding [FCR^{+}17]. Automata [BKKL10, PBU16, RNP13, VDC17, WSW^{+}15, WSWQ17]. Automated [AKL13, ASBZ15, ASBZ17, BTS17, BDR^{+}20, BI15, BP16a, BWB12, CBD^{+}17, GSWH15a, GRH17, HWM18, LHS^{+}15, LRCL14, LXCL14, LWY^{+}19, MBP^{+}17, NRDB17, PZTM13, PKT18, PFN^{+}14, POM13, PC10, RBH18, RBK^{+}13, SER^{+}10, SGM13, TMHM19, XLD^{+}17, dCBGU12]. Automatic [CLP17, CK15, GMM19, JPAB19, KS12, LNFW12, LLNL16, MM16, MWMK19, MvDR12, PLM15, PBM^{+}19, RMM14, VBP12, WYZ^{+}18, XCZ^{+}19, XMD^{+}17, KDO14]. Automatically [DKM^{+}12, DLWF19, Egy11, dSMST17, ZKB16]. Automating [GITT17, YXS^{+}19]. Autonomic [BRDG16, KGDS13]. Autonomous [CDGC12]. AutoSense [NRDB17]. available [Ano11q, Ano11r]. Average [KRX11]. Aware [BG18, CGMM16, HYJY14, HBS^{+}17, Kim14, KT10, LZ12, MC10, MSK^{+}18, PZF^{+}19, SER^{+}10, SYB18, SMJ19, TdlRSCB16, WLX^{+}16, YZL^{+}15, Yil13]. Awareness [dSR11]. Away [TPB^{+}17].

Back [Ano10c, Ano10f, Ano10d, Ano10h, Ano10g, Ano10e, Ano11h, Ano11e, Ano11d, Ano11c, Ano11g, Ano11f, Ano12e, Ano12g, Ano12f, Ano12c, Ano12d, Ano12h, Ano12m, Ano11s, Ano12q, Ano12r, Ano12p, Ano12-31]. Backbone [XJRL12]. Backbone-Based [XJRL12]. Background [PWSG19]. Bad [FRA^{+}19, LMSN12, TPB^{+}17]. Bailout [BBD17]. Balancing [PMGZ13]. Based [ABHPW10, AHG13, AM13, AGR19, BSM17, BB15, BJGB10, BKP^{+}18, BDM18, BCD^{+}18, BBL^{+}11, BPR19, BKBR12, BMB^{+}15, BRLM11, CGK^{+}11, CBCB12, CAC^{+}17, CZ13, CK15, CB17, CNKM19, CCP13, CFMS12, DLvL14, DNT14, DSP11, DLGRCBC15, EEM13, FPTO11, FFM^{+}10, GRPF13, Gar10, GRH17, HM10a, HM10b, HHY^{+}14, HYJY14, HZZ^{+}17, HBS^{+}17, JGP13, JWDX17, KKS^{+}14, KMBK12, KMD15, LMST10, LW13, Li12, LLHF17, LO17, LL14, LP13a, LP13b, LCNM12, LGS13, LLNL16, LMW^{+}17, MJC17, MHL^{+}12, MC10, MBK^{+}13, MvDR12, MM13, MSY14, MB14, MBCC^{+}20, MBLD11, PZTM13, PRT10, PG13, RRS19, RCKH16, RRL^{+}13, SWAK13, SSP18, SCP14, SZ10, SCF19, TZL^{+}15, WWY^{+}17, WHYY18, WMW11, XJRL12, YZL^{+}15, YM10, NB13]. Baseline [CNKM19, KM19]. Batched [JMC13]. Bayesian [DVB13, KKMY10, MB14, WWY^{+}17]. Be [HZZ^{+}16, Nus10c]. Behavior [BGP^{+}19, DCS^{+}18, KKR10, LBK^{+}15, LYC14, MNBB19, YLQ^{+}18, ZKZ10, ZM15]. Behavioral [IRBW16, SGM13, WMW11]. Behaviors [GRH17]. Behaviour [CMM^{+}19]. Beliefs [BGP^{+}19]. Bellwethers [KM19]. Benchmark [HTG18, HTG19a, LBFW17, YBL16]. Benchmarks [CKKR12, LHS^{+}15]. Benefits [KZN14, VAJ16, WNM12]. Best [RZ10]. Better [HLP10]. Between [DLWF19, PPZ^{+}18, BVD15, CSF^{+}16, COLY15, FRGCH20, GS12]. Bias [SBH14, SHB18, TMHM16]. Biases [JG12]. Bidirectional [BDP16]. Binary [HYH^{+}17, LMW^{+}17, QSM16, XXCL19, XXCL21]. Birthmarking [TLZ^{+}18]. Birthmarks [TZL^{+}15]. Bisimulation [BSM17]. Bitstate [IJK13]. Black [SM16]. Black-Box [SM16]. Blind [ARM18]. BLISS [RGA^{+}15]. Boolean [YT18]. Both [JWDX17]. Bound [CGC19]. Bounded [CFMS12, GRPF13, RGA^{+}15]. Bounds [GRPF13, LP13a, SFCR16]. Box [SM16]. BPEL [BG11]. Branch [BDP16]. Breakdowns [CH13]. Breaking [GRPF13]. Bristlecone [DS11]. Bug [DPD^{+}18, HOLL19, KTKZ13, MHZBN15, PZF^{+}19, PTYN19, RMM14, SWAK13, TNBH13, XLD^{+}17, YBL16, ZPB^{+}10, dCMS^{+}17]. Bug-Introducing [dCMS^{+}17]. Bugs [AKD^{+}10, BLZ^{+}19, BLZ^{+}20, CMM^{+}11, GW19, HMMB18, LPZ12, NWCS20, OBPM17, WWY^{+}18, XMD^{+}17, ZWL^{+}16]. Build [KSD11]. Building [HZZ^{+}17]. BURN [CKKR12]. Burstiness [CKKR12, CMCS12]. Bursting [KGDS13]. Bursts [KRX11]. Business [GBvBD^{+}18, HJ14, MPO^{+}19]. Bypassing [HPP^{+}14].

C [CFMS12, DCMS14, LHS^{+}15]. C-O [DCMS14]. CACheck [DXCW17]. Caching [MN17]. Calculus [PMR11]. Callback [WLX^{+}16]. Callback-Aware [WLX^{+}16]. Calling [SZWZ12]. Callstack [HLP10]. Callstack-Sensitive [HLP10]. Candidate [BCD^{+}18]. Candidate-Based [BCD^{+}18]. Capabilities [LED13, SWEK15]. Capability [CDT^{+}18, COLY15]. Capsule [MM13]. Capsule-Based [MM13]. Case [ABHPW10, AA11, ACCA15, BSH^{+}18, BWRR17, CWW^{+}14, DMTR10, ET16, FCC13, GSWH15b, HZZ^{+}16, KGM^{+}16, KKK11, LMZP19, MK18, MFL^{+}16, MBK^{+}13, MLS^{+}14, MAT12, OF13, POPD15, PKT18, PBS^{+}18, SM16, Thi15, Yil13, YT18, dBFP^{+}13, CPTT12]. Case-Aware [Yil13]. Case-Based [MBK^{+}13]. Cases [AM13, CWG^{+}18, DKM^{+}12, MIA^{+}16, MHZ^{+}12, YM10]. categorieS [CPTT12]. Cause [RE13]. Causes [OBPM17]. Causing [NNL^{+}20]. CCBot [CLP17]. CEGAR [TDD14]. Cell [DXCW17, YT18]. Centered [DPD^{+}18]. Centric [KSD^{+}13]. Centroidal [STM13]. Chain [BPR19]. Chains [BDR^{+}20, BP16b]. Challenge [SZF^{+}17]. Challenges [BKML14, BSH^{+}18, KZN14, KZDB18, LBK^{+}15, RNP13]. Change [BLVBC^{+}15, BWRR17, SWAK13, YXS^{+}19, dlVBWM16]. Change- [BLVBC^{+}15]. Change-Based [SWAK13]. Change-Level [YXS^{+}19]. Changes [DR14, KNGW13, MK18, SRvdH12, dCMS^{+}17]. Channel [LR12, LP13b]. Channels [SZF^{+}17]. Characterization [JJW^{+}15]. Checkers [BSM17, KKK11]. Checking [AL10, AKRU13, ASBZ15, AKD^{+}10, BYW^{+}18, CSF^{+}16, CCS^{+}13, CFMS12, DPS12, GFM^{+}15, GBvBD^{+}18, IJK13, LW13, LAH^{+}14, LCL^{+}13, MDB13, NZ16, PHV16, SFCR16, TDD14, WSW^{+}15, WSWQ17]. Checks [BKML14]. Chief [Kra10, Med18]. CHiP [MYK19]. choicE [CPTT12]. choiceS [CPTT12]. Choosing [PBS^{+}18]. Chord [Zav17]. Churn [SMWO11]. Cina [XXC^{+}15]. Claims [BBL^{+}11]. Clarifications [LBFW17]. Class [Al 11, BKP^{+}18, BBL10, CSF^{+}16, CT13, CGC19, JMC13, JWDX17, POM13, SPG10, ZNKZ18]. Class-Imbalance [JWDX17]. Class-Level [CT13]. Class-names [ZNKZ18]. Classes [FRGCH20, RLR14]. Classification [CSVC11, FDMM18, MJC17, PBM^{+}19]. Classifier [TNBH13]. Classifiers [DVB13]. Client [MSCC13, OBPM17]. Client-Side [MSCC13, OBPM17]. Clocks [RNP13, WWY^{+}18]. Clockwork [Hat14].

Clone [NNP^{+}12]. Clones [TMK15]. Clonesat [WDZ^{+}14]. Cloud [CB17, KGDS13, NB13]. Cloud-Based [CB17]. Clustered [HOLL19]. Clustering [ASBZ17, BWB12, MA13, PHY11, WYZ^{+}18]. Co [HKB17]. Co-Evolution [HKB17]. Cochange [GS12]. Code [AA18, ANVZ14, BCB^{+}17, BW10, CCK14, DDO11, EKAYW16, FRGCH20, FCR^{+}17, KKS^{+}14, KNGW13, Kim14, LW12, LLNL16, LMW^{+}17, LCG^{+}18, MSCC13, MSS19, MM16, MGP^{+}12, MCH16, MGDL10, PBP^{+}15, PHLW18, QSM16, RLMM15, SPG15, SRvdH12, SMWO11, SWAK13, SYA^{+}13, TPB^{+}17, UR18, WDZ^{+}14, WNM12, XXCL19, XXCL21, ZKB16, ZNKZ18]. Code-Smells [KKS^{+}14]. Cognitive [SL12]. Cohesion [Al 11, YZL^{+}15]. Collaboration [BHEN13]. Collaborative [FDMM18, LF12, TS12]. Colony [CZ13]. Combination [TNBH13]. Combinatorial [HYJY14, HPP^{+}14, PCHY15, WNP^{+}20, Yil13, YDCP14]. Combined [DNT14]. Combining [MSS19, MBLD11, TZM^{+}17]. Coming [Ano12n, Ano12x]. Command [HA16]. Comments [BP16b, Her17, SSSM13, SHB18, TMHM16]. Commonality [FAHCC14]. Communicating [BKKL10, TS12]. Communication [SZF^{+}17]. Community [DBF^{+}16, DR16, ZM15]. Companies [GPU^{+}16, SGR^{+}12]. Company [PMGZ13]. Comparative [DVMB12, HTG18, HTG19a, KKK11, MOTJ14]. Comparing [BWB12, MBLD11, WNM12]. Comparison [ARM18, BJGB10, BBA^{+}15, FPTO11, KRG^{+}13, LMW^{+}17, PPZ^{+}18, Pre11, SBGM17, TMHM17, WNP^{+}20, Her17]. Comparisons [MA13]. Competition [SCF19]. Competition-Based [SCF19]. Compiler [Cha10, KPJ^{+}18]. Complete [GBvBD^{+}18, Hie15]. Completeness [RM17]. Complex [GRH13, RO13]. Complexity [LED13, RKH12, RLR14, SMWO11]. Component [BI15, BKBR12, BMB^{+}15, CSVC11, CGdL14, DFG16, KS12, Li12, PMM18, PBS^{+}18, RRL^{+}13, TLB14]. Component-Based [BMB^{+}15, Li12, RRL^{+}13]. Components [BGMM17, DBM19, LH12, SWHJ14]. Composability [ZB15]. Composite [MRD12, OLJM16]. Composition [AKL13]. Compositional [BSGM15, CPV13, LAL^{+}14, XLP^{+}16]. Comprehensible [DVB13]. Comprehension [AGI^{+}13, ARM18, BP16b, CZvD11, RDT^{+}10, SPA^{+}17, XBL^{+}18]. Comprehensive [STST12, SGS19]. Compromising [NR19]. Computational [PCPS15]. Computer [AHG13, Ano10b, Ano10o, Ano11q, Ano11r, Ano12m, Ano12n]. Computer-Based [AHG13]. Computing [CBCB12, PWSG19]. Concerns [TS12, ZJ12]. Concurrency [WWY^{+}18, YWHH19, ZWL^{+}16]. Concurrent [ABL14, BMP18, PHV16, SBL12, YHW20, ZCW15]. Conditional [XMD^{+}17]. Conducted [LLLW15]. Conference [AI12, MS10]. Confidence [KKRD12]. Configurable [MYK19, TLLV20]. Configuration [CDGC12, LKB18, NBKC15, TNBH13, XZH^{+}15]. Configurations [HPP^{+}14, SPF14]. Conflicts [BHEN13, SRvdH12]. Conformance [ASBZ15, GBvBD^{+}18, MDB13]. Congruence [KSD11]. Connecting [DBF^{+}16]. ConPredictor [YWHH19]. conquer [CPTT12]. Consequences [OBPM17]. Conservation [Hat14]. Conservative [BBL^{+}11, LP13a, LP13b]. Considering [IRBW16]. Consistency [BGMM17, CSF^{+}16, CYH^{+}11, MS18, WDZ^{+}14, WMW11]. Consistency-Maintenance [WDZ^{+}14]. Consistent [KDO14]. Constant [OLJM16, WZX^{+}13]. Constrained [MAT12]. Constraint [TSBB20]. Constraints [AIAB13, CL16, DMTR10, LWY^{+}19, NBKC15]. Construction [BSD14, LBFW17]. Constructor [MYK19]. Contemporary [BCB^{+}17]. content [KMK^{+}13]. Contents [Ano12-27, Ano12-30, Ano12-29, Ano12z, Ano12-28, Ano12y]. ConTesa [YHW20]. Context [BWRR17, KT10, MM16, SER^{+}10, SZWZ12, XXC^{+}15, YCM11]. Context-Aware [KT10, SER^{+}10]. Contexts [KPJ^{+}16]. Continuous [ABM19, MBEN15]. Contract [CLP17, dCBGU12]. Contracts [ABL14, PFN^{+}14]. Control [BRDG16, CBD^{+}17, MMPT19, SBWM18, TNTN12]. Control-Theoretical [SBWM18]. Controllable [Hie15]. Controlled [ADJ16, CZvD11, DMTR10, JG11, PR11]. Conversion [EKAYW16]. Cooperative [KKS^{+}14]. Coordination [BSH^{+}18, BVD15, CH13, KSD11, LED13]. Coping [ZB12b]. Copy [WDZ^{+}14]. Copy-and-Paste [WDZ^{+}14]. Correct [CD12, Zav17]. Correction [HTG19a]. Corrections [BLZ^{+}20]. Correctness [RHB^{+}12]. Cost [HHY^{+}14, MSS19, MS18, MA13, MHS^{+}13, WHYY18]. Cost-Effective [HHY^{+}14, WHYY18]. Costs [DLWF19]. COTS [PBS^{+}18]. Counterexamples [AL10]. Counting [FAHCC14]. Country [LBK^{+}15]. Coupling [PR11]. cover [Ano10c, Ano10f, Ano10d, Ano10h, Ano10g, Ano10e, Ano10k, Ano10m, Ano10i, Ano10l, Ano10n, Ano10j, Ano10q, Ano10r, Ano10t, Ano10p, Ano10s, Ano10u, Ano11h, Ano11e, Ano11d, Ano11c, Ano11g, Ano11f, Ano11n, Ano11m, Ano11k, Ano11p, Ano11o, Ano11l, Ano11s, Ano11v, Ano11y, Ano11x, Ano11u, Ano11w, Ano11t, Ano12e, Ano12g, Ano12f, Ano12c, Ano12d, Ano12h, Ano12k, Ano12l, Ano12q, Ano12r, Ano12p, Ano12m, Ano12u, Ano12s, Ano12t, Ano12-27, Ano12-30, Ano12-29, Ano12z, Ano12-28, Ano12y, Ano12-31]. Cover2 [Ano12i]. Cover3 [Ano12j]. Coverage [Arc12, BPR19, GSWH15b, IJK13, MSS19, SPG15, SPF14, TdlRSCB16]. Coverage-Aware [TdlRSCB16]. Coverage-Based [BPR19]. Coverage-Directed [GSWH15b]. Covering [MYK19, YT18]. COVERT [BSGM15]. Crash [CK15]. Crashes [KWK^{+}11]. Creating [MOd19]. Criteria [CT13, SBM13]. Criterion [SYB18]. Critical [BH13, BWRR17, CKMR13, DLvL14, MG20]. Criticality [BBD17, WHYY18]. Criticality-Based [WHYY18]. Critiquing [AHG13]. Cross [EKAYW16, HTG18, HTG19a, HTG19b, JWDX17, PMGZ13, XLP^{+}16, XXCL19, XXCL21]. Cross-Architecture [XXCL19, XXCL21]. Cross-Company [PMGZ13]. Cross-OS [XXCL21, XXCL19]. Cross-Platform [EKAYW16]. Cross-Project [HTG18, HTG19a, JWDX17, XLP^{+}16]. Crosschecking [CCK14]. Crosscutting [ZJ12]. Crossover [VAJ16]. Crowd [SBR^{+}19]. Crowdsourcing [SCF19]. CS [SMG11]. CS/SE [SMG11]. CSDP [Ano10o, Ano11j, Ano11i]. Culture [ABM19, SZF^{+}17]. Customer [KRX11, SCF19]. Customized [CKKR12]. Customizing [SWEK15]. Cyber [BDM18, FRA^{+}19]. Cyber-Physical [BDM18, FRA^{+}19]. Cycle [CRV11]. Cycles [WSC^{+}15].

Data [AIAB13, BHKZ13, BCE^{+}14, CSJ^{+}16, CCK14, DVMB12, GSWH15a, KZDB18, MHL^{+}12, MBK^{+}13, NRDB17, SCF^{+}12, SRS12, SSSM13, SMJ19, TWL13, ZCW15, KMK^{+}13]. Data-Aware [SMJ19]. Data-Management [BCE^{+}14]. Data-Parallel [CCK14]. Database [BBH17, MWMK19, TdlRSCB16]. Datasets [SSSM13]. Date [DSB14]. David [Nus13d]. Deadlocks [CC14, CL16, SBL12]. Dealing [CMCS12, VBJM14]. Debt [dSMST17, RK14, RK19, YXS^{+}19]. Debug [GZS^{+}15, LBB^{+}13]. Debugging [CTZ11, HLP10, KWK^{+}11, LLHF17, LL14]. DEC [KKRD12]. Decade [Nus10b]. DECAF [HYH^{+}17]. Decentralized [CGMM16, NB13]. Decision [DLvL14, MOA15, MB14, SDMW14, WGC^{+}16]. Decision-Based [DLvL14]. Decision-Making [MOA15, MB14, SDMW14]. Decisions [FRA^{+}19, MBK^{+}13]. Declarative [BBA^{+}15]. Decomposition [AGR19]. Decomposition-Based [AGR19]. DECOR [MGDL10]. Deep [CDT^{+}19]. Defect [BKP^{+}18, HTG18, HTG19a, HTG19b, JWDX17, LNH^{+}16, LJZ^{+}19, MK18, MBC^{+}13, NFK^{+}18, PMGZ13, SSSM13, SBH14, SHB18, SJS^{+}11, SGS19, SNAH15, TMHM16, TMHM17, TMHM19, WNM12, XLP^{+}16, YWHH19, ZHMZ17, Her17]. Defect-Proneness [SJS^{+}11, SNAH15]. Defects [HKS12, RM17, YGV17, dBFP^{+}13]. Defining [RT12]. Delay [MBEN15]. Delivery [CDT^{+}18]. Demand [KKRD12, LP13b]. Dense [HPSW16]. Dependability [BBL^{+}11, GPK10]. Dependence [BPH10, Cai18, WLG^{+}17]. Dependencies [KL11, LCG^{+}18]. Dependency [GS12, eZHM13, XCZ^{+}19]. Deployment [CDGC12, MMMR12]. Derivation [DLGRCBC15, Kim14]. Deriving [BSM17, WZ11]. Descartes [HBS^{+}17]. Description [YLQ^{+}18]. Description-to-Behavior [YLQ^{+}18]. Design [ACCA15, CXK^{+}19, CMSSS13, DFG16, Egy11, HMC19, JGP13, KRG^{+}13, KPJ^{+}16, LKR14, MSY14, MGDL10, MHZBN15, NDMX16, PLN10, RE13, SPG10, SZ10, ZB12a, ZB15]. Designers [MLPvdH15]. Designing [BRDG16, WDS10]. Designs [VAJ16]. DESSERT [CPTT12]. Detect [dSMST17]. Detecting [BLZ^{+}19, BLZ^{+}20, DXCW17, Egy11, KPJ^{+}18, MCH16, MGS^{+}18, PBP^{+}15, SYX14, WSC^{+}15, ZWL^{+}16]. Detection [ABL14, AB12, BH10, BBH17, BHEN13, CC14, JJW^{+}15, KDO14, KKS^{+}14, KS12, LMSN12, LLNL16, LMW^{+}17, MWMK19, MGDL10, PMG19, PCHY15, PLN10, PLM15, SRvdH12, TZM^{+}17, TZL^{+}15, TLZ^{+}18, WLX^{+}16, XXC^{+}15]. Detectors [ANN^{+}19]. Determination [YXS^{+}19]. Determined [SFCR16]. Determining [RE13]. Developed [CSJ^{+}16]. Developer [BGP^{+}19, DSSP17, DCS^{+}18, DPD^{+}18, Lat14, LNH^{+}16, RDT^{+}10, SMWO11]. Developers [BVD15, MBM^{+}17, MHZBN15, SJMH^{+}19]. Developing [BCE^{+}14, BSM11]. Development [BSH^{+}18, BCB^{+}17, BWB12, CBCB12, CEM11, CH13, CDT^{+}18, DSB14, EKAYW16, FET^{+}17, GPU^{+}16, HA16, HNM13, JH11, JG11, JG12, Lat14, LED13, MDB13, MBEN15, NDMX16, PBM^{+}19, Pre11, RM13, Ram14, SRvdH12, SCF19, SZF^{+}17, TS12, TRD15, WNM12, ZB12b]. Devices [BHCC11, LRCL14]. Diagnose [SJMH^{+}19]. Diagnosing [MPP11]. Diagnosis [ABL14, BPH10, LXCL14, TZM^{+}17]. Diagram [KRG^{+}13]. Diagrams [AGI^{+}13, CGC19, DCMS14, KKMY10]. Differences [IRBW16, LBK^{+}15]. Different [PHLW18]. Differential [TGH12]. Digital [VW12]. Directed [AL10, GSWH15b, YHW20]. Directives [MRG^{+}18]. Discipline [MRG^{+}18]. Disclosed [DBM19]. Discovering [BLZ^{+}19, BLZ^{+}20, LH14, MOd19, SPF14, SZ10]. Discovery [ANC^{+}10, MN14, ZKZ10, ZSDS13]. Discriminative [Al 11]. Dispersion [LED13]. Display [dSR11]. Dissatisfaction [NR19]. Dissection [FET^{+}17]. Distinguish [Ano11j, Ano11i]. Distributed [BGMM17, BHKZ13, ByBEM15, Hie12, Hie15, LAH^{+}14, MMMR12, MS18, Ram14, RNP13, VDC17].

Distribution [CMM^{+}11]. Distributions [GRH13, KA11]. Diverse [LR12, LLL^{+}18]. Diversity [BKP^{+}18, POPD15, SYB18]. Diversity-Aware [SYB18]. DividE [CPTT12]. DividE-and-conquer [CPTT12]. Do [LMZP19, NBKC15, ZB12a]. Documentation [DR14, KGM^{+}16, MR13, MJK^{+}18, TRD15]. Does [FET^{+}17, HKS12, KSD11, LKTC14]. Domain [ACDO12, AA11, ADJ16, ABD^{+}16, BGRJ16, BSL^{+}13, GHL^{+}13, HCHM^{+}13, MHL^{+}12, MFL^{+}16, MVDS20, MRD12]. Domain-Specific [ACDO12, AA11, BGRJ16, GHL^{+}13, MFL^{+}16, MVDS20, MRD12]. Domains [PW12]. Double [ABM19]. Downstream [PHLW18]. Driven [ABY16, BCE^{+}14, BMB^{+}15, BSM11, CCG^{+}12, CFMS15, FDMM18, FZ12, FET^{+}17, Lat14, LH14, MDB13, NM14, RM13, TFB14, WNM12, YDCP14]. DTMC [KA11]. Duplicate [RBH18]. during [SZ10]. Dynamic [AKD^{+}10, ADTP12, BTS17, BGMM17, CL16, Cai18, CGK^{+}11, CWG^{+}18, CYH^{+}11, GFM^{+}15, HSH^{+}12, HYH^{+}17, HCBB12, KS12, LM11, Li12, LGN12, LLNL16, LMZP19, MPP11, MNBB19, PKT18, RO13, RHB^{+}12, SMBY16, SMJ19, TZL^{+}15, WWY^{+}17]. Dynamics [BCB^{+}17].

Early [ABD^{+}16, BHEN13, KWK^{+}11, LH14, PCHY15, SRvdH12]. Early-Stage [ABD^{+}16]. EARMO [MSK^{+}18]. Eclipse [CMM^{+}11, DSB14]. Edit [LKKS15]. Editor [AI12, Kra10, Med18]. Editorial [Dwy15, DR16, Kra10, Med18, Med19, Nus10a, Nus10c, Nus10b, Nus10d, Nus11a, Nus11b, Nus13a, Nus13b, Nus13c, OT12]. Editors [Nus10d, Nus13b, CEM11, GRI10, GPK10, MS10]. Edits [ASM^{+}18]. Education [SMG11]. EDZL [LS13]. Effect [ACCA15, ADJ16, BTEGF12, KSD11, MHL^{+}12, SPA^{+}17, SYA^{+}13]. Effective [BDP16, Cai18, GKZ^{+}14, HHY^{+}14, TSBB20, WHYY18]. Effectively [HWM18, LKTC14]. Effectiveness [AGI^{+}13, ADJ16, ABL14, MHS^{+}13, ZB12a]. Effects [DMTR10, HAES10, KGM^{+}16, Lat14, NWCS20, RM13, SWEK15, YDCP14]. Efficiency [BCBR18, BP16a, HC18, LRCL14, PW12, PCHY15]. Efficient [BGMM17, BVD15, GRPF13, MPS17, PC10, TDD14, WMW11, WSC^{+}15]. Efficiently [SPF14]. Effort [DVMB12, JG11, KMBK12, KMK12, LMSN12, MJK^{+}18, MBC^{+}13, SYA^{+}13, YZL^{+}15, KMK^{+}13]. Effort-Aware [YZL^{+}15]. Efforts [KWK^{+}11]. Elaborating [AKRU13]. Elaboration [DLR13]. Electronic [BBC^{+}10, SRE10]. Eleven [SGR^{+}12]. Elicitation [ADJ16, CWW^{+}14, LF12, MLS^{+}14, DJ11]. Eliminating [YYG^{+}18]. Embedded [BBD17, BHCC11, CKMR13, CFMS12, LH12, MDB13]. Embedding [BGS^{+}13, CFMS15]. Emergent [Pre11]. Empirical [AA18, ABHPW10, Arc12, BH13, FCC13, Gar10, GS12, HM10b, HJ14, KZN14, MSKM10, MN14, NBKC15, PCHY15, POM13, SMG11, SPA^{+}17, SDdS16, SYB18, TMHM17, WNP^{+}20, YZL^{+}15, ZM19, DJ11, Her17, KSA^{+}13]. Empirically [SFCR16]. Employing [CC15]. Emulation [XXCL19, XXCL21]. Enabling [CKKR12]. Encoding [SZWZ12]. End [GKZ^{+}14, NDMX16]. Energy [BCBR18, BPHB19, DLWF19, LXCL14, MSK^{+}18]. Energy-Aware [MSK^{+}18]. Energy-Efficiency [BCBR18]. EnergyPatch [BCBR18]. Enforcement [MMPT19]. Enforcing [BGRJ16]. Engine [CDGC12, MGP^{+}12]. Engineering [ABY16, CWG^{+}18, Cha10, CNKM19, DBF^{+}16, DR16, Dwy17, EEM13, FDMM18, HBB^{+}12, HM10a, KKS^{+}14, KSD^{+}13, KMD15, KKR10, LBK^{+}15, LZX10, MSKM10, MSJ^{+}17, MG20, MB14, MS10, Ral19, RRS19, RCKH16, SDdS16, VAJ16, VCMG20, WGC^{+}16, Ano19a, AI12, Ano15a, Ano17a, Ano18a, HM10a]. Engineering-Utilizing [WGC^{+}16]. Engineers [FdS20, KBZ^{+}19]. Engines [MGS^{+}18, SGM13, ZXC16]. English [AGL^{+}15]. Enhanced [BBD17, EKAYW16]. Enhancing [YLQ^{+}18]. Ensemble [KMK12]. Enterprise [RK14]. Entropy [SSP18]. Entry [Thi15]. Environment [BDM18, CEM11, DSB14, LED13]. Environments [CGMD19, CBD^{+}17, RO13]. ePUB [Ano11q, Ano11r]. Equality [CT13]. Equally [PG13]. Equals [CT13]. Equivalence [BSM17, CT13]. Equivalences [KPJ^{+}18]. Equivalent [FCC13, MOTJ14]. ERP [SL12]. Erratum [XXCL21]. Error [XZH^{+}15]. Errors [HPSW16, MHB12]. Essential [KMBK12, KMK^{+}13]. Estimates [JG11]. Estimating [CDT^{+}19, PCPS15]. Estimation [DVMB12, IJK13, JG12, KKRD12, KMBK12, KMK12, LH12, MBC^{+}13, MA13, KMK^{+}13]. Ethnographic [SDdS16]. Evaluate [MN14]. Evaluating [BCD^{+}18, CSJ^{+}16, HSH^{+}12, KGM^{+}16, MOA15, PR11, RT12, SMWO11, SNAH15, dCMS^{+}17]. Evaluation [AA18, ANN^{+}19, BH13, BMB^{+}15, Gar10, GPK10, HBC13, MGB12, PLN10, RBH18, UGI^{+}12, YBL16, ZSDS13]. Event [BB15, BSM11, CZ13, CCP13, MC10, NM14, YJ13, YM10, YCM11]. Event-Based [BB15, CZ13, MC10]. Event-Driven [BSM11, NM14]. Evidence [GS12, MB14, dlVBWM16]. Evidence-Based [MB14]. Evolution [ABP12, CC15, CFMS15, DR14, HKB17, KA11, POM13, SMJ19]. Evolutionary [LKS10, MSY14, SFH^{+}17, SPG10]. Evolving [KGRX12, NNP^{+}12]. Exact [FAHCC14, LHWZ17]. Exception [BGRJ16, BG18, FFM^{+}10, GRI10, LCO^{+}10, SGH10]. Execution [CGMD19, CK15, HJ14, JMC13, LZX10, MSS19, QSM16, RGA^{+}15, SCP14, WLG^{+}17, YYG^{+}18]. Exemplar [MGP^{+}12]. Exercise* [NM14]. Existing [NDMX16, ZB12b]. Expanding [LLLW15, ZNKZ18]. Expectations [BLZ^{+}19, BLZ^{+}20]. Experience [AKL13, BBC^{+}10, Lat14, RDT^{+}10]. Experiment [ADJ16, BP16b, CZvD11, JG11, KSD^{+}13, KPJ^{+}16, MOTJ14, PR11, Pre11]. Experimental [RKH12]. Experiments [AGI^{+}13, DMTR10, JGP13, KSD^{+}13, MBLD11, RDT^{+}10, VAJ16]. Expert [HWM18]. Experts [SGH10]. Explicit [AL10, AKD^{+}10, WSW^{+}15]. Explicit-State [AKD^{+}10, WSW^{+}15]. Exploit [JPAB19]. Exploiting [BB15, KMBK12, RHB^{+}12]. Exploratory [IML13, Pre11]. Exploring [COLY15, DLWF19, NDMX16]. Explosion [HPP^{+}14]. Exposure [YJ13]. Expression [FPTO11]. Expression-Based [FPTO11]. Expressions [KGRX12, YT18]. Extended [ACH^{+}13, FAHCC14, SGR^{+}12]. Extending [EALK^{+}15, LGN12]. Extensibility [BBA^{+}15]. Extensible [MMMR12]. Extension [BSM17]. Extensive [LMZP19, PPZ^{+}18]. External [RM13]. Externalize [WZX^{+}13]. Extract [CAC^{+}17]. Extracted [HCBB12]. Extracting [BKML14, TRD15, VBP12]. Extraction [ASBZ17, NBKC15]. Eye [RLMM15]. Eye-Tracking [RLMM15].

Facilitating [BVD15]. Factors [NDMX16]. Failure [LP13b, NNL^{+}20, OY16]. Failure-Causing [NNL^{+}20]. Failures [CH13, CCP13]. Family [AGI^{+}13]. Farming [MBK^{+}13]. Fault [AB12, ADTP12, BPH10, BLVBC^{+}15, BTWV15, DVB13, GRH13, HBB^{+}12, MHS^{+}13, NCDM13, PW12, PCHY15, SER^{+}10, WHYY18, WGL^{+}16, YZL^{+}15]. Fault-Proneness [BLVBC^{+}15, YZL^{+}15]. Faults [ABL14, LLHF17, MPP11, NNL^{+}20]. Feature [EEM13, FAHCC14, MJC17, NSC^{+}12, YBL16]. Feature-Based [MJC17]. Feature-Oriented [EEM13]. Featured [CCS^{+}13]. FeatureHouse [AKL13]. Features [HCHM^{+}13, KDO14, MSCC13, PMR11, SWAK13]. Feedback [LLNL16, MBEN15, YDCP14, YM10]. Feedback-Based [LLNL16]. Fermat [WZ11]. Fidelity [DLWF19, YLQ^{+}18]. Field [HKS12, JG11, RK19, XBL^{+}18]. Files [YBL16]. Filtering [LF12, PTYN19]. Finding [AKD^{+}10, CSJ^{+}16, KMK^{+}13, LPZ12, MGP^{+}12]. Findings [PCHY15]. Fine [YBL16]. Fine-Grained [YBL16]. Finite [ACH^{+}13, HT16, Hie17]. First [DR16, FET^{+}17, GZS^{+}15, KWK^{+}11, NZ16]. Five [AGI^{+}13]. Fix [HMMB18, KWK^{+}11, KTKZ13, MK18]. Fix-Inducing [MK18]. Fixes [MHZBN15, XZH^{+}15]. Fixing [PFN^{+}14, SHCX16]. Flexible [CGMD19, EB10]. Floating [CCK14]. Floating-Point [CCK14]. FLOSS [ZM15]. Flow [LMST10, QSM16, SYX14]. Flowchart [YF11]. FlowTalk [BHCC11]. Fluid [HBC13, TDGH12, Tri13]. Fly [MPS12]. Foraging [LBB^{+}13]. Forecasting [CC15, SL12]. Formal [AB12, CMM^{+}19, LSL^{+}18, LCNM12]. Formalism [BBL^{+}11]. format [Ano11q, Ano11r]. Formulating [HHY^{+}14, WHYY18]. FOSS [DBM19]. Foundation [HPSW16]. Foundations [CCS^{+}13, MS10]. Four [RDT^{+}10]. Fragments [PBM^{+}19]. Framework [BKML14, CCG^{+}12, CSVC11, EEM13, FDMM18, HCBB12, JWDX17, KT10, LSL^{+}18, MMMR12, MMPT19, MVDS20, NRDB17, PWSG19, RK19, SJS^{+}11, TLB14, TLLV20, ZSDS13, dCMS^{+}17]. Framework-Usage [HCBB12]. Frameworks [AA11, CSJ^{+}16]. Friendly [MHB12]. Front [Ano10k, Ano10m, Ano10i, Ano10l, Ano10n, Ano10j, Ano11n, Ano11m, Ano11k, Ano11p, Ano11o, Ano11l, Ano12k, Ano12l, Ano12-27, Ano12-30, Ano12-29, Ano12z, Ano12-28, Ano12y, Ano10q, Ano10r, Ano10t, Ano10p, Ano10s, Ano10u, Ano11v, Ano11y, Ano11x, Ano11u, Ano11w, Ano11t, Ano12u, Ano12s, Ano12t]. Frontier [LLHF17]. FSC [WGC^{+}16]. Full [SYX14]. Full-Sparse [SYX14]. Fully [CGMM16, CGMD19, TFB14]. Functional [AGI^{+}13, ABL14, CAC^{+}17, eZHM13]. Functions [QSM16]. Future [DCS^{+}18]. Fuzzing [BPR19]. Fuzzy [PSB^{+}17, SL12].

Gain [BTEGF12]. GALE [KMD15]. Game [FRA^{+}19, HPSW16]. Game-Theoretic [HPSW16]. Gaussian [TWL13]. GEA [LH14]. General [RLR14, SJS^{+}11]. Generalize [KPJ^{+}16]. Generalized [WZX^{+}13]. Generate [AA11, HPP^{+}14, MBH^{+}18]. Generating [AIAB13, AM13, DKM^{+}12, GHL^{+}13, Hie15, HT16, YM10]. Generation [ABHPW10, AL10, AGR19, AKD^{+}10, BCD^{+}18, CRV13, FZ12, FA13, GSWH15b, JPAB19, LMP14, MNBB19, MHL^{+}12, MBP^{+}17, PZTM13, PKT18, SM16, UKB10, YT18, CPTT12]. Generator [Gar10, SPG15]. Generic [GITT17, LNFW12]. Genetic [AML11, BBL10, Gar10, KKR10, POPD15, PHLW18]. GenProg [LNFW12]. Geometric [KMD15]. GitHub [LMZP19]. GK [MPS17]. GK-Tail [MPS17]. Global [BG18, HM10b, LED13, MHL^{+}12, MBC^{+}13]. Global-Aware [BG18]. Globally [Ram14]. Glossary [ASBZ17]. Go [TPB^{+}17]. Goal [CWW^{+}14, DLR13, LH14]. Goal-Driven [LH14]. Goal-Obstacle [DLR13]. GoF [ACCA15]. Good [FRA^{+}19, ZPB^{+}10]. GoPrime [CGMM16]. Gossip [TLB14]. GossipKit [TLB14]. Governed [KA11]. Governing [Ram14]. Grained [YBL16]. Grammar [AGL^{+}15, ZKZ10]. Graph [BPH10, QSM16, ZKZ10]. Graphical [MBCC^{+}20]. Great [Ano10b, KBZ^{+}19]. GreenDroid [LXCL14]. Greenfield [GPU^{+}16]. Greybox [BPR19]. Guest [OT12, AI12, CEM11, GRI10, GPK10, MS10]. GUI [GCZM16, YM10, YCM11]. Guided [MMP15, WLG^{+}17]. Guidelines [Ral19].

Handling [ANVZ14, BGRJ16, BG18, FFM^{+}10, GRI10, LCO^{+}10, SGH10]. Hardware [WDS10]. Harmonised [Hie17]. Harrold [Nus13e]. Hashing [IJK13]. Heads [PSA^{+}20]. heaps [MSSF13]. Heavy [JGP13]. Heavy-Tailed [JGP13]. Hell [MRG^{+}18]. Heterogeneous [LJZ^{+}19, NFK^{+}18]. Hidden [HMMB18, Hat14]. Hierarchical [CPV13]. High [SPF14]. High-Coverage [SPF14]. Higher [SMG11]. Highly [MGP^{+}12, TLLV20]. Hinders [ZHMZ17]. Historical [FRGCH20]. Histories [LKKS15, LZRC18, PBP^{+}15]. House [PBS^{+}18]. Human [KSD^{+}13, SDMW14]. Human-Centric [KSD^{+}13]. Human-in-the-Loop [SDMW14]. Hybrid [BSH^{+}18, Cai18, HM10b, MHL^{+}12, MYK19, SBM13, TSBB20]. HYDRA [XLP^{+}16].

I/O [WSC^{+}15]. ICPMD [EKAYW16]. IDE [BGP^{+}19, DCS^{+}18, DSB14]. Identification [HC18, PBM^{+}19, SER^{+}10, SE13]. Identifier [AEP^{+}14, Zav17]. Identify [FRGCH20, QSM16]. Identifying [CMM^{+}19, CAC^{+}17, KNGW13, LLLW15, MSCC13, NNL^{+}20, dCMS^{+}17, CPTT12]. IDEs [RHB^{+}12]. Idioms [ABB^{+}18]. IDKL [ABM19]. IEEE [Ano10b, Ano10o, Ano11q, Ano11r, Ano12m, Ano12n, Ano12o, Ano19a, Ano15a, Ano17a, Ano18a, HM10a]. Imbalance [BKP^{+}18, JWDX17]. Imbalanced [SGS19]. Impact [AA18, BLVBC^{+}15, BWRR17, Cai18, CH13, CSJ^{+}16, JG11, LKKS15, LCG^{+}18, MFL^{+}16, PR11, RM17, SL12, SE13, TMHM19, TNBH13, dlVBWM16]. Impacted [FRGCH20]. Impediments [SWEK15]. Implementation [CSF^{+}16, MOA15].

Implementations [SCP14]. Implementing [BCD^{+}18]. Implications [AIB12]. Imprecise [PSB^{+}17]. Improve [AGA13, BCBR18, BBA^{+}15, CPR16, HC18, LW12, SPG15, SWAK13]. Improved [JWDX17, MSY14, RGA^{+}15]. Improvement [GPK10, HKS12, PHLW18, UGI^{+}12]. Improvements [ABM19, Ram14]. Improves [RHB^{+}12]. Improving [BH13, CKMR13, DDO11, Dwy17, MMMR12, POPD15, Thi15, XLD^{+}17]. In-House [PBS^{+}18]. Inclusion [WSWQ17]. Inconsistencies [Egy11]. Inconsistency [RE13, XXC^{+}15]. Incorporating [HWM18, YCM11]. Incremental [UKB10]. Independence [LP13a]. Independent [AKL13]. Index [Ano11a, Ano12a, Ano13a, Ano15a, Ano17a, Ano18a, Ano19a]. Indicators [SMWO11]. Individual [MSCC13]. Inducing [MK18]. Inductive [AKRU13]. Industrial [BWRR17, BCB^{+}17, FCC13, KKK11, MFL^{+}16, SGR^{+}12, dlVBWM16, dBFP^{+}13]. Industry [MLM^{+}13, MG20]. Ineffective [MWMK19]. Inefficiency [LXCL14]. Inference [BBA^{+}15, KNGW13, RBK^{+}13]. Inferring [GFM^{+}15]. Infinite [TZM^{+}17]. Infinite-Precision [TZM^{+}17]. Influence [RDT^{+}10]. Information [Ano10v, Ano10-27, Ano10y, Ano10z, Ano10x, Ano10w, Ano11-27, Ano11-31, Ano11-30, Ano11-29, Ano11-28, DDO11, Hat14, JG11, LMST10, LBB^{+}13, RHB^{+}12, SCF^{+}12, ZB12b, Ano12-31]. Initial [ZM15]. Initialization [RGA^{+}15]. Injecting [POPD15]. Injection [ANC^{+}10, JPAB19, NCDM13, TSBB20]. Inner [RCKH16]. Input [CRV13, CDF14, LZX10, MHL^{+}12, NM14]. Input-Sensitive [CDF14]. Insertion [CLP17]. Insights [BCB^{+}17]. Inspection [ASM^{+}18, LCNM12, WNM12]. Instability [TZM^{+}17]. Instance [SPG15, SMJ19].

Instant [LGS13]. Instruction [TZL^{+}15]. Instrument [BSD14]. Integer [JLR15, LW13]. Integrated [BDM18, CTZ11, EKAYW16, TSBB20]. Integrating [RK19]. Integration [MPP11, MPO^{+}19, SCF^{+}12]. Integrative [ABM19]. Intelligence [CD12]. Intensive [CCS^{+}13, PBS^{+}18]. Inter [BSGM15, WLX^{+}16]. Inter-App [BSGM15]. Inter-Procedural [WLX^{+}16]. Interact [MLPvdH15]. Interaction [AB12, CBD^{+}17, LNH^{+}16, PCHY15, RRS19, SPF14, Yil13, YDCP14, YCM11]. Interactions [DSSP17, PMM18]. Interactive [BXXL19, GKZ^{+}14, SPG10, XZH^{+}15]. Interdependence [JG12]. Interface [HMC19, MM13, PBU16]. Interfaces [ABY16, MBCC^{+}20, NWCS20, Thi15]. Interleaving [LPZ12]. Internally [ADJ16]. International [OT12, AI12]. Internationalization [WZX^{+}13]. Interoperability [BI15, MMPT10]. Interpretable [GBvBD^{+}18]. Interview [MG20, SGR^{+}12]. IntroClass [LHS^{+}15]. Introducing [MFL^{+}16, dCMS^{+}17]. Introduction [Ano12v, AI12, CEM11, GRI10, GPK10, HM10a, MS10, RZ10]. Invariant [BCD^{+}18, MvDR12]. Invariant-Based [MvDR12]. Invariants [GFM^{+}15]. Investigating [LBK^{+}15]. Investigation [ABHPW10, Sha10, SGS19]. Irrelevant [JG11, MHL^{+}12]. Isolation [PW12]. ISSTA [RZ10]. Issue [Ano11z, ANVZ14, BKP^{+}18, HM10a, RBH18]. Issues [LJZ^{+}19, Nus10c]. Item [TS12]. Iterative [CDT^{+}18, HYJY14]. iTree [SPF14].

Java [BH10, CC15, CSF^{+}16, LZ12, MM16, RLMM15, STST12, TNTN12, XMD^{+}17]. JavaScript [MMP15, OBPM17, PWSG19]. Jean [Nus13e]. Joining [Ano10b]. Joint [KPJ^{+}16]. Journal [Dwy15, DR16, Dwy18, Kra10, Med19, Nus11a, Nus12, Nus13c, Nus14]. Journal-First [DR16]. JUnit [MHZ^{+}12]. Just [MK18, KSA^{+}13]. Just-In-Time [MK18, KSA^{+}13].

Kaizen [ABM19]. KAOS [CWW^{+}14]. Keeping [DSB14]. Key [TZL^{+}15]. Keyword [HZZ^{+}17]. Kits [ACDO12]. Know [MHPB12, ZB12a]. Knowledge [ADJ16, HWM18, IML13, MR13].

Language [AKL13, ASBZ15, BGRJ16, BHCC11, BRLM11, DFG16, DS11, FCC13, GHL^{+}13, HJ14, HBS^{+}17, LMP14, LGN12, dSMST17, SCF^{+}12, WSWQ17, YF11]. Language-Based [BRLM11]. Language-Independent [AKL13]. Languages [AA11, BJGB10, BSL^{+}13, MMPT10, MLM^{+}13]. Large [BKML14, BSH^{+}18, CC14, JH15, KGM^{+}16, LF12, LLL^{+}18, MM13, WGC^{+}16, XBL^{+}18, XJRL12, dBFP^{+}13, KSA^{+}13]. Large-Scale [BSH^{+}18, CC14, JH15, LF12, LLL^{+}18, MM13, WGC^{+}16, XBL^{+}18, KSA^{+}13]. Last [FET^{+}17]. Latency [BHCC11]. layered [Tri13]. Lazy [RGA^{+}15]. Leading [SBL12]. Leak [WLX^{+}16]. Leakage [BSGM15]. Leaks [BCBR18, SYX14]. Learn [MPS17]. Learning [ABM19, AKRU13, BKKL10, BW10, CDT^{+}19, EEM13, GKZ^{+}14, KMD15, KM19, Lat14, LAL^{+}14, MBK^{+}13, MBCC^{+}20, PSA13, SBH14, SHB18, SGS19, TMHM16, KMK^{+}13]. Learning-Based [EEM13, MBCC^{+}20]. Legacy [MGB12]. Legal [MOA15]. LEILA [CMM^{+}19]. Length [Arc12]. Lens [BGM11]. Lessons [MBC^{+}13]. Level [CT13, MUA13, MN17, SRE10, WYZ^{+}18, YXS^{+}19]. Levels [Sha10]. Leveraging [FRGCH20]. Lexical [PLM15]. Lexicographical [ET16]. Lexicon [DDO11]. Library [QSM16]. Life [CRV11, MBM^{+}17]. Light [WLX^{+}16]. Light-Weight [WLX^{+}16]. Lightweight [OY16, ZWL^{+}16]. Like [SE13]. Line [ACDO12, KDO14]. Linear [LW13, LLC^{+}16]. Linearizability [LCL^{+}13, ZCW15]. Lines [HPP^{+}14, UKB10, VBP12]. Link [GS12]. Links [AGA13, DR14]. List [Ano10a, Ano11b, Ano12b, Ano13b, Ano14, Ano19b, Ano16]. List* [Ano15b, Ano17b, Ano18b]. Listening [SBR^{+}19]. Listings [HCHM^{+}13]. Literature [AA18, ABG^{+}13, GWT^{+}14, HBB^{+}12, HTG19b, MOTJ14, SMG11, SBWM18, UGI^{+}12, YGV17, vH19]. Live [GITT17]. Load [JH15, LKB18]. Load-Time [LKB18]. Loadings [KS12]. Local [HM10b, MHL^{+}12, MBC^{+}13, WWY^{+}18]. Localization [ADTP12, BTWV15, HOLL19, TNBH13, WGL^{+}16]. Locating [GW19, LLHF17, WZX^{+}13]. Location [MC10]. Location-Aware [MC10]. Logic [NZ16, PSB^{+}17]. Logs [CCP13]. Long [BHCC11]. Long-Latency [BHCC11]. Longitudinal [HKS12, MK18]. Look [PSA^{+}20]. Loop [ABM19, ABB^{+}18, GFM^{+}15, SDMW14, XCZ^{+}19]. Loupe [BGM11]. LTL [CCS^{+}13].

Machine [GKZ^{+}14, MBCC^{+}20, MBLD11, PSA13, SBH14, SHB18, TMHM16]. Machine-Based [MBLD11]. Machines [ACH^{+}13, CGMD19, HT16, Hie17]. MADMatch [KRG^{+}13]. Magazines [Ano11q, Ano11r]. Magiclock [CC14]. Magnifying [BGM11]. MAHAKIL [BKP^{+}18]. Maintainability [PR11]. Maintenance [GRI10, MFL^{+}16, RHB^{+}12, SL12, SYA^{+}13, WDZ^{+}14]. Make [Zav17]. Makes [KBZ^{+}19, Nus11b, ZPB^{+}10]. makeSense [MPO^{+}19]. Making [MOA15, MB14, SDMW14, TDD14]. Malicious [CMM^{+}19]. Management [BCE^{+}14, BRDG16, CGK^{+}11, HBS^{+}17, KL11, MBK^{+}13, NNP^{+}12, RK19, RO13, RRL^{+}13, SCF^{+}12]. Manager [KBZ^{+}19].

Managing [RK14]. Manual [ASM^{+}18]. Many [FTK^{+}17, KRG^{+}13, PKT18]. Many-Objective [FTK^{+}17, PKT18]. Many-to-Many [KRG^{+}13]. ManyBugs [LHS^{+}15, LBFW17]. Map [FDMM18]. Mapping [AHG13, BTS17, CSJ^{+}16, MTHG15, YBL16]. Maps [SL12]. Markov [BPR19, HMMB18]. MARTE [SBL12]. Mary [Nus13e]. Mashups [SE13]. Masking [YDCP14]. Massive [UR18]. Massively [XLP^{+}16]. Matching [KKMY10, KRG^{+}13, NSC^{+}12, PSB^{+}17]. Matter [FET^{+}17]. Matters [MRG^{+}18]. Maturity [SRS12]. Maximum [HPSW16]. MDD [VBJM14]. Measure [RT12]. Measurement [UGI^{+}12, WMW11]. Measuring [Al 11, BSD14, LW12, LCG^{+}18, XBL^{+}18]. mechanism [NB13]. Mediators [BI15]. Meets [OF13]. Memoriam [Nus10e, Nus13d, Nus13e]. Memory [ADG16, SYX14]. Merging [NSC^{+}12]. Message [BRLM11]. Meta [HTG19b, RM13]. Meta-Analysis [HTG19b, RM13]. Metamodels [HKB17]. Metamorphic [LKTC14, SFSRC16, SPTRC18, ZXC16]. Metering [CWW^{+}14]. Method [BOG^{+}14, BP16b, CAC^{+}17, CTZ11, KM19, LNFW12, MGDL10, PG13, SMBY16, SBL12, SCF19, WYZ^{+}18, KRX11]. Method-Level [WYZ^{+}18]. Methodbook [BOG^{+}14]. Methodological [Ral19]. Methodology [ABD^{+}16, BCE^{+}14, CBCB12, DSP11, KL11, MN14, CPTT12]. Methods [ABG^{+}13, BWB12, MM16]. Metric [BW10, MW11]. Metrics [Al 11, LNH^{+}16, Sha10, SMWO11, TWL13, YZL^{+}15, ZHMZ17]. Micro [LNH^{+}16]. Microsoft [BCB^{+}17, KZN14]. Microtask [LDR^{+}19]. Middleware [CGMM16, MC10, SCF^{+}12]. Migration [SMJ19]. Milner [Nus10e]. Minimum [LLHF17]. Mining [AGA13, ABB^{+}18, DKM^{+}12, DSSP17, DVMB12, HCHM^{+}13, KDO14, LW12, PBP^{+}15, SWHJ14, SRZ15, ZJ12]. Misalignment [BSH^{+}18]. Misleading [JG11]. Mission [CKMR13, HA16]. Mission-Critical [CKMR13]. Misuse [ANN^{+}19]. Mixed [BBD17]. Mobile [CMM^{+}19, CKMR13, EKAYW16, LBK^{+}15, MC10, MSK^{+}18, MBCC^{+}20, NDMX16, SBR^{+}19]. Modal [HOLL19, SBUK13]. Model [ABY16, AL10, AKRU13, AGR19, AKD^{+}10, BYW^{+}18, BDR^{+}20, BCE^{+}14, BB15, BDM18, BBA^{+}15, BKBR12, BMB^{+}15, BSM11, BTWV15, CXK^{+}19, CSF^{+}16, CYH^{+}11, CDT^{+}19, CCS^{+}13, CFMS12, CGdL14, CGdL17, CFMS15, DPD^{+}18, DPS12, DNT14, EALK^{+}15, FAHCC14, FTK^{+}17, FDMM18, GRH17, GPU^{+}16, HMMB18, HBS^{+}17, IJK13, KKK11, KTKZ13, LKR14, LMP14, LAH^{+}14, LO17, LL14, MJC17, MMPT10, MDB13, NM14, NZ16, PZTM13, PZF^{+}19, PHV16, POM13, RE13, SRS12, SBL12, SFCR16, TMHM17, TDD14, VBJM14, XLP^{+}16, XLD^{+}17, YBL16, Her17, Tri13, CRV11]. Model-Based [AGR19, BDM18, DNT14, GRH17, HBS^{+}17, LO17, PZTM13]. Model-Driven [ABY16, BCE^{+}14, BMB^{+}15, CFMS15, FDMM18, MDB13]. Model-Inference [BBA^{+}15]. Model-Transformation [LKR14]. Modeling [ACDO12, BJGB10, CB17, HBS^{+}17, LCO^{+}10, LKS10, MM13, RNP13, SWEK15, SDMW14, TLLV20, TWL13, VDC17, ZM15]. Modelling [MFL^{+}16, MVDS20]. Models [AM13, ABD^{+}16, BOG^{+}14, BSL^{+}13, CMCS12, CC15, CBD^{+}17, CSVC11, DCS^{+}18, DVB13, Egy11, FAHCC14, GAB^{+}10, HKB17, Kim14, KKR10, LM11, MPS17, MN14, MNBB19, MA13, PLM15, SWEK15, SRZ15, TMHM17, TMHM19, TGH12, WMW11, ZHMZ17, Her17].

Moderator [KPJ^{+}16]. Modern

[LMZP19, MvDR12, ZKB16]. Modification [VW12]. Modular [LAH^{+}14]. Modularization [FTK^{+}17]. Module [PHY11, VCMG20]. Monitor [LGS13, dSR11]. Monitor-Based [LGS13]. Monitoring [BHKZ13, BBH17, HHY^{+}14, MCH16]. Morphology [BB15]. MOSES [CCG^{+}12]. Motifs [WWY^{+}17]. Motifs-Based [WWY^{+}17]. Motivation [FdS20]. Move [BOG^{+}14]. Moving [MK18]. MSCs [BKKL10]. MSeer [GW19]. Multi [BBL10, CMCS12, HYJY14, HOLL19, KPJ^{+}16, MIA^{+}16, POPD15, PCPS15, PHY11, SFH^{+}17, SM16, SSP18, SCF19, TFB14, WHYY18]. Multi-Attribute [HYJY14]. Multi-Method [SCF19]. Multi-Modal [HOLL19]. Multi-Objective [BBL10, MIA^{+}16, POPD15, PHY11, SFH^{+}17, SM16, TFB14]. Multi-Site [KPJ^{+}16]. Multi-Tenant [WHYY18]. Multi-Threaded [PCPS15]. Multi-Tier [CMCS12]. Multi-Version [SSP18]. Multicore [ADG16, LS13]. Multicriteria [MAT12]. Multilevel [XJRL12]. Multimedia [AA11]. Multiple [GW19, LJZ^{+}19, LKS10, MA13, NNL^{+}20]. Multithreaded [CC14, Kim14, TLZ^{+}18]. Mutable [PW12]. Mutant [KPJ^{+}18, MOTJ14]. Mutants [MWMK19]. Mutation [BH13, FZ12, GFM^{+}15, JH11, MOTJ14, MUA13, MWMK19, MMP15, SYB18, ZZH^{+}19]. Mutation-Driven [FZ12]. My [dSR11].

Name [PG13]. Name-Based [PG13]. names [ZNKZ18]. Naming [STST12]. NASA [SSSM13]. Natural [ASBZ15, FCC13, LMP14, dSMST17]. Nature [AEP^{+}14]. Navigate [MHZBN15, TRD15]. Need [WZX^{+}13]. Need-to-Externalize [WZX^{+}13]. Needs [MLM^{+}13]. Nets [CRV11, DSP11].

Network [BRLM11, CC15, DVB13, HOLL19, SCP14, dSR11]. Network-Clustered [HOLL19]. Networks [LF12, MB14, MPO^{+}19, WWY^{+}17, WYZ^{+}18, Tri13]. Neutral [HYH^{+}17]. Newsletter [Ano12w]. Next [XJRL12]. NLP [CWW^{+}14]. NLP-KAOS [CWW^{+}14]. Non [WSW^{+}15, WSWQ17]. Non-Zenoness [WSWQ17, WSW^{+}15]. Nonequivalence [CT13]. Nopol [XMD^{+}17]. Normative [DCMS14, RK19]. Notation [EALK^{+}15]. Notations [vH19]. Notes [MBP^{+}17]. Notkin [Nus13d]. Novices [SGH10]. Number [Thi15]. Numerical [TZM^{+}17].

O [DCMS14, WSC^{+}15]. OBEY [JMC13]. Obfuscation [LMW^{+}17]. Obfuscation-Resilient [LMW^{+}17]. Object [Al 11, AA18, ABL14, BTS17, BBL10, CT13, CSJ^{+}16, Kim14, LZ12, MTHG15, POM13, SPG15, Sha10, SPG10]. Object-Oriented [Al 11, AA18, ABL14, BBL10, CT13, Kim14, MTHG15, POM13, Sha10, SPG10]. Object-Relational [BTS17, CSJ^{+}16]. Objective [BBL10, FTK^{+}17, MIA^{+}16, POPD15, PKT18, PHY11, SFH^{+}17, SM16, TFB14]. Objects [VW12]. Observable [CBD^{+}17]. Observation [ADG16]. Observe [NM14]. Observe-Model-Exercise* [NM14]. Obstacle [DLR13]. Occurrence [PLN10]. OCL [AIAB13, CGC19, LWY^{+}19]. Off [GAB^{+}10, Nus14]. On-the-Fly [MPS12]. One [LR12, LP13b]. Online [CB17, HCHM^{+}13, OY16, WWY^{+}17]. OnlinePlus [Ano12n, Ano12v, Ano12x]. Only [GKZ^{+}14]. Ontology [DLGRCBC15]. Ontology-Based [DLGRCBC15]. Open [Ano12o, BCB^{+}17, KGM^{+}16, LL14, PMM18, RT12, Sha10, SSP18]. Open-Source [Sha10]. Operations [BHCC11]. Opportunities [CAC^{+}17, LLLW15]. Optimal [BDR^{+}20, HZZ^{+}16, JMC13]. Optimisation [PKT18]. Optimisations [KPJ^{+}18]. Optimization [ABG^{+}13, CGK^{+}11, CZ13, FTK^{+}17, LKS10, MAT12, SM16, TMHM19]. Optimized [LCL^{+}13]. Optimizer [CNKM19]. Optimizing [KGDS13]. Options [LKB18]. Oracle [BHM^{+}15, GSWH15a, GKZ^{+}14, GZS^{+}15, LO17, LKTC14, PZTM13]. Oracles [ABL14, BPHB19, FZ12, GRH17, Hie12, LRCL14]. Order [MOTJ14, MUA13, NZ16]. Ordered [KGDS13]. Ordering [ET16]. Organisational [COLY15]. Organizations [ABM19]. Organizing [HNM13]. Oriented [Al 11, AA18, ABL14, BH10, BBL10, CCG^{+}12, CT13, EEM13, KGRX12, Kim14, MTHG15, MGB12, PR11, POM13, SPG15, Sha10, SPG10, WWY^{+}17]. Origins [PBS^{+}18]. OSS [PBS^{+}18]. Output [DLWF19, HLP10]. Outsourcing [PBS^{+}18]. Overcoming [MOTJ14]. Overly [MGS^{+}18]. Oversampling [BKP^{+}18]. Overtime [SFH^{+}17].

Packages [RK14]. Pair

[HAES10, SMG11, dBFP^{+}13]. Palantir [SRvdH12]. Palladio [BKBR12]. Papers [RZ10]. Paradigm [MLS^{+}14, NM14]. Parallel [ADG16, CCK14, GW19, HT16, KKS^{+}14, MYK19, SRvdH12]. Parameter [JLR15, TMHM19]. Parameterization [CMCS12]. Parameters [SFCR16]. Parametric [KKR10]. Partial [CBD^{+}17, Hie17]. Participant [ZM15]. Participatory [SZF^{+}17]. Paste [WDZ^{+}14]. Path [BSM17, BDR^{+}20, XCZ^{+}19, YYG^{+}18]. Pattern [PLN10]. Patterns [ACCA15, AGL^{+}15, BGP^{+}19, BBH17, KPJ^{+}16, LKR14, LCO^{+}10, LLL^{+}18, MR13, SER^{+}10, WGC^{+}16, ZB12a, ZB15]. PCM [DSP11]. PCM-Based [DSP11]. Peer [ZKB16]. Perceived [MBM^{+}17, SWEK15]. Perfect [LR12, LP13b]. Perform [LMZP19]. Performance [ADG16, ANVZ14, BMB^{+}15, CSJ^{+}16, COLY15, HBB^{+}12, HBC13, HBS^{+}17, KRX11, KKR10, MTHG15, RBH18, ZHMZ17]. Perils [VAJ16]. PerLa [SCF^{+}12]. Permission [BSGM15, BKML14]. Persistence [LZ12]. Persistent [POM13]. Persona [BWB12]. Personality [HAES10]. Personas [MOd19]. Perspective [LBB^{+}13, SCF19]. Perspectives [IRBW16]. Pert [LZ12]. Perturbation [SFCR16]. Pervasive [CBCB12, SCF^{+}12]. Petri [CRV11, DSP11]. pfd [LP13a]. Phase [HC18, KTKZ13]. Physical [BDM18, FRA^{+}19]. Physics [vH19]. Pipe [SE13]. Pipe-Like [SE13]. Plagiarism [JJW^{+}15, LMW^{+}17, TZL^{+}15, TLZ^{+}18]. Plan [JMC13]. Planning [BSH^{+}18, NR19, SFH^{+}17, SBR^{+}19]. Plat Forms [Pre11]. Platform [EKAYW16, HYH^{+}17, Pre11, RCKH16]. Platform-Based [RCKH16]. Platform-Neutral [HYH^{+}17]. Pnueli [Nus10e]. Point [CCK14]. Pointcut [KGRX12]. Points [CDT^{+}19]. Policies [BGRJ16, BTEGF12, MMPT19]. Policy [YLQ^{+}18]. Polychrony [BGS^{+}13]. Polynomial [TFB14]. Positive [SPA^{+}17]. Possible [GKZ^{+}14]. Possibly [LR12, LP13b]. Post [YZL^{+}15]. Post-Release [YZL^{+}15]. Postconditioned [YYG^{+}18]. Potential [CC14, SJMH^{+}19]. Power [Al 11]. Practical [AIB12, LH12, PCHY15]. Practice [BCB^{+}17, SGR^{+}12, dlVBWM16]. Practices [DSB14, VCMG20]. Practitioners [MG20]. Precise [SZWZ12]. Precision [TZM^{+}17]. Preconditions [LLC^{+}16, MGS^{+}18]. Predict [HMMB18]. Predicting [CDT^{+}18, DCS^{+}18, KWK^{+}11, KRX11, SWHJ14, WDZ^{+}14]. Prediction [BKP^{+}18, BKBR12, Cai18, DVB13, DPD^{+}18, HBB^{+}12, HTG18, HTG19a, HTG19b, JWDX17, KKR10, LNH^{+}16, LJZ^{+}19, MK18, MBC^{+}13, MHS^{+}13, NFK^{+}18, OY16, PZF^{+}19, PMGZ13, PTYN19, SBH14, SHB18, SWAK13, SJS^{+}11, SGS19, TMHM16, TMHM17, TMHM19, WWY^{+}17, XLP^{+}16, YZL^{+}15, YWHH19, ZHMZ17, Her17]. Predictive [ZZH^{+}19]. Preemptive [CRV11]. Preprocessor [MRG^{+}18]. Presence [NNL^{+}20]. Preservation [LJZ^{+}19]. Preventing [LYJC11, RM17]. Principles [FCC13]. Prioritization [BSM11, DMTR10, ET16, eZHM13, HZZ^{+}16, LMZP19, MNBB19, PSA13]. Prioritize [HPP^{+}14, KWK^{+}11, MIA^{+}16]. Prioritizing [MJK^{+}18, MHZ^{+}12]. Privacy [LJZ^{+}19, PMGZ13, YLQ^{+}18]. Proactive [CKMR13, SMBY16, ZSDS13]. Proactive-Rescheduling [SMBY16]. Probabilistic [AGL^{+}15, BPH10, BP16a, PHV16, PBU16, SFCR16]. Probability [AB12, KA11, LMST10, LP13a, LP13b, SFCR16]. Probability-Based [LMST10]. Probes [HBC13]. Problem [BHM^{+}15, BBL10, FTK^{+}17, LKTC14, MOTJ14, MSY14, PKT18, PHY11, SPG15, XJRL12]. Problems [ADG16, EB10, JWDX17]. Procedural [WLX^{+}16]. Process [BJGB10, BCB^{+}17, COLY15, FET^{+}17, HKS12, HJ14, LED13, LMP14, LCO^{+}10, MPS12, PLM15, Ral19, Ram14, SWEK15, SRS12, TDGH12, TGH12, UGI^{+}12, WMW11]. Processes [BG11, DLvL14, FFM^{+}10, GBvBD^{+}18, MPO^{+}19, RK19, SDMW14, SMJ19, TWL13]. Processing [ASBZ15, BRLM11, FCC13, dSMST17]. Product [ACDO12, DLGRCBC15, HCHM^{+}13, HPP^{+}14, KDO14, Ram14, RCKH16, UKB10, VBP12]. Product-Line [KDO14]. Productivity [CH13, MBM^{+}17, RM13]. Professionals [XBL^{+}18]. Profiles [LLL^{+}18, WMW11]. Profiling [BBH17, CDF14]. Program [ARM18, BPH10, Cai18, CTZ11, CZvD11, DFG16, GRH17, JJW^{+}15, LZX10, NRDB17, OLJM16, SBGM17, TLZ^{+}18, XBL^{+}18, ZKZ10, MSSF13]. Programmer [MHB12, RKH12]. Programmer-Friendly [MHB12]. Programmers [ARM18, LBB^{+}13, PSA^{+}20, RLMM15]. Programming [BXXL19, BSD14, HAES10, KT10, LDR^{+}19, LW13, MS18, OF13, SMG11, SPA^{+}17, YF11, dBFP^{+}13]. Programming-Based [LW13]. Programs [BBH17, CC14, GITT17, LHS^{+}15, LCNM12, LLC^{+}16, PFN^{+}14, STST12, XMD^{+}17]. Project [CZ13, HTG18, HTG19a, HTG19b, JWDX17, KGM^{+}16, KSD11, KL11, MBK^{+}13, MSY14, RT12, SMBY16, XLP^{+}16]. Projects [CEM11, LMZP19, SFH^{+}17]. Proneness [BLVBC^{+}15, SJS^{+}11, SNAH15, YZL^{+}15]. Proofs [BNR^{+}10]. Propagation [OLJM16]. Properties [GAB^{+}10, Pre11, RLR14]. Property [AGL^{+}15, LW13, RBK^{+}13]. Protection [PW12]. Protocol [BBD17, BRLM11, MDB13, SCP14]. Protocols [LSL^{+}18, MPS12]. Prototyping [CD12, MBCC^{+}20]. Provably [CD12]. Provenance [MBH^{+}18]. Providing [MMPT10]. Proving [CTZ11, CTZ11]. Publication [DR16, Nus11b]. Publish [BGM11, VDC17]. Publishing [Ano12o, Dwy17]. Putting [CRV11]. PVS [BDM18]. PVS-Simulink [BDM18].

QoS [CGK^{+}11, CCG^{+}12, CB17, Li12]. Qualitative [AGL^{+}15, MN17, SBGM17]. Quality [AA18, ANVZ14, BH13, HYJY14, KSD^{+}13, LW12, LB10, LKS10, RM13, RK19, RM17, SSSM13, SGR^{+}12, TFB14, ZXC16, KSA^{+}13]. Quality-Aware [HYJY14]. Quality-Driven [TFB14]. Quantifying [SYA^{+}13]. Quantitative [BMB^{+}15, CRV13, FTG16, GRH13, Sha10, TLLV20]. Quantum [YF11]. Quasi [KSD^{+}13, ZCW15]. Quasi-Experiment [KSD^{+}13]. Queries [ZNKZ18]. Queueing [LL14, Tri13].

Race [BH10]. Random [AB12, AIB12, STM13, WNP^{+}20, ZJ12]. Randomized [AML11, JG11]. Range [XZH^{+}15]. Ranking [MA13, PTYN19, YBL16]. Rapid [CD12]. Rate [LL14]. Rate-Based [LL14]. Ratings [BLVBC^{+}15]. Re [SNAH15, TLLV20]. Re-Evaluating [SNAH15]. Reactive [BRDG16, LW13, MS18, SPA^{+}17, ZSDS13]. Readability [BP16b, BW10]. Readers [Nus10d]. Readiness [MOA15]. Real [AM13, AGL^{+}15, BBC^{+}10, BTEGF12, CPV13, CRV13, EB10, GRH17, JLR15, Kim14, LS13, YWHH19]. Real-Time [AM13, AGL^{+}15, BTEGF12, CPV13, CRV13, EB10, JLR15, Kim14, LS13]. Real-World [BBC^{+}10, YWHH19]. Really [FET^{+}17]. Reasoning [LR12, LP13b, MBK^{+}13, Zav17]. Reasons [Ano10b]. Recommend [DR14]. Recommendation [BWRR17, KTKZ13]. Recommendations [BG18, BVD15, LKKS15]. Recommending [BOG^{+}14, HCHM^{+}13, ZKB16]. Reconfiguration [Li12]. Recovering [KGRX12]. Recovery [LCG^{+}18]. Redistribution [JMC13]. Reduce [HKS12]. Reduced [QSM16]. Reducing [MBEN15, SWAK13, VW12, YDCP14]. Reduction [MSS19, MPS12, MHL^{+}12, TdlRSCB16, WNM12]. Reductions [RLR14]. Redundancy [YYG^{+}18]. Redundant [CSJ^{+}16]. Refactor [MHPB12]. Refactorability [TMK15]. Refactoring [AA18, ASM^{+}18, CAC^{+}17, HC18, JMC13, KZN14, LGS13, LWY^{+}19, MRG^{+}18, MGS^{+}18, MSK^{+}18, MGB12, MHB12, STST12, SGM13, SE13, WYZ^{+}18]. Refactorings [BOG^{+}14, LLLW15]. Reference [MR13]. Refinement [ABD^{+}16, LCL^{+}13]. Reflecting [IRBW16]. Reflective [CGMD19]. Regenerative [PHV16]. Regression [MAT12, SBM13]. Rejuvenation [KGRX12]. RELAI [CPR16]. Related [ZNKZ18]. Relation [ANVZ14]. Relational [BTS17, BOG^{+}14, CSJ^{+}16, MWMK19]. Relations [BSM17, CPTT12]. Relationship [COLY15]. Relative [SNAH15]. Relaxed [CYH^{+}11]. Release [MBP^{+}17, NR19, SBR^{+}19, XJRL12, YZL^{+}15]. Relevance [CAC^{+}17]. Relevant [MGP^{+}12, YBL16]. Reliability [BKBR12, ByBEM15, CKMR13, CPR16, JGP13, LR12, LLC^{+}16, LYC14, PW12, PRT10, SSP18, TWL13, WWY^{+}17]. Reliable [MSKM10]. Remodularisation [HWM18]. Removal [MHL^{+}12, MWMK19]. Rename [LLLW15]. Renaming [LLLW15]. Renamings [AEP^{+}14]. Repair [FFM^{+}10, GCZM16, GMM19, LNFW12, LHS^{+}15, MJC17, XMD^{+}17]. Repairing [BCBR18, BG18, DXCW17]. REPENT [AEP^{+}14]. Replicated [ADJ16, GRH13]. Replicating [SNAH15]. Replication [KPJ^{+}16]. Reply [SHB18]. REPM [VCMG20]. Report [PTYN19, ZPB^{+}10]. Reporting [YGV17]. Reports [RBH18, RMM14, YBL16]. Repositories [AGA13, LKS10]. Representation [SPG15, SBM13, SWEK15]. Representative [MOd19]. Representativeness [NCDM13]. Representing [CXK^{+}19]. Reproduce [WWY^{+}18]. Reproduction [CK15]. Requirement [AGA13, WDZ^{+}14]. Requirements [AGI^{+}13, AKRU13, ASBZ15, ASBZ17, DLR13, FCC13, FRGCH20, GRI10, Gar10, IRBW16, LHWZ17, LF12, MSS19, MG20, MLS^{+}14, PCPS15, PSA13, PSB^{+}17, RM17, SGR^{+}12, VCMG20, WGC^{+}16]. Rescheduling [SMBY16]. Research [Dwy17, FDMM18]. Researcher [SBH14, SHB18, TMHM16]. Resilience [HPSW16]. Resilient [LMW^{+}17]. Resolution [LMSN12, PLM15, XZH^{+}15]. Resource [BCBR18, HBS^{+}17, RO13, WLX^{+}16].

Resources [VDC17]. response [Her17]. Responsibility [BBL10, JMC13]. RESTful [SPTRC18]. Results [AGI^{+}13, AIB12, dCMS^{+}17]. Retrieval [DDO11, RBH18]. Retrieving [FCC13]. Reusability [LRCL14]. Reverse [KKR10, LZX10]. Review [AA18, ABG^{+}13, ABHPW10, BCB^{+}17, GWT^{+}14, HBB^{+}12, HTG19b, MOTJ14, RRS19, SMG11, SBWM18, UGI^{+}12, YGV17, ZKB16, vH19, DJ11]. Reviewers [Ano10a, Ano11b, Ano12b, Ano13b, Ano14, Ano15b, Ano17b, Ano18b, Ano19b, Nus10d, ZKB16, Ano16]. Reviews [MSKM10]. Revisit [CT13]. Revisited [LBB^{+}13]. Revisiting [RBH18]. Reviving [TLZ^{+}18]. Revolution [BXXL19]. Rewards [TDGH12]. Rich [MOd19]. Rigorous [MMPT19]. Risk [KL11, SL12, Sha10]. Risks [BHEN13, GSWH15b]. Robin [Nus10e]. Robust [DS11, KT10]. Role [Arc12, BP16b, IML13, JG12, SDdS16, SGS19]. Roles [HNM13]. Round [ZCW15]. Round-Up [ZCW15]. Rule [CXK^{+}19, CCP13, KNGW13, SCP14]. Rule-Based [CCP13, SCP14]. Run [FTG16]. Runtime [CCG^{+}12, JJW^{+}15, MSY14, YM10, ZCW15, ZSDS13, MSSF13].

Safer [Thi15]. Safety [BH13, BWRR17, HSH^{+}12, MG20, VCMG20, dlVBWM16]. Safety-Critical [BH13, BWRR17, MG20]. Sampling [CNKM19]. SAT [GRPF13, RGA^{+}15]. SAT-Based [GRPF13]. Satisfaction [FdS20, NR19]. Save [LMSN12]. Scalability [DLR13]. Scalable [CC14, FAHCC14, TGH12, XXCL19, XXCL21]. Scale [BKML14, BSH^{+}18, CC14, JH15, KGM^{+}16, LF12, LLL^{+}18, MM13, UR18, WGC^{+}16, XBL^{+}18, XJRL12, KSA^{+}13]. Scenarios [EB10, SBL12, SBUK13]. Scent [PPZ^{+}18]. Schedulability [Kim14, LS13]. Schedulability-Aware [Kim14]. Schedule [LMSN12]. Scheduler [CZ13]. Schedulers [KGDS13]. Scheduling [BTEGF12, CPV13, CZ13, LS13, MSY14, SMBY16]. Schema [POM13]. Schemas [MWMK19, NNL^{+}20]. Scheme [TFB14]. Scientific [LYJC11]. Scientists [KZDB18]. Scope [WGC^{+}16]. ScottKnottESD [Her17]. Screening [DBM19]. Script [GCZM16]. Scrum [HA16]. SDA [JWDX17]. SE [SMG11]. Search [ABHPW10, AIAB13, AL10, BPHB19, CNKM19, EB10, HM10a, HM10b, HZZ^{+}17, KKS^{+}14, KMD15, KKR10, LWY^{+}19, MGP^{+}12, MHL^{+}12, PC10, PHY11, RRS19, SPG10, XXCL19, XXCL21, ZNKZ18, ZXC16]. Search-Based [ABHPW10, CNKM19, HM10b, KKS^{+}14, KMD15, MHL^{+}12, RRS19]. Searching [Pre11]. Second [GRH13, MOTJ14, MUA13]. Second-Order [MUA13]. Section [CEM11, GRI10, GPK10, OT12]. Secure [BCE^{+}14]. Security [BBC^{+}10, EALK^{+}15, FRA^{+}19, GAB^{+}10, LSL^{+}18, PTYN19, SBGM17, SDMW14, SJMH^{+}19, TSBB20]. Seer [OY16]. Selection [BDR^{+}20, CGC19, DNT14, GSWH15a, GKZ^{+}14, HYJY14, LHWZ17, LB10, MAT12, MRD12, POPD15, PKT18, TFB14]. Self [BG11, ByBEM15, CWG^{+}18, CB17, CKMR13, EEM13, FTG16, HNM13, HBS^{+}17, dSMST17, RRL^{+}13, YXS^{+}19, NB13]. Self-Adaptation [CKMR13, FTG16, NB13]. Self-Adapting [ByBEM15]. Self-Adaptive [CWG^{+}18, CB17, EEM13]. Self-Admitted [dSMST17, YXS^{+}19]. Self-Aware [HBS^{+}17]. Self-Management [RRL^{+}13]. Self-Organizing [HNM13]. Self-Supervising [BG11]. Semantic [ABB^{+}18, LZRC18]. Semantically [DPS12, ZNKZ18]. Semantics [Cha10, DFG16, LGN12, LMW^{+}17, MS18]. Semantics-Based [LMW^{+}17]. Semi [BWB12, KDO14, VBP12, CTZ11]. Semi-Automated [BWB12].

Semi-Automatic [VBP12, KDO14]. Semi-Proving [CTZ11]. Sensitive [CDF14, HLP10]. Sensitivity [FTG16, NRDB17]. Sensor [MPO^{+}19]. Sequence [AGI^{+}13, Arc12, YM10]. Sequence-Based [YM10]. Sequences [DSSP17, HT16, TZL^{+}15]. Sequential [TLZ^{+}18]. Series [DMTR10, MBLD11, RDT^{+}10]. Server [GITT17]. Service [CGK^{+}11, CGMM16, CCG^{+}12, CKKR12, FFM^{+}10, HHY^{+}14, HYJY14, HZZ^{+}17, KKRD12, LB10, LLL^{+}18, MPS12, MRD12, PMG19, PR11, SRE10, SZ10, TFB14, WWY^{+}17, WHYY18, ZSDS13, NB13]. Service-Based [CGK^{+}11, FFM^{+}10, HHY^{+}14, HYJY14, HZZ^{+}17, SZ10, WHYY18, NB13]. Service-Level [SRE10]. Service-Oriented [CCG^{+}12, PR11, WWY^{+}17]. Services [ABP12, CB17, CDGC12, DPS12, MRD12, PSB^{+}17, SRE10, SZ10]. Serving [DBF^{+}16]. Session [JGP13]. Sessions [KRX11]. Set [LLHF17]. Settings [BSH^{+}18]. Severity [HKS12]. Shape [SZF^{+}17]. Shared [ADG16]. Shared-Memory [ADG16]. Shortest [BDR^{+}20]. Shortest-Path [BDR^{+}20]. Should [KWK^{+}11, KTKZ13, Nus10c, dSR11]. Side [MSCC13, OBPM17]. Sighted [ARM18]. Signing [Nus14]. Similarity [HPP^{+}14, LMW^{+}17]. Simplifying [MPO^{+}19]. Simulation [BMB^{+}15, LL14]. Simulator [MLS^{+}14]. Simulink [BDM18, MNBB19]. Single [BSM11]. SITAR [GCZM16]. Site [KPJ^{+}16]. Situated [DSB14]. Six [BJGB10]. Size [LH12, MAT12]. Size-Constrained [MAT12]. Sketches [MLPvdH15]. Skill [BSD14]. Slice [YZL^{+}15]. Slice-Based [YZL^{+}15]. Slicing [ACH^{+}13, HLP10, LM11, LZRC18, TSBB20, WZ11]. Smart [CWW^{+}14, CGC19]. Smartphone [LXCL14]. Smell [LMSN12, LLNL16, PPZ^{+}18, PZF^{+}19, TPB^{+}17]. Smell-Aware [PZF^{+}19]. Smells [DSSP17, KKS^{+}14, MGDL10, PBP^{+}15, PPZ^{+}18, SYA^{+}13, SE13, TPB^{+}17]. SMT [CFMS12]. SMT-Based [CFMS12]. SOA [YJ13]. Social [BCB^{+}17, LF12, PMM18, SZF^{+}17]. Society [Ano10o, Ano12n, Ano10b, Ano11q, Ano11r, Ano12m]. Socio [CEM11, KSD11]. Socio-Technical [CEM11, KSD11]. Soft [BTEGF12]. Soft-Task [BTEGF12]. Software [ABM19, ABG^{+}13, AGA13, ACDO12, Ano15a, Ano17a, Ano18a, Ano19a, AKL13, ABL14, Arc12, AI12, BH13, BHM^{+}15, BBD17, BJGB10, BKP^{+}18, BSL^{+}13, BMP18, BSH^{+}18, BBL^{+}11, BVD15, BP16a, BP16b, BPHB19, ByBEM15, BSM11, CXK^{+}19, CWG^{+}18, CMSSS13, CEM11, CH13, CC15, CGMD19, CYH^{+}11, CZ13, CT13, CB17, CNKM19, CDT^{+}18, CCP13, COLY15, CKMR13, CFMS12, CPR16, CSVC11, DVMB12, DVB13, DS11, DLWF19, DBF^{+}16, DR16, Dwy17, Egy11, EEM13, FdS20, FDMM18, GWT^{+}14, GRI10, GMM19, GPU^{+}16, GPK10, GRH13, HBB^{+}12, HWM18, HM10a, HMC19, HKS12, HA16, Hat14, HSH^{+}12, HPP^{+}14, HNM13, HPSW16, IML13, JJW^{+}15, JH15, JG11, JG12, KBZ^{+}19, KKS^{+}14, KGRX12, KKK11, KZDB18, KSD^{+}13, KMD15, KSD11, LNFW12, LED13, LNH^{+}16, LAH^{+}14, Li12, LLHF17, LZRC18, LBK^{+}15, LB10, LL14, LH12, LP13a]. Software [LP13b, LKS10, LGS13, LLC^{+}16, LMW^{+}17, LMZP19, LCG^{+}18, LYC14, MSKM10, MMMR12, MLPvdH15, MTHG15, MPS17, MSJ^{+}17, MFL^{+}16, MB14, MA13, MDB13, MS10, MBEN15, NCDM13, NWCS20, NNP^{+}12, NZ16, OT12, PR11, PSA13, PBS^{+}18, PHLW18, POM13, PRT10, PSB^{+}17, PBM^{+}19, PC10, PHY11, Ral19, Ram14, RK14, RK19, RRS19, RM17, RHB^{+}12, SBGM17, SPA^{+}17, SFH^{+}17, SWHJ14, SDdS16, SMBY16, SSSM13, SBH14, SHB18, SBWM18, SMWO11, SSP18, SJS^{+}11, SGS19, SCF19, SZF^{+}17, TZM^{+}17, TMHM16, TDD14, TZL^{+}15, TLZ^{+}18, TWL13, TS12, TRD15, TMK15, UGI^{+}12, UKB10, VBP12, VAJ16, WZX^{+}13, WYZ^{+}18, WDS10, WGL^{+}16, XZH^{+}15, YHW20, ZB12a, ZHMZ17, ZXC16, dBFP^{+}13, KMK^{+}13]. Software-Based [BBL^{+}11, LP13a, LP13b]. Solutions [BKML14, EB10, RNP13]. Solvers [NZ16]. Solving [BBL10, TSBB20, XJRL12]. Some [SSSM13]. Soon [Ano12n, Ano12x]. Sound [Cha10]. Source [BCB^{+}17, DDO11, FRGCH20, FCR^{+}17, KGM^{+}16, LL14, MM16, MGP^{+}12, PMM18, RT12, RCKH16, RLMM15, Sha10, SSP18, UR18]. Sources [LJZ^{+}19]. Space [AL10, MHZBN15]. Spaces [BTS17, BPHB19, CXK^{+}19, NM14, Zav17]. Sparse [SYX14]. Special [CEM11, GRI10, GPK10, HM10a, Nus10c, OT12]. Specialising [PHLW18]. Specialized [XLD^{+}17]. Specific [ACDO12, AA11, BGRJ16, GHL^{+}13, MFL^{+}16, MVDS20, MRD12]. Specification [AGL^{+}15, BBA^{+}15, DKM^{+}12, DCMS14, HBC13, LW12, LSL^{+}18, LCNM12, MMPT19, MLS^{+}14, MGDL10, PMG19]. Specification-Based [LCNM12]. Specifications [GHL^{+}13, NSC^{+}12, NZ16, PSB^{+}17, SCP14]. Specifying [LGN12]. Speed [RHB^{+}12]. Spreadsheets [CFMS15, DXCW17]. Stability [ACCA15]. Stack [CK15]. Staffing [CZ13]. Stage [ABD^{+}16, KWK^{+}11]. Stakeholder [IRBW16]. StakeRare [LF12]. STAR [CK15]. Starts [TPB^{+}17]. Startup [GPU^{+}16]. Starvation [SBL12]. State [AL10, ACH^{+}13, AKD^{+}10, Dwy15, Dwy18, HT16, Hie17, KZDB18, Kra10, Med19, MBLD11, Nus11a, Nus12, Nus13c, Nus14, WSW^{+}15, YM10]. State-Space [AL10].

Statecharts [EALK^{+}15]. Statement [XMD^{+}17]. Static [ANN^{+}19, BKML14, BTWV15, CGdL17, GFM^{+}15, LMZP19, MHZ^{+}12, SJMH^{+}19]. Statically [SYX14]. Statistical [PC10]. Stay [ZM15]. Steering [GRH17]. Stem [NBKC15]. Stereotypes [RDT^{+}10]. Stochastic [AL10, CRV13, HBC13, PMR11, TZM^{+}17, TDGH12]. Store [MSJ^{+}17]. Story [CDT^{+}19]. Strategies [ARM18, BSM11, FPTO11, HHY^{+}14, LO17, WHYY18]. Strategy [LYC14, RKH12]. Stress [Gar10]. Stressing [EB10]. String [SM16, WZX^{+}13]. String-Taint [WZX^{+}13]. Strings [WZX^{+}13]. Strong [MGS^{+}18]. Structural [Arc12, MHL^{+}12, MBLD11, PPZ^{+}18, RKH12]. Structure [LZX10]. Structured [AGL^{+}15]. Structures [eZHM13, VDC17, ZCW15]. Studies [HCBB12, SMG11, SDdS16, DJ11]. Studio [DSSP17]. Study [AA11, ACCA15, BSL^{+}13, BSH^{+}18, BVD15, BWRR17, CWW^{+}14, DVMB12, DSB14, FCC13, FRA^{+}19, HM10b, HKS12, HTG18, HTG19a, KGM^{+}16, KKK11, KZN14, KSD11, LMZP19, MG20, MK18, MFL^{+}16, MBK^{+}13, MN17, MLS^{+}14, NBKC15, OBPM17, OF13, PMM18, POM13, RLMM15, SPA^{+}17, SYB18, SCF19, SGR^{+}12, Thi15, WSW^{+}15, XBL^{+}18, YZL^{+}15, ZM19, ZXC16, dBFP^{+}13, KSA^{+}13]. Subjective [LP13a]. Subscribe [BGM11, VDC17]. Success [KSD11]. Suite [FA13, MSS19, YHW20]. Suites [eZHM13, Hie15]. Summarization [FCR^{+}17, MM16, RMM14, RLMM15, XCZ^{+}19]. Summarizing [KNGW13]. Summation [ZHMZ17]. Supervising [BG11]. Support [ASM^{+}18, BI15, BHCC11, DS11, GSWH15a, MVDS20, POM13, RGA^{+}15]. Supported [ABD^{+}16, RDT^{+}10]. Supporting [BWRR17, DNT14, FTG16, HCHM^{+}13, LMP14, WGC^{+}16]. Suppressing [XXC^{+}15]. Surface [MW11]. Survey [BHM^{+}15, BMP18, GMM19, HKB17, JH11, JH15, MLM^{+}13, MSJ^{+}17, PBS^{+}18, SFSRC16, WGL^{+}16, dlVBWM16]. Survivability [RT12]. Sustainable [ABM19]. SW [CRV11]. Swarm [HJG11]. Switches [MBM^{+}17]. SymbexNet [SCP14]. Symbolic [AM13, BDP16, CK15, CCK14, RGA^{+}15, SCP14, WLG^{+}17, YYG^{+}18]. Symmetry [GRPF13]. Symposium [OT12]. Synchronization [SHCX16]. Synchrony [BGS^{+}13]. Synergy [BPHB19]. Syntactic [LZX10]. Synthesis [BTS17, BI15, JLR15]. Synthesizing [Kim14, SBUK13]. System [BXXL19, BWRR17, BBH17, CWW^{+}14, CMM^{+}11, GAB^{+}10, HYH^{+}17, LP13a, LP13b, MMMR12, MUA13, MBH^{+}18, PW12, SZ10, ZWL^{+}16]. Systematic [AA18, ABG^{+}13, ABHPW10, ANN^{+}19, DJ11, DLR13, GWT^{+}14, HBB^{+}12, HSH^{+}12, HTG19b, KNGW13, MSKM10, MOTJ14, MTHG15, RRS19, SMG11, SBWM18, UGI^{+}12, WSW^{+}15, YGV17, vH19]. Systems [AM13, ADG16, BBC^{+}10, BHKZ13, BSL^{+}13, BDM18, BRDG16, BMP18, BBL^{+}11, ByBEM15, BTEGF12, CGK^{+}11, CCG^{+}12, CPV13, CRV13, CWW^{+}14, CCS^{+}13, EEM13, FRA^{+}19, GWT^{+}14, GRH13, GKZ^{+}14, HHY^{+}14, HYJY14, HZZ^{+}17, JGP13, JH15, JLR15, KRX11, LW13, LAH^{+}14, Li12, LAL^{+}14, LR12, LMZP19, MG20, MDB13, NM14, PZTM13, PHV16, PBS^{+}18, RNP13, SCF^{+}12, Sha10, SBL12, SBUK13, TLLV20, WWY^{+}17, WHYY18, ZB12b]. SZZ [dCMS^{+}17].

Table [Ano12-27, Ano12-30, Ano12-29, Ano12z, Ano12-28, Ano12y]. Tabular [FPTO11]. TACO [GRPF13]. Tactics [MCH16]. Tagging [TS12]. Tail [MPS17]. Tailed [JGP13]. Tailoring [LZ12]. Taint [WZX^{+}13]. Tandem [WDS10]. Target [MK18]. Targeted [HA16]. Targets [PKT18]. Task [BTEGF12, EB10, LED13]. Tasks [PWSG19, RDT^{+}10, RHB^{+}12, TRD15]. Taxonomies [Ral19]. Taxonomy [AHG13, ADG16, SBGM17]. Teaching [SMG11]. Team [LED13, RKH12]. Teams [HNM13, KZDB18]. Technical [CEM11, KSD11, dSMST17, RK14, RK19, YXS^{+}19]. Technique [CPR16, GW19, LM11, MIA^{+}16, WDS10]. Techniques [AIAB13, BRDG16, BVD15, DVMB12, DNT14, FCC13, HJG11, LMZP19, LCG^{+}18, MPS12, RBK^{+}13, SBGM17, TMHM17, DJ11, Her17]. Technologies [MMPT10]. Templates [ASBZ15, HMC19, HCBB12]. Templating [MBH^{+}18]. Temporal [LYJC11, MVDS20, PMR11]. Tenant [WHYY18]. Terms [ASBZ17]. Tessellations [STM13]. Test [ABHPW10, AIAB13, AM13, ABL14, AGR19, Arc12, AKD^{+}10, ANVZ14, BH13, BSM11, CPTT12, DKM^{+}12, DBM19, DMTR10, ET16, FA13, FET^{+}17, GCZM16, Gar10, GSWH15b, GRH17, GKZ^{+}14, GZS^{+}15, eZHM13, HZZ^{+}16, HPP^{+}14, Hie15, Lat14, LO17, LRCL14, LMZP19, MIA^{+}16, MSS19, MNBB19, MHL^{+}12, MHZ^{+}12, MAT12, NM14, PZTM13, POPD15, PKT18, RM13, RKH12, SM16, TdlRSCB16, UKB10, WNM12, Yil13, YT18, YHW20, YM10]. Test-Case [HZZ^{+}16]. Test-Driven [FET^{+}17, Lat14, RM13, WNM12]. Test-First [FET^{+}17]. Test-Last [FET^{+}17]. Testability [RLR14]. Tester [IML13]. Testing [AML11, AB12, AIB12, Arc12, BH13, BDP16, BBC^{+}10, BHM^{+}15, BGP^{+}19, BB15, BMP18, BP16a, CL16, CRV13, CTZ11, CT13, CPR16, DNT14, FPTO11, HM10b, HSH^{+}12, Hie12, Hie15, HT16, Hie17, IML13, JH11, JH15, KKK11, LO17, LKTC14, LPZ12, LYC14, MvDR12, MMP15, MHS^{+}13, MBLD11, OT12, PZTM13, PCHY15, PRT10, PC10, RLR14, SBM13, SFSRC16, SPTRC18, STM13, SGM13, SCP14, TZM^{+}17, WNP^{+}20, YJ13, Yil13, YDCP14, YCM11, ZZH^{+}19, ZXC16]. Tests [BNR^{+}10, FZ12, JPAB19, RK19]. Text [PTYN19, SWHJ14]. Texts [DCMS14]. Textual [PPZ^{+}18]. Their [BLZ^{+}19, BLZ^{+}20, CMCS12, CH13, CCS^{+}13, SER^{+}10]. Theoretic [HPSW16]. Theoretical [AIB12, Arc12, HM10b, SBWM18, SYB18]. Theories [Ral19]. Theory [LBB^{+}13, RLR14, SNAH15]. Thread [Kim14]. Threaded [PCPS15]. Threshold [LLNL16]. Throughput [KGDS13]. Tier [CMCS12]. Tight [GRPF13]. Time [AM13, AGL^{+}15, BTEGF12, CRV11, CPV13, CRV13, DMTR10, EB10, FTG16, HMMB18, JLR15, Kim14, LMST10, LS13, LKB18, LB10, MK18, PRT10, TFB14, WDZ^{+}14, KSA^{+}13]. Timed [LSL^{+}18, LAL^{+}14, RNP13, VDC17, WSW^{+}15, WSWQ17]. Timeliness [Dwy17]. Timely [BVD15]. Tolerance [WHYY18]. Tolerating [ZWL^{+}16]. Tool [ABD^{+}16, CMM^{+}19, CBCB12, Gar10, SJMH^{+}19]. Tool-Based [CBCB12]. Tool-Supported [ABD^{+}16]. Tools [AHG13, CD12, GHL^{+}13, KKK11, MMPT10, MBEN15]. Top [KWK^{+}11]. Topic [BOG^{+}14, DCS^{+}18, XLD^{+}17]. Trace [CK15, CZvD11]. Traceability [AGA13, DR14, DDO11, RM17, VBJM14]. Traces [HCBB12, Hie17]. Tracing [HLP10, MCH16]. Tracking [Egy11, LKB18, RLMM15, WGC^{+}16]. Trade [GAB^{+}10]. Trade-Off [GAB^{+}10]. Tradeoff [BTS17]. Tradeoffs [DLWF19]. Transactions [Ano11q, Ano11r, Ano11z, Ano11-32, Ano12w, Ano12-32, Ano15a, Ano17a, Ano18a, Ano19a, HM10a]. Transfer [KM19]. Transformation [BDR^{+}20, FTK^{+}17, LKR14, MMPT10]. Transformations [BTWV15, CGdL14, CGdL17, VBJM14, WZ11]. Transition [CCS^{+}13, SBUK13]. Transplantation [PHLW18]. Trees [SPF14]. Trends [BMP18, CC15, KSD^{+}13]. Triaging [XLD^{+}17]. Triggered [SBUK13]. Trivial [KPJ^{+}18]. Trustrace [AGA13]. Trustworthiness [LB10]. Trustworthy [CWG^{+}18]. TSE [Ano10v, Ano10-27, Ano10y, Ano10z, Ano10x, Ano10w, Ano11-27, Ano11-31, Ano11-30, Ano11-29, Ano11-28, Ano12n, Ano12x, Ano12-31, Nus10b]. Tuning [PMR11]. Turnaround [MLS^{+}14]. Tutorial [BXXL19, PBM^{+}19]. Two [DFG16, HC18, HCBB12, KTKZ13, LR12]. Two-Channel [LR12]. Two-Component [DFG16]. Two-Phase [HC18, KTKZ13]. Typed [PG13].

Ugly [FRA^{+}19]. UIO [HT16]. UML [BJGB10, CSF^{+}16, CGC19, DSP11, EALK^{+}15, GAB^{+}10, LM11, RDT^{+}10, SBL12, SZ10]. UML-Based [BJGB10]. UML/MARTE [SBL12]. UML/OCL [CGC19]. Unauthorized [VW12]. Uncovering [SBL12]. Understanding [BBA^{+}15, LLL^{+}18, SWEK15, SGH10, MSSF13]. Undetermined [NM14]. Unequals [CT13]. Uni [VCMG20]. Uni-REPM [VCMG20]. UniDoSA [PMG19]. Unified [HBC13, PMG19, TLB14]. Uniform [SBM13]. Unit [AML11, FZ12, KKK11, Lat14, PZTM13]. Unsafe [KS12]. Unserializable [LPZ12]. Unstable [XXC^{+}15]. Update [GITT17, HSH^{+}12]. Updates [BGMM17]. Updating [CYH^{+}11]. Upstream [SPG10]. Usability [CMSSS13, YGV17]. Usage [BHKZ13, DSSP17, HJ14, HCBB12, LLL^{+}18]. Usages [ZM19]. Use [HMMB18, LBFW17, NDMX16, PMM18, SBH14, SHB18, TMHM16, ZHMZ17]. Useful [YZL^{+}15]. User [ABY16, BLVBC^{+}15, HMC19, LBK^{+}15, MM13, MBCC^{+}20, NDMX16, Thi15]. Users [GKZ^{+}14]. Using [AKRU13, ACDO12, AB12, ASBZ15, AKD^{+}10, AGL^{+}15, BRDG16, BBA^{+}15, BWRR17, CYH^{+}11, CSJ^{+}16, DR14, DCS^{+}18, DVB13, DCMS14, ET16, GRPF13, eZHM13, HSH^{+}12, HPP^{+}14, HBS^{+}17, JJW^{+}15, KGDS13, KPJ^{+}16, KKR10, LF12, LLC^{+}16, dSMST17, MPS12, MLS^{+}14, MAT12, MDB13, NZ16, PHLW18, PSB^{+}17, PC10, QSM16, RNP13, SPF14, SZ10, VDC17, WWY^{+}18, YM10, ZNKZ18]. Utility [CGMM16, PMGZ13]. Utility-Aware [CGMM16]. Utilizing [WGC^{+}16].

V [CRV11]. V-Model [CRV11]. Validating [MUA13]. Validation [ABD^{+}16, BSD14, CFMS15, LMP14, TMHM17, dCBGU12, Her17]. Validity [SMJ19]. Value [KMK12, LHWZ17, SYX14]. Value-Flow [SYX14]. Values [JJW^{+}15]. Variability [BSL^{+}13, CCS^{+}13, GWT^{+}14, IRBW16, KDO14]. Variability-Intensive [CCS^{+}13]. Variable [MHL^{+}12]. Variables [KPJ^{+}16]. Variant [NSC^{+}12]. Verification [CPV13, CGC19, DCMS14, FTG16, GRPF13, GAB^{+}10, HMC19, HJG11, LSL^{+}18, LAL^{+}14, LCNM12, MVDS20, PC10, VDC17, ZCW15, ZKZ10]. Verifying [BGM11, Cha10, CCS^{+}13, KA11, LCL^{+}13, MDB13, SHCX16]. Version [BGMM17, LZRC18, PBP^{+}15, SSP18]. versus [MBK^{+}13, MBC^{+}13]. Very [WGC^{+}16]. via [BDR^{+}20, BOG^{+}14, CL16, CK15, DDO11, FCC13, FTG16, HLP10, KNGW13, Kim14, KPJ^{+}18, LCL^{+}13, MBEN15, SWHJ14, TWL13, WWY^{+}17, WYZ^{+}18, WZ11, XCZ^{+}19, YJ13, YYG^{+}18]. Video [Ano12v, BXXL19, PBM^{+}19]. View [LKKS15]. Viewpoints [SGH10]. Violation [LPZ12, SHCX16]. Violations [BG18, BLZ^{+}19, BLZ^{+}20, LYJC11]. Virtual [CGMD19]. Visibility [Dwy17]. Visual [DSSP17, GHL^{+}13]. Visualization [CZvD11, WGC^{+}16]. Visualizations [WGC^{+}16]. Vol [Ano15a, Ano17a, Ano18a, Ano19a]. Voronoi [STM13]. Voting [BBC^{+}10]. VT [BXXL19]. VT-Revolution [BXXL19]. Vulnerabilities [DBM19, JPAB19, SMWO11, SJMH^{+}19]. Vulnerability [ANC^{+}10, MN14, TSBB20]. Vulnerable [SWHJ14].

Walks [ZJ12]. WAM [KRX11]. Warehousing [SRS12]. Watching [BXXL19]. Way [LMSN12]. Weakest [LLC^{+}16]. Web [JPAB19, SPTRC18, TSBB20, AKD^{+}10, ADTP12, JGP13, MSCC13, MvDR12, MMP15, Pre11, RDT^{+}10, SRZ15, SE13]. Weight [WLX^{+}16]. Weighted [WYZ^{+}18, KRX11]. Where [KTKZ13, LYJC11, NBKC15]. Whether [TPB^{+}17]. Which [KWK^{+}11, LR12, LP13b]. Whiteboard [MLPvdH15]. Whitening [YJ13]. Who [ZM15]. Whole [FA13, HYH^{+}17]. Whole-System [HYH^{+}17]. Whom [dSR11]. Whose [dSR11]. Will [ZM15]. Wireless [MPO^{+}19]. Wise [HPP^{+}14]. Within [JWDX17, PMR11]. Within-Project [JWDX17]. without [Hie17]. Work [CRV11, MBM^{+}17, TS12]. Workflow [SRZ15]. Workflows [LYJC11]. Workload [CKKR12]. Workloads [JGP13]. World [BBC^{+}10, YWHH19]. Worse [LP13a]. Writers [Nus10d].

XML [JPAB19].

yourself [Ano11j, Ano11i].

Zebu [BRLM11]. Zenoness [WSWQ17, WSW^{+}15].

References

Amatriain:2011:FGD

[AA11] X. Amatriain and P. Arumi. Frameworks generate domainspecific languages: A case study in the multimedia domain. IEEE Transactions on Software Engineering, 37(4):544–558, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                                      http://ieeexplore.ieee.
                                                      org/stamp/stamp.jsp?arnumber=
                                                      5441292.

AlDallal:2018:EEI

[AA18] J. Al Dallal and A. Abdin. Empirical evaluation of the impact of object-oriented code refactoring on quality attributes: A systematic literature review. IEEE Transactions on Software Engineering, 44(1):44–69, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                                      http://ieeexplore.ieee.
                                                      org/stamp/stamp.jsp?arnumber=
                                                      7833023.

Arcuri:2012:FAP

[AB12] A. Arcuri and L. Briand. Formal analysis of the probability of interaction fault detection using random testing. IEEE Transactions on Software Engineering, 38(5):1088–1099, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5999671.

Allamanis:2018:MSL

[ABB^{+}18] Miltiadis Allamanis, Earl T.

Barr, Christian Bird, Premkumar Devanbu, Mark Marron, and Charles Sutton. Mining semantic loop idioms. IEEE Transactions on Software Engineering, 44(7):651–668, July 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8355713/.

Autili:2016:TSM

[ABD^{+}16] M. Autili, A. Bertolino, G. De Angelis, D. D. Ruscio, and A. D. Sandro. A tool-supported methodology for validation and refinement of early-stage domain models. IEEE Transactions on Software Engineering, 42(1):2–25, January 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7132782.

Aleti:2013:SAO

[ABG^{+}13] A. Aleti, B. Buhnova, L. Grunske, A. Koziolek, and I. Meedeniya. Software architecture optimization methods: A systematic literature review. IEEE Transactions on Software Engineering, 39(5):658–683, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6311410.

Ali:2010:SRA

[ABHPW10] S. Ali, L. C. Briand, H. Hemmati, and R. K. Panesar-Walawege. A systematic review of the application and empirical investigation of search-based test case generation. IEEE Transactions on Software Engineering, 36(6):742–762, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5210118.

Araujo:2014:ECT

[ABL14] W. Araujo, L. C. Briand, and Y. Labiche. On the effectiveness of contracts as test oracles in the detection and diagnosis of functional faults in concurrent object-oriented software. IEEE Transactions on Software Engineering, 40(10):971–992, October 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6857355.

Al-Baik:2019:IDK

[ABM19] O. Al-Baik and J. Miller.

Integrative Double Kaizen Loop (IDKL): Towards a culture of continuous learning and sustainable improvements for software organizations. IEEE Transactions on Software Engineering, 45(12): 1189–1210, December 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Andrikopoulos:2012:ES

[ABP12] V. Andrikopoulos, S. Benbernou, and M. P. Papazoglou. On the evolution of services. IEEE Transactions on Software Engineering, 38(3):609–628, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5728828.

Akiki:2016:EAM

[ABY16] P. A. Akiki, A. K. Bandara, and Y. Yu. Engineering adaptive model-driven user interfaces. IEEE Transactions on Software Engineering, 42 (12):1118–1147, December 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7451279.

Ampatzoglou:2015:EGD

[ACCA15] A. Ampatzoglou, A. Chatzigeorgiou, S. Charalampidou, and P. Avgeriou. The effect of GoF design patterns on stability: A case study. IEEE Transactions on Software Engineering, 41(8):781–802, August 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7066925.

Altintas:2012:MPL

[ACDO12] N. I. Altintas, S. Cetin, A. H. Dogru, and H. Oguztuzun. Modeling product line software assets using domain-specific kits. IEEE Transactions on Software Engineering, 38(6):1376–1402, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6065739.

Androutsopoulos:2013:ASE

[ACH^{+}13] K. Androutsopoulos, D. Clark, M. Harman, R. M. Hierons, Z. Li, and L. Tratt. Amorphous slicing of extended finite state machines. IEEE Transactions on Software Engineering, 39(7):892–909, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6374192.

Atachiants:2016:PPP

[ADG16] R. Atachiants, G. Doherty, and D. Gregg. Parallel performance problems on shared-memory multicore systems: Taxonomy and observation. IEEE Transactions on Software Engineering, 42(8):764–785, August 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7386691.

Aranda:2016:EDK

[ADJ16] A. M. Aranda, O. Dieste, and N. Juristo. Effect of domain knowledge on elicitation effectiveness: An internally replicated controlled experiment. IEEE Transactions on Software Engineering, 42(5):427–451, May 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7307191.

Artzi:2012:FLD

[ADTP12] S. Artzi, J. Dolby, F. Tip, and M. Pistoia. Fault localization for dynamic Web applications. IEEE Transactions on Software Engineering, 38(2):314–335, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5975173.

Arnaoudova:2014:RAN

[AEP^{+}14] V. Arnaoudova, L. M. Eshkevari, M. D. Penta, R. Oliveto, G. Antoniol, and Y. G. Guéhéneuc. REPENT: Analyzing the nature of identifier renamings. IEEE Transactions on Software Engineering, 40(5):502–532, May 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6776542.

Ali:2013:TMS

[AGA13] N. Ali, Y. G. Guéhéneuc, and G. Antoniol. Trustrace: Mining software repositories to improve the accuracy of requirement traceability links. IEEE Transactions on Software Engineering, 39(5):725–741, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6341764.

Abrahao:2013:AES

[AGI^{+}13] S. Abrahão, C. Gravino, E. Insfran, G. Scanniello, and G. Tortora. Assessing the effectiveness of sequence diagrams in the comprehension of functional requirements: Results from a family of five experiments. IEEE Transactions on Software Engineering, 39(3):327–342, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6193111.

Autili:2015:AQR

[AGL^{+}15] M. Autili, L. Grunske, M. Lumpe, P. Pelliccione, and A. Tang. Aligning qualitative, real-time, and probabilistic property specification patterns using a structured English grammar. IEEE Transactions on Software Engineering, 41(7):620–638, July 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7029714.

Arcaini:2019:DBA

[AGR19] P. Arcaini, A. Gargantini, and E. Riccobene. Decomposition-based approach for model-based test generation. IEEE Transactions on Software Engineering, 45(5): 507–520, May 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Ali:2013:TMC

[AHG13] N. M. Ali, J. Hosking, and J. Grundy. A taxonomy and mapping of computer-based critiquing tools. IEEE Transactions on Software Engineering, 39 (11):1494–1520, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6570472.

Atlee:2012:GEI

[AI12] J. M. Atlee and P. Inverardi. Guest Editor’s introduction: International Conference on Software Engineering. IEEE Transactions on Software Engineering, 38 (1):3–4, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6141071.

Ali:2013:GTD

[AIAB13] S. Ali, M. Zohaib Iqbal, A. Arcuri, and L. C. Briand. Generating test data from OCL constraints with search techniques. IEEE Transactions on Software Engineering, 39(10):1376–1402, October 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6491405.

Arcuri:2012:RTT

[AIB12] A. Arcuri, M. Z. Iqbal, and L. Briand. Random testing: Theoretical results and practical implications. IEEE Transactions on Software Engineering, 38(2):258–277, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6104067.

Artzi:2010:FBW

[AKD^{+}10] S. Artzi, A. Kiezun, J. Dolby, F. Tip, D. Dig, A. Paradkar, and M. D. Ernst. Finding bugs in Web applications using dynamic test generation and explicit-state model checking. IEEE Transactions on Software Engineering, 36(4):474–494, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5416728.

Apel:2013:LIA

[AKL13] S. Apel, C. Kästner, and C. Lengauer. Languageindependent and automated software composition: The FeatureHouse experience. IEEE Transactions on Software Engineering, 39(1):63– 79, January 2013. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6095570.

Alrajeh:2013:ERU

[AKRU13] D. Alrajeh, J. Kramer, A. Russo, and S. Uchitel.

Elaborating requirements using model checking and inductive learning. IEEE Transactions on Software Engineering, 39(3):361–383, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6216384.

AlDallal:2011:MDP

[Al 11]  J. Al Dallal.  Measur-

ing the discriminative power of object-oriented class cohesion metrics. IEEE Transactions on Software Engineering, 37(6):788–804, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5639020.

Aljazzar:2010:DES

[AL10] H. Aljazzar and S. Leue. Directed explicit state-space search in the generation of counterexamples for stochastic model checking. IEEE Transactions on Software Engineering, 36(1):37–60, January/February 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5262946.

Andrade:2013:GTC

[AM13] W. L. Andrade and P. D. L.

Machado. Generating test cases for real-time systems based on symbolic models. IEEE Transactions on Software Engineering, 39(9):1216–1229, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6475130.

Andrews:2011:GAR

[AML11] J. H. Andrews, T. Menzies, and F. C. H. Li. Genetic algorithms for randomized unit testing. IEEE Transactions on Software Engineering, 37 (1):80–94, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5704237.

Antunes:2010:VDA

[ANC^{+}10] J. Antunes, N. Neves, M. Correia, P. Verissimo, and R. Neves. Vulnerability discovery with attack injection. IEEE Transactions on Software Engineering, 36(3):357–370, May/ June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5374427.

Amann:2019:SES

[ANN^{+}19] S. Amann, H. A. Nguyen, S. Nadi, T. N. Nguyen, and M. Mezini. A systematic evaluation of static API-misuse detectors. IEEE Transactions on Software Engineering, 45 (12):1170–1188, December 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Anonymous:2010:RL

[Ano10a] Anonymous. 2009 reviewers list. IEEE Transactions on Software Engineering, 36(1): 141–143, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5401366.

Anonymous:2010:GRJ

[Ano10b] Anonymous. 7 great reasons for joining the IEEE Computer Society [advertisement]. IEEE Transactions on Software Engineering, 36 (1):144, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5401367.

Anonymous:2010:BCa

[Ano10c] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 36(1): c4, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5401369.

Anonymous:2010:BCc

[Ano10d] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 36(2):c4, March/April 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5439571.

Anonymous:2010:BCf

[Ano10e] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 36(3): c4, May/June 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5473902.

Anonymous:2010:BCb

[Ano10f] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 36(4):c4, July/August 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5532340.

Anonymous:2010:BCe

[Ano10g] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 36(5): c4, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5593047.

Anonymous:2010:BCd

[Ano10h] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 36(6):c4, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5644738.

Anonymous:2010:FCc

[Ano10i] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 36(1): c1, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5401362.

Anonymous:2010:FCf

[Ano10j] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 36(2):c1, March/April 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5439566.

Anonymous:2010:FCa

[Ano10k] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 36(3): c1, May/June 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5473897.

Anonymous:2010:FCd

[Ano10l] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 36(4):c1, July/August 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5532335.

Anonymous:2010:FCb

[Ano10m] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 36(5): c1, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5593042.

Anonymous:2010:FCe

[Ano10n] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 36(6):c1, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5644732.

Anonymous:2010:ICS

[Ano10o] Anonymous. IEEE Computer Society CSDP [advertisement]. IEEE Transactions on Software Engineering, 36 (6):878, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5644736.

Anonymous:2010:IFCd

[Ano10p] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 36(1):c2, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5401363.

Anonymous:2010:IFCa

[Ano10q] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 36(2):c2, March/April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5439567.

Anonymous:2010:IFCb

[Ano10r] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 36(3):c2, May/June 2010. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5473898.

Anonymous:2010:IFCe

[Ano10s] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 36(4):c2, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5532336.

Anonymous:2010:IFCc

[Ano10t] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 36 (5):c2, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5593043.

Anonymous:2010:IFCf

[Ano10u] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 36 (6):c2, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5644733.

Anonymous:2010:TIAa

[Ano10v]  Anonymous.  TSE infor-

mation for authors. IEEE Transactions on Software Engineering, 36(1):c3, January/February 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5401368.

Anonymous:2010:TIAf

[Ano10w] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 36(2):c3, March/April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5439570.

Anonymous:2010:TIAe

[Ano10x] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 36(3):c3, May/June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5473901.

Anonymous:2010:TIAc

[Ano10y] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 36(4):c3, July/August 2010. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5532339.

Anonymous:2010:TIAd

[Ano10z] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 36(5):c3, September/October 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5593046.

Anonymous:2010:TIAb

[Ano10-27]  Anonymous.  TSE infor-

mation for authors. IEEE Transactions on Software Engineering, 36(6):c3, November/December 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5644737.

Anonymous:2011:AI

[Ano11a] Anonymous. 2010 annual index. IEEE Transactions on Software Engineering, 37(1):Online Only, January/February 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5704239.

Anonymous:2011:RL

[Ano11b] Anonymous. 2010 reviewers list. IEEE Transactions on Software Engineering, 37(1): 142–144, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5704238.

Anonymous:2011:BCd

[Ano11c] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 37(1): c4, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5704241.

Anonymous:2011:BCc

[Ano11d] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 37(2):c4, March/April 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5739159.

Anonymous:2011:BCb

[Ano11e] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 37(3): c4, May/June 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5779019.

Anonymous:2011:BCf

[Ano11f] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 37(4):c4, July/August 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5966996.

Anonymous:2011:BCe

[Ano11g] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 37(5): c4, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6030118.

Anonymous:2011:BCa

[Ano11h] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 37(6):c4, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6095286.

Anonymous:2011:DYCb

[Ano11i]  Anonymous.  Distinguish

yourself with the CSDP [advertisement]. IEEE Transactions on Software Engineering, 37(3):448, May/June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5779017.

Anonymous:2011:DYCa

[Ano11j]  Anonymous.  Distinguish

yourself with the CSDP [advertisement]. IEEE Transactions on Software Engineering, 37(4):592, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5966994.

Anonymous:2011:FCc

[Ano11k] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 37(1): c1, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5704234.

Anonymous:2011:FCf

[Ano11l] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 37(2):c1, March/April 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5739155.

Anonymous:2011:FCb

[Ano11m] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 37(3): c1, May/June 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5779014.

Anonymous:2011:FCa

[Ano11n] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 37(4):c1, July/August 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5966992.

Anonymous:2011:FCe

[Ano11o] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 37(5): c1, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6030114.

Anonymous:2011:FCd

[Ano11p] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 37(6):c1, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6095280.

Anonymous:2011:ICSa

[Ano11q] Anonymous. IEEE Computer Society magazines and transactions available in ePUB format [advertisement]. IEEE Transactions on Software Engineering, 37 (5):736, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6030116.

Anonymous:2011:ICSb

[Ano11r] Anonymous. IEEE Computer Society magazines and transactions available in ePUB format [advertisement]. IEEE Transactions on Software Engineering, 37 (6):878, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6095282.

Anonymous:2011:IBC

[Ano11s] Anonymous. [inside back cover]. IEEE Transactions on Software Engineering, 37 (6):c3, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
6095285.

Anonymous:2011:IFCf

[Ano11t] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 37(1):c2, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5704235.

Anonymous:2011:IFCd

[Ano11u] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 37(2):c2, March/April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5739156.

Anonymous:2011:IFCa

[Ano11v] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 37(3):c2, May/June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5779015.

Anonymous:2011:IFCe

[Ano11w] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 37(4):c2, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5966993.

Anonymous:2011:IFCc

[Ano11x] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 37 (5):c2, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6030115.

Anonymous:2011:IFCb

[Ano11y] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 37 (6):c2, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6095281.

Anonymous:2011:NTI

[Ano11z] Anonymous. New transactions issue alerts. IEEE Transactions on Software Engineering, 37(6):879, November/December 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6095283.

Anonymous:2011:TIAa

[Ano11-27]  Anonymous.  TSE infor-

mation for authors. IEEE Transactions on Software Engineering, 37(1):c3, January/February 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5704240.

Anonymous:2011:TIAe

[Ano11-28] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 37(2):c3, March/April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5739158.

Anonymous:2011:TIAd

[Ano11-29] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 37(3):c3, May/June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5779018.

Anonymous:2011:TIAc

[Ano11-30] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 37(4):c3, July/August 2011. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5966995.

Anonymous:2011:TIAb

[Ano11-31] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 37(5):c3, September/October 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6030117.

Anonymous:2011:WNT

[Ano11-32] Anonymous. What’s new in transactions [advertisement]. IEEE Transactions on Software Engineering, 37 (6):880, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6095284.

Anonymous:2012:AI

[Ano12a]  Anonymous.  2011 an-

nual index. IEEE Transactions on Software Engineering, 38(1):Online Only– ??, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141073.

Anonymous:2012:RL

[Ano12b] Anonymous. 2011 reviewers list. IEEE Transactions on Software Engineering, 38(1): 236–238, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141072.

Anonymous:2012:BCd

[Ano12c] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 38(1): c4, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141077.

Anonymous:2012:BCe

[Ano12d] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 38(2):c4, March/April 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6173079.

Anonymous:2012:BCa

[Ano12e] Anonymous. [back cover].

IEEE Transactions on Software Engineering, 38(3): c4, May/June 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6205672.

Anonymous:2012:BCc

[Ano12f] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 38(4):c4, July/August 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6249694.

Anonymous:2012:BCb

[Ano12g] Anonymous. [back cover]. IEEE Transactions on Software Engineering, 38(6):c4, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6363460.

Anonymous:2012:BIC

[Ano12h] Anonymous. [back inside cover]. IEEE Transactions on Software Engineering, 38 (6):c3, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6363459.

Anonymous:2012:Ca

[Ano12i] Anonymous. [cover2]. IEEE Transactions on Software Engineering, 38(2):c2, March/ April 2012. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6173076.

Anonymous:2012:Cb

[Ano12j] Anonymous. [cover3]. IEEE Transactions on Software Engineering, 38(1):c3, January/February 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6141076.

Anonymous:2012:FICa

[Ano12k] Anonymous. [front inside cover]. IEEE Transactions on Software Engineering, 38(1):c2, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141069.

Anonymous:2012:FICb

[Ano12l] Anonymous. [front inside cover]. IEEE Transactions on Software Engineering, 38 (6):c2, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6363458.

Anonymous:2012:ICSa

[Ano12m] Anonymous. IEEE Computer Society [back cover].

IEEE Transactions on Software Engineering, 38(5): c4, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311394.

Anonymous:2012:ICSb

[Ano12n] Anonymous. IEEE Computer Society OnlinePlus coming soon to TSE. IEEE Transactions on Software Engineering, 38(3):736, May/June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6205670.

Anonymous:2012:IOA

[Ano12o]  Anonymous.  IEEE open

access publishing [advertisement]. IEEE Transactions on Software Engineering, 38 (5):1232, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311396.

Anonymous:2012:IBCc

[Ano12p] Anonymous. [inside back cover]. IEEE Transactions on Software Engineering, 38(2):c3, March/April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6173078.

Anonymous:2012:IBCa

[Ano12q] Anonymous. [inside back cover]. IEEE Transactions on Software Engineering, 38(3):c3, May/June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6205671.

Anonymous:2012:IBCb

[Ano12r] Anonymous. [inside back cover]. IEEE Transactions on Software Engineering, 38(4):c3, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6249693.

Anonymous:2012:IFCb

[Ano12s] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 38(3):c2, May/June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6205669.

Anonymous:2012:IFCc

[Ano12t] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 38(4):c2, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6249692.

Anonymous:2012:IFCa

[Ano12u] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 38 (5):c2, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311392.

Anonymous:2012:IOV

[Ano12v] Anonymous. Introduction to OnlinePlus video [advertisement]. IEEE Transactions on Software Engineering, 38 (1):239, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141074.

Anonymous:2012:NTN

[Ano12w]  Anonymous.  New trans-

actions newsletter [advertisement]. IEEE Transactions on Software Engineering, 38(4):992, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
6249695.

Anonymous:2012:OCS

[Ano12x] Anonymous. OnlinePlus coming soon to TSE [advertisement]. IEEE Transactions on Software Engineering, 38(6): 1488, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6363463.

Anonymous:2012:TCFf

[Ano12y]  Anonymous.  [table of

contents — front cover]. IEEE Transactions on Software Engineering, 38(6):c1, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6363457.

Anonymous:2012:TCFd

[Ano12z] Anonymous. Table of contents [front cover]. IEEE Transactions on Software Engineering, 38(1):c1, January/February 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6141068.

Anonymous:2012:TCFa

[Ano12-27] Anonymous. Table of contents [front cover]. IEEE Transactions on Software Engineering, 38(2):c1, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6173075.

Anonymous:2012:TCFe

[Ano12-28] Anonymous. Table of contents [front cover]. IEEE Transactions on Software Engineering, 38(3):c1, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6205668.

Anonymous:2012:TCFc

[Ano12-29] Anonymous. Table of contents [front cover]. IEEE Transactions on Software Engineering, 38(4):c1, July/ August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6249798.

Anonymous:2012:TCFb

[Ano12-30] Anonymous. Table of contents [front cover]. IEEE Transactions on Software Engineering, 38(5):c1, September/October 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6311484.

Anonymous:2012:TIA

[Ano12-31] Anonymous. TSE information for authors [inside back cover]. IEEE Transactions on Software Engineering, 38 (5):c3, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311393.

Anonymous:2012:WNT

[Ano12-32] Anonymous. What’s new in transactions [advertisement]. IEEE Transactions on Software Engineering, 38 (1):240, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141075.

Anonymous:2013:AI

[Ano13a] Anonymous. 2012 annual index. IEEE Transactions on Software Engineering, 39 (1):Not in print, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6397558.

Anonymous:2013:RL

[Ano13b] Anonymous. 2012 reviewers list. IEEE Transactions on Software Engineering, 39(1):141–144, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6397559.

Anonymous:2014:RL

[Ano14] Anonymous. 2013 reviewers list. IEEE Transactions on Software Engineering, 40(1):103–106, January 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6755631.

Anonymous:2015:IIT

[Ano15a]  Anonymous.  2014 in-

dex IEEE Transactions on Software Engineering vol.

40. IEEE Transactions on Software Engineering, 41 (1):104–112, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7014339.

Anonymous:2015:RL

[Ano15b] Anonymous. 2014 reviewers list*. IEEE Transactions on Software Engineering, 41(1):100–103, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7004099.

Anonymous:2016:RL

[Ano16] Anonymous. 2015 reviewers list. IEEE Transactions on Software Engineering, 42(1):100–102, January 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7374786.

Anonymous:2017:IIT

[Ano17a]  Anonymous.  2016 in-

dex IEEE Transactions on Software Engineering vol.

42. IEEE Transactions on Software Engineering, 43(1):1–8, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7809252.

Anonymous:2017:RL

[Ano17b] Anonymous. 2016 reviewers list*. IEEE Transactions on Software Engineering, 43(1):100–103, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7809238.

Anonymous:2018:IIT

[Ano18a]  Anonymous.  2017 in-

dex IEEE Transactions on Software Engineering vol.

43. IEEE Transactions on Software Engineering, 44(1):1–9, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        8249567.

Anonymous:2018:RL

[Ano18b] Anonymous. 2017 reviewers list*. IEEE Transactions on Software Engineering, 44(1):100–102, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        8249563.

Anonymous:2019:IIT

[Ano19a] Anonymous. 2018 index IEEE Transactions on Software Engineering vol. 44. IEEE Transactions on Software Engineering, 45(1):1–9, January 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Anonymous:2019:RL

[Ano19b] Anonymous. 2018 reviewers list. IEEE Transactions on Software Engineering, 45(1): 107–110, January 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Athanasiou:2014:TCQ

[ANVZ14] D. Athanasiou, A. Nugroho, J. Visser, and A. Zaidman.

Test code quality and its relation to issue handling performance. IEEE Transactions on Software Engineering, 40 (11):1100–1125, November 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6862882.

Arcuri:2012:TEA

[Arc12] A. Arcuri. A theoretical and empirical analysis of the role of test sequence length in software testing for structural coverage. IEEE Transactions on Software Engineering, 38(3):497–519, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5750005.

Armaly:2018:CPC

[ARM18] Ameer Armaly, Paige Rodeghero, and Collin McMillan. A comparison of program comprehension strategies by blind and sighted programmers. IEEE Transactions on Software Engineering, 44 (8):712–724, August 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7987041/.

Arora:2015:ACC

[ASBZ15] C. Arora, M. Sabetzadeh, L. Briand, and F. Zimmer. Automated checking of conformance to requirements templates using natural language processing. IEEE Transactions on Software Engineering, 41(10):944–968, October 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7100933.

Arora:2017:AEC

[ASBZ17] C. Arora, M. Sabetzadeh, L. Briand, and F. Zimmer. Automated extraction and clustering of requirements glossary terms. IEEE Transactions on Software Engineering, 43(10):918–945, October 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7765062.

Alves:2018:RIS

[ASM^{+}18] Everton L. G. Alves, Myoungkyu Song, Tiago Massoni, Patrícia D. L. Machado, and Miryung Kim. Refactoring inspection support for manual refactoring edits. IEEE Transactions on Software Engineering, 44(4):365–383, April 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
7874212.

Belli:2015:EMM

[BB15] F. Belli and M. Beyazıt.

Exploiting model morphology for event-based testing. IEEE Transactions on Software Engineering, 41 (2):113–134, February 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6915728.

Beschastnikh:2015:UDS

[BBA^{+}15] I. Beschastnikh, Y. Brun, J. Abrahamson, M. D. Ernst, and A. Krishnamurthy. Using declarative specification to improve the understanding, extensibility, and comparison of model-inference algorithms. IEEE Transactions on Software Engineering, 41(4):408–428, April 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6951474.

Balzarotti:2010:ETS

[BBC^{+}10] D. Balzarotti, G. Banks, M. Cova, V. Felmetsger, R. Kemmerer, W. Robertson, F. Valeur, and G. Vigna. An experience in testing the security of realworld electronic voting systems. IEEE Transactions on Software Engineering, 36(4):453–473, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5210119.

Bate:2017:EBP

[BBD17] I. Bate, A. Burns, and R. I. Davis. An enhanced bailout protocol for mixed criticality embedded software. IEEE Transactions on Software Engineering, 43(4):298–320, April 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7516652.

Bossi:2017:SPM

[BBH17] L. Bossi, E. Bertino, and S. R. Hussain. A system for profiling and monitoring database access patterns by application programs for anomaly detection. IEEE Transactions on Software Engineering, 43(5):415–431, May 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7534833.

Bowman:2010:SCR

[BBL10] M. Bowman, L. C. Briand, and Y. Labiche. Solving the class responsibility assignment problem in objectoriented analysis with multiobjective genetic algorithms. IEEE Transactions on Software Engineering, 36(6):817– 837, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5530324.

Bishop:2011:TFC

[BBL^{+}11] P. Bishop, R. Bloomfield, B. Littlewood, A. Povyakalo, and D. Wright. Toward a formalism for conservative claims about the dependability of software-based systems. IEEE Transactions on Software Engineering, 37(5): 708–717, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5492693.

Bosu:2017:PAS

[BCB^{+}17] A. Bosu, J. C. Carver, C. Bird, J. Orbeck, and C. Chockley. Process aspects and social dynamics of contemporary code review: Insights from open source development and industrial practice at Microsoft. IEEE Transactions on Software Engineering, 43(1):56–75, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?arnumber=
                                        7484733.

Banerjee:2018:ERR

[BCBR18] Abhijeet Banerjee, Lee Kee Chong, Clément Ballabriga, and Abhik Roychoudhury. EnergyPatch: Repairing resource leaks to improve energy-efficiency of Android apps. IEEE Transactions on Software Engineering, 44(5):470–490, May 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7889026/.

Betts:2018:IEC

[BCD^{+}18] Adam Betts, Nathan Chong, Pantazis Deligiannis, Alastair F. Donaldson, and Jeroen Ketema. Implementing and evaluating candidate-based invariant generation. IEEE Transactions on Software Engineering, 44(7):631–650, July 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7955079/.

Basin:2014:MDM

[BCE^{+}14] D. Basin, M. Clavel, M. Egea, M. A. G. de Dios, and C. Dania. A model-driven methodology for developing secure data-management applications. IEEE Transactions on Software Engineering, 40(4):324–337, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6698396.

Bernardeschi:2018:PSI

[BDM18] Cinzia Bernardeschi, Andrea Domenici, and Paolo Masci. A PVS-Simulink integrated environment for model-based analysis of cyber-physical systems. IEEE Transactions on Software Engineering, 44(6):512–533, June 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7900400/.

Baluda:2016:BSA

[BDP16] M. Baluda, G. Denaro, and M. Pezzè. Bidirectional symbolic analysis for effective branch testing. IEEE Transactions on Software Engineering, 42(5):403–426, May 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7296670.

Basciani:2020:ASO

[BDR^{+}20] F. Basciani, M. D’Emidio, D. D. Ruscio, D. Frigioni, L. Iovino, and A. Pierantonio. Automated selection of optimal model transformation chains via shortest-path algorithms. IEEE Transactions on Software Engineering, 46(3):251–279, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Baresi:2011:SSB

[BG11] L. Baresi and S. Guinea. Self-supervising BPEL processes. IEEE Transactions on Software Engineering, 37(2):247–263, March/ April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5432226.

Barbosa:2018:GAR

[BG18] Eiji Adachi Barbosa and Alessandro Garcia. Globalaware recommendations for repairing violations in exception handling. IEEE Transactions on Software Engineering, 44(9):855–873, September 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7953550/.

Baresi:2011:LVP

[BGM11] L. Baresi, C. Ghezzi, and L. Mottola. Loupe: Verifying publish–subscribe architectures with a magnifying lens. IEEE Transactions on Software Engineering, 37(2):228–246, March/ April 2011. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5432228.

Baresi:2017:EDU

[BGMM17] L. Baresi, C. Ghezzi, X. Ma, and V. P. L. Manna. Efficient dynamic updates of distributed components through version consistency. IEEE Transactions on Software Engineering, 43(4):340–358, April 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7516718.

Beller:2019:DTI

[BGP^{+}19] Moritz Beller, Georgios Gousios, Annibale Panichella, Sebastian Proksch, Sven Amann, and Andy Zaidman. Developer testing in the IDE: Patterns, beliefs, and behavior. IEEE Transactions on Software Engineering, 45(3):261–284, March 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8116886/.

Barbosa:2016:EEH

[BGRJ16] E. A. Barbosa, A. Garcia, M. P. Robillard, and B. Jakobus. Enforcing exception handling policies with a domain-specific language. IEEE Transactions on Software Engineering, 42(6):559–584, June 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7348692.

Brandt:2013:EPS

[BGS^{+}13] J. Brandt, M. Gemünde, K. Schneider, S. K. Shukla, and J. P. Talpin. Embedding polychrony into synchrony. IEEE Transactions on Software Engineering, 39(7):917–929, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6381420.

Bodden:2010:AOR

[BH10] E. Bodden and K. Havelund. Aspect-oriented race detection in Java. IEEE Transactions on Software Engineering, 36(4):509–527, July/ August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5406531.

Baker:2013:EEM

[BH13] R. Baker and I. Habli. An empirical evaluation of mutation testing for improving the test quality of safetycritical software. IEEE Transactions on Software Engineering, 39(6):787–805, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6298894.

Bergel:2011:FLS

[BHCC11] A. Bergel, W. Harrison, V. Cahill, and S. Clarke. FlowTalk: Language support for long-latency operations in embedded devices. IEEE Transactions on Software Engineering, 37(4):526–543, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5492692.

Brun:2013:EDC

[BHEN13] Y. Brun, R. Holmes, M. D.

Ernst, and D. Notkin. Early detection of collaboration conflicts and risks. IEEE Transactions on Software Engineering, 39(10): 1358–1375, October 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6520859.

Basin:2013:MDU

[BHKZ13] D. Basin, M. Harvan, F. Klaedtke, and E. Zalinescu. Monitoring data usage in distributed systems. IEEE Transactions on Software Engineering, 39(10):1403–1426, October 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6493331.

Barr:2015:OPS

[BHM^{+}15] E. T. Barr, M. Harman, P. McMinn, M. Shahbaz, and S. Yoo. The oracle problem in software testing: A survey. IEEE Transactions on Software Engineering, 41(5):507–525, May 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6963470.

Bennaceur:2015:ASM

[BI15] A. Bennaceur and V. Issarny. Automated synthesis of mediators to support component interoperability. IEEE Transactions on Software Engineering, 41(3):221–240, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6936339.

Bendraou:2010:CSU

[BJGB10] R. Bendraou, J. M. Jezequel, M. P. Gervais, and X. Blanc. A comparison of six UML-based languages for software process modeling. IEEE Transactions on Software Engineering, 36(5): 662–675, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5593045.

Brosch:2012:ABR

[BKBR12] F. Brosch, H. Koziolek, B. Buhnova, and R. Reussner. Architecture-based reliability prediction with the Palladio component model. IEEE Transactions on Software Engineering, 38(6):1319–1339, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6018968.

Bollig:2010:LCA

[BKKL10] B. Bollig, J. P. Katoen, C. Kern, and M. Leucker. Learning communicating automata from MSCs. IEEE Transactions on Software Engineering, 36(3):390–408, May/June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5374425.

Bartel:2014:SAE

[BKML14] A. Bartel, J. Klein, M. Monperrus, and Y. Le Traon. Static analysis for extracting permission checks of a large scale framework: The challenges and solutions for analyzing android. IEEE Transactions on Software Engineering, 40(6):617–632, June 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6813664.

Bennin:2018:MDB

[BKP^{+}18] Kwabena Ebo Bennin, Jacky Keung, Passakorn Phannachitta, Akito Monden, and Solomon Mensah. MA-HAKIL: Diversity based oversampling approach to alleviate the class imbalance issue in software defect prediction. IEEE Transactions on Software Engineering, 44(6):534–550, June 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7990590/.

Bavota:2015:IAC

[BLVBC^{+}15] G. Bavota, M. Linares-Vásquez, C. E. Bernal-Cárdenas, M. D. Penta, R. Oliveto, and D. Poshyvanyk. The impact of API change- and fault-proneness on the user ratings of Android apps. IEEE Transactions on Software Engineering, 41(4):384–407, April 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6945855.

Bian:2019:DBD

[BLZ^{+}19] P. Bian, B. Liang, Y. Zhang, C. Yang, W. Shi, and Y. Cai. Detecting bugs by discovering expectations and their violations. IEEE Transactions on Software Engineering, 45(10):984–1001, October 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). See corrections [BLZ^{+}20].

Bian:2020:CDB

[BLZ^{+}20] P. Bian, B. Liang, Y. Zhang, C. Yang, W. Shi, and Y. Cai. Corrections to “Detecting Bugs by Discovering Expectations and Their Violations”. IEEE Transactions on Software Engineering, 46 (1):113, January 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). See [BLZ^{+}19].

Brosig:2015:QEM

[BMB^{+}15] F. Brosig, P. Meier, S. Becker, A. Koziolek, H. Koziolek, and S. Kounev. Quantitative evaluation of model-driven performance analysis and simulation of component-based architectures. IEEE Transactions on Software Engineering, 41(2):157–175, February 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6920061.

Bianchi:2018:SRT

[BMP18] Francesco Adalberto Bianchi, Alessandro Margara, and Mauro Pezzè. A survey of recent trends in testing concurrent software systems. IEEE Transactions on Software Engineering, 44(8):747–783, August 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7932530/.

Beckman:2010:PT

[BNR^{+}10] N. E. Beckman, A. V. Nori, S. K. Rajamani, R. J. Simmons, S. D. Tetali, and A. V. Thakur. Proofs from tests. IEEE Transactions on Software Engineering, 36(4):495–508, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5444886.

Bavota:2014:MRM

[BOG^{+}14] G. Bavota, R. Oliveto, M. Gethers, D. Poshyvanyk, and A. De Lucia. Methodbook: Recommending move method refactorings via relational topic models. IEEE Transactions on Software Engineering, 40(7):671–694, July 2014. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6684534.

Bohme:2016:PAE

[BP16a] M. Böhme and S. Paul. A probabilistic analysis of the efficiency of automated software testing. IEEE Transactions on Software Engineering, 42(4):345–360, April 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7289448.

Borstler:2016:RMC

[BP16b] J. Börstler and B. Paech. The role of method chains and comments in software readability and comprehension — an experiment. IEEE Transactions on Software Engineering, 42(9):886–898, September 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7404062.

Baah:2010:PPD

[BPH10] G. K. Baah, A. Podgurski, and M. J. Harrold. The probabilistic program dependence graph and its application to fault diagnosis. IEEE Transactions on Software Engineering, 36(4):528–545, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5374423.

Bruce:2019:AOS

[BPHB19] B. R. Bruce, J. Petke, M. Harman, and E. T. Barr. Approximate oracles and synergy in software energy search spaces. IEEE Transactions on Software Engineering, 45(11): 1150–1169, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Bohme:2019:CBG

[BPR19] M. Böhme, V. Pham, and A. Roychoudhury. Coveragebased greybox fuzzing as Markov chain. IEEE Transactions on Software Engineering, 45(5):489–506, May 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Berthier:2016:DAM

[BRDG16] N. Berthier, É. Rutten, N. De Palma, and S. M. K. Gueye. Designing autonomic management systems by using reactive control techniques. IEEE Transactions on Software Engineering, 42(7):640–657, July 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
7360217.

Burgy:2011:ZLB

[BRLM11] L. Burgy, L. Reveillere, J. Lawall, and G. Muller. Zebu: A language-based approach for network protocol message processing. IEEE Transactions on Software Engineering, 37(4): 575–591, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5487528.

Bergersen:2014:CVI

[BSD14] G. R. Bergersen, D. I. K.

Sjøberg, and T. Dybå. Construction and validation of an instrument for measuring programming skill. IEEE Transactions on Software Engineering, 40(12): 1163–1184, December 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6882243.

Bagheri:2015:CCA

[BSGM15] H. Bagheri, A. Sadeghi, J. Garcia, and S. Malek. COVERT: Compositional analysis of Android inter-app permission leakage. IEEE Transactions on Software Engineering, 41(9):866–886, September 2015. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7079508.

Bick:2018:CCL

[BSH^{+}18] Saskia Bick, Kai Spohrer, Rashina Hoda, Alexander Scheerer, and Armin Heinzl. Coordination challenges in large-scale software development: A case study of planning misalignment in hybrid settings. IEEE Transactions on Software Engineering, 44(10):932–950, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7990187/.

Berger:2013:SVM

[BSL^{+}13] T. Berger, S. She, R. Lotufo, A. Wasowski, and K. Czarnecki. A study of variability models and languages in the systems software domain. IEEE Transactions on Software Engineering, 39 (12):1611–1640, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6572787.

Bryce:2011:DSM

[BSM11] R. C. Bryce, S. Sampath, and A. M. Memon. Developing a single model and test prioritization strategies for event-driven software. IEEE Transactions on Software Engineering, 37 (1):48–64, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5401169.

Banerjee:2017:DBR

[BSM17] K. Banerjee, D. Sarkar, and C. Mandal. Deriving bisimulation relations from path extension based equivalence checkers. IEEE Transactions on Software Engineering, 43(10):946–953, October 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7801117.

Burdalo:2012:AEG

[BTEGF12] L. Búrdalo, A. Terrasa, A. Espinosa, and A. García-Fornes. Analyzing the effect of gain time on softtask scheduling policies in real-time systems. IEEE Transactions on Software Engineering, 38(6):1305–1318, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6025357.

Bagheri:2017:ASD

[BTS17] H. Bagheri, C. Tang, and K. Sullivan. Automated synthesis and dynamic analysis of tradeoff spaces for object-relational mapping. IEEE Transactions on Software Engineering, 43(2): 145–163, February 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7506009.

Burgueno:2015:SFL

[BTWV15] L. Burgueño, J. Troya, M. Wimmer, and A. Vallecillo. Static fault localization in model transformations. IEEE Transactions on Software Engineering, 41(5):490–506, May 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6967841.

Blincoe:2015:FCB

[BVD15] K. Blincoe, G. Valetto, and D. Damian. Facilitating coordination between software developers: A study and techniques for timely and efficient recommendations. IEEE Transactions on Software Engineering, 41 (10):969–985, October 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7105409.

Buse:2010:LMC

[BW10] R. P. L. Buse and W. R.

Weimer. Learning a metric for code readability. IEEE Transactions on Software Engineering, 36(4): 546–558, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5332232.

Brickey:2012:CSA

[BWB12] J. Brickey, S. Walczak, and T. Burgess. Comparing semi-automated clustering methods for persona development. IEEE Transactions on Software Engineering, 38(3):537–546, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5928355.

Borg:2017:SCI

[BWRR17] M. Borg, K. Wnuk, B. Regnell, and P. Runeson. Supporting change impact analysis using a recommendation system: An industrial case study in a safetycritical context. IEEE Transactions on Software Engineering, 43(7):675–700, July 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7637029.

Bao:2019:VRI

[BXXL19] L. Bao, Z. Xing, X. Xia, and D. Lo. VT-Revolution: Interactive programming video tutorial authoring and watching system. IEEE Transactions on Software Engineering, 45(8):823–838, August 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Brun:2015:SAR

[ByBEM15] Y. Brun, J. y. Bang, G. Edwards, and N. Medvidovic. Self-adapting reliability in distributed software systems. IEEE Transactions on Software Engineering, 41 (8):764–780, August 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7058381.

Bai:2018:TMC

[BYW^{+}18] Guangdong Bai, Quanqi Ye, Yongzheng Wu, Heila Botha, Jun Sun, Yang Liu, Jin Song Dong, and Willem Visser. Towards model checking Android applications. IEEE Transactions on Software Engineering, 44(6):595–612, June 2018. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7911333/.

Charalampidou:2017:IEM

[CAC^{+}17] S. Charalampidou, A. Ampatzoglou, A. Chatzigeorgiou, A. Gkortzis, and P. Avgeriou. Identifying extract method refactoring opportunities based on functional relevance. IEEE Transactions on Software Engineering, 43(10):954–974, October 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7801138.

Cai:2018:HPD

[Cai18] Haipeng Cai. Hybrid program dependence approximation for effective dynamic impact prediction. IEEE Transactions on Software Engineering, 44(4):334–364, April 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7895205/.

Chen:2017:SAO

[CB17] T. Chen and R. Bahsoon.

Self-adaptive and online QoS modeling for cloud-based software services. IEEE Transactions on Software Engineering, 43(5):453–475, May 2017. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7572219.

Cassou:2012:TTB

[CBCB12] D. Cassou, J. Bruneau, C. Consel, and E. Balland. Toward a toolbased development methodology for pervasive computing applications. IEEE Transactions on Software Engineering, 38(6):1445–1463, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6051438.

Ciolek:2017:IMA

[CBD^{+}17] D. Ciolek, V. Braberman, N. D’Ippolito, N. Piterman, and S. Uchitel. Interaction models and automated control under partial observable environments. IEEE Transactions on Software Engineering, 43(1):19–33, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7466810.

Cai:2014:MSD

[CC14] Y. Cai and W. K. Chan. Magiclock: Scalable detection of potential deadlocks in large-scale multithreaded programs. IEEE Transactions on Software Engineering, 40(3):266–281, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6718069.

Chaikalis:2015:FJS

[CC15] T. Chaikalis and A. Chatzigeorgiou. Forecasting Java software evolution trends employing network models. IEEE Transactions on Software Engineering, 41(6):582–602, June 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6985636.

Cardellini:2012:MFQ

[CCG^{+}12] V. Cardellini, E. Casalicchio, V. Grassi, S. Iannucci, F. L. Presti, and R. Mirandola. MOSES: A framework for QoS driven runtime adaptation of serviceoriented systems. IEEE Transactions on Software Engineering, 38(5):1138–1159, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5963694.

Collingbourne:2014:SCD

[CCK14] P. Collingbourne, C. Cadar, and P. H. J. Kelly. Symbolic crosschecking of data-parallel floating-point code. IEEE Transactions on Software Engineering, 40(7):710–737, July 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6698391.

Cinque:2013:ELA

[CCP13] M. Cinque, D. Cotroneo, and A. Pecchia. Event logs for the analysis of software failures: A rule-based approach. IEEE Transactions on Software Engineering, 39(6):806–821, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6320555.

Classen:2013:FTS

[CCS^{+}13] A. Classen, M. Cordy, P. Y. Schobbens, P. Heymans, A. Legay, and J. F. Raskin. Featured transition systems: Foundations for verifying variability-intensive systems and their application to LTL model checking. IEEE Transactions on Software Engineering, 39(8):1069–1089, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6389685.

Coronato:2012:TRP

[CD12] A. Coronato and G. De Pietro. Tools for the rapid prototyping of provably correct ambient intelligence applications. IEEE Transactions on Software Engineering, 38(4):975–991, July/August 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5963693.

Coppa:2014:ISP

[CDF14] E. Coppa, C. Demetrescu, and I. Finocchi. Inputsensitive profiling. IEEE Transactions on Software Engineering, 40(12):1185– 1205, December 2014. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6858059.

Cuadrado:2012:AES

[CDGC12] F. Cuadrado, J. C. Duenas, and R. Garcia-Carmona. An autonomous engine for services configuration and deployment. IEEE Transactions on Software Engineering, 38(3):520–536, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5728830.

Choetkiertikul:2018:PDC

[CDT^{+}18] Morakot Choetkiertikul, Hoa Khanh Dam, Truyen Tran, Aditya Ghose, and John Grundy. Predicting delivery capability in iterative software development. IEEE Transactions on Software Engineering, 44(6):551–573, June 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7898472/.

Choetkiertikul:2019:DLM

[CDT^{+}19] M. Choetkiertikul, H. K. Dam, T. Tran, T. Pham, A. Ghose, and T. Menzies. A deep learning model for estimating story points. IEEE Transactions on Software Engineering, 45(7):637–656, July 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Cataldo:2011:GEI

[CEM11] M. Cataldo, K. Ehrlich, and A. Mockus. Guest Editors’ introduction: Special section on the socio-technical environment of software development projects. IEEE Transactions on Software Engineering, 37(3):305–306, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5779016.

Cordeiro:2012:SBB

[CFMS12] L. Cordeiro, B. Fischer, and J. Marques-Silva. SMTbased bounded model checking for embedded ANSI-C software. IEEE Transactions on Software Engineering, 38(4):957–974, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5928354.

Cunha:2015:EEV

[CFMS15] J. Cunha, J. P. Fernandes, J. Mendes, and J. Saraiva. Embedding, evolution, and validation of model-driven spreadsheets. IEEE Transactions on Software Engineering, 41(3):241–263, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6915751.

Clariso:2019:SBS

[CGC19] R. Clarisó, C. A. González, and J. Cabot. Smart bound selection for the verification of UML/OCL class diagrams. IEEE Transactions on Software Engineering, 45 (4):412–426, April 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Cuadrado:2014:CMM

[CGdL14] J. S. Cuadrado, E. Guerra, and J. de Lara. A component model for model transformations. IEEE Transactions on Software Engineering, 40 (11):1042–1060, November 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6858077.

Cuadrado:2017:SAM

[CGdL17] J. S. Cuadrado, E. Guerra, and J. de Lara. Static analysis of model transformations. IEEE Transactions on Software Engineering, 43(9):868–897, September 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7765073.

Calinescu:2011:DQM

[CGK^{+}11] R. Calinescu, L. Grunske, M. Kwiatkowska, R. Mirandola, and G. Tamburrelli. Dynamic QoS management and optimization in servicebased systems. IEEE Transactions on Software Engineering, 37(3):387–409, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5611553.

Chari:2019:FRE

[CGMD19] G. Chari, D. Garbervetsky, S. Marr, and S. Ducasse. Fully reflective execution environments: Virtual machines for more flexible software. IEEE Transactions on Software Engineering, 45(9):858– 876, September 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Caporuscio:2016:GFD

[CGMM16] M. Caporuscio, V. Grassi, M. Marzolla, and R. Mirandola. GoPrime: A fully decentralized middleware for utility-aware service assembly. IEEE Transactions on Software Engineering, 42(2):136–152, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7243346.

Cataldo:2013:CBT

[CH13] M. Cataldo and J. D.

Herbsleb. Coordination breakdowns and their impact on development productivity and software failures. IEEE Transactions on Software Engineering, 39 (3):343–360, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
6205767.

Chalin:2010:ESA

[Cha10]  P. Chalin.  Engineering

a sound assertion semantics for the verifying compiler. IEEE Transactions on Software Engineering, 36(2):275–287, March/April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5262949.

Chen:2015:SST

[CK15] N. Chen and S. Kim. STAR: Stack trace based automatic crash reproduction via symbolic execution. IEEE Transactions on Software Engineering, 41(2):198–220, February 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6926857.

Casale:2012:BEW

[CKKR12] G. Casale, A. Kalbasi, D. Krishnamurthy, and J. Rolia. BURN: Enabling workload burstiness in customized service benchmarks. IEEE Transactions on Software Engineering, 38(4): 778–793, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
5928353.

Cooray:2013:PSA

[CKMR13] D. Cooray, E. Kouroshfar, S. Malek, and R. Roshandel. Proactive self-adaptation for improving the reliability of mission-critical, embedded, and mobile software. IEEE Transactions on Software Engineering, 39 (12):1714–1735, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6574866.

Cai:2016:DTD

[CL16] Y. Cai and Q. Lu. Dynamic testing for deadlocks via constraints. IEEE Transactions on Software Engineering, 42(9):825–842, September 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7423814.

Carr:2017:ACI

[CLP17] S. A. Carr, F. Logozzo, and M. Payer. Automatic contract insertion with CCBot. IEEE Transactions on Software Engineering, 43(8):701–714, August 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
7736073.

Casale:2012:DBM

[CMCS12] G. Casale, N. Mi, L. Cherkasova, and E. Smirni. Dealing with burstiness in multi-tier applications: Models and their parameterization. IEEE Transactions on Software Engineering, 38(5):1040–1053, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311395.

Concas:2011:DBE

[CMM^{+}11] G. Concas, M. Marchesi, A. Murgia, R. Tonelli, and I. Turnu. On the distribution of bugs in the eclipse system. IEEE Transactions on Software Engineering, 37(6):872– 877, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5928349.

Canfora:2019:LFT

[CMM^{+}19] G. Canfora, F. Martinelli, F. Mercaldo, V. Nardone, A. Santone, and C. A. Visaggio. LEILA: Formal tool for identifying mobile malicious behaviour. IEEE Transactions on Software Engineering, 45(12):1230– 1252, December 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Carvajal:2013:UTS

[CMSSS13] L. Carvajal, A. M. Moreno, M. I. Sánchez-Segura, and A. Seffah. Usability through software design. IEEE Transactions on Software Engineering, 39(11):1582–1596, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6523225.

Chen:2019:SBO

[CNKM19] J. Chen, V. Nair, R. Krishna, and T. Menzies. Sampling as a baseline optimizer for search-based software engineering. IEEE Transactions on Software Engineering, 45 (6):597–614, June 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Clarke:2015:ERB

[COLY15] P. Clarke, R. V. O’Connor, B. Leavy, and M. Yilmaz. Exploring the relationship between software process adaptive capability and organisational performance. IEEE Transactions on Software Engineering, 41 (12):1169–1183, December 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7214314.

Cotroneo:2016:RTT

[CPR16] D. Cotroneo, R. Pietrantuono, and S. Russo. RE-LAI testing: A technique to assess and improve software reliability. IEEE Transactions on Software Engineering, 42(5):452–475, May 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7299696.

Chen:2012:DDC

[CPTT12] T. Y. Chen, P. L. Poon, S. F. Tang, and T. H. Tse. DESSERT: a DividEand-conquer methodology for identifying categorieS, choiceS, and choicE Relations for Test case generation. IEEE Transactions on Software Engineering, 38(4):794–809, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5963695.

Carnevali:2013:CVH

[CPV13] L. Carnevali, A. Pinzuti, and E. Vicario. Compositional verification for hierarchical scheduling of realtime systems. IEEE Transactions on Software Engineering, 39(5):638–657, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6264049.

Carnevali:2011:PPT

[CRV11] L. Carnevali, L. Ridi, and E. Vicario. Putting preemptive time Petri nets to work in a V-Model SW life cycle. IEEE Transactions on Software Engineering, 37(6):826– 844, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5680913.

Carnevali:2013:QAI

[CRV13] L. Carnevali, L. Ridi, and E. Vicario. A quantitative approach to input generation in real-time testing of stochastic systems. IEEE Transactions on Software Engineering, 39(3):292–304, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6226426.

Chavez:2016:ACC

[CSF^{+}16] H. M. Chavez, W. Shen, R. B.

France, B. A. Mechling, and G. Li. An approach to checking consistency between UML class model and its Java implementation. IEEE Transactions on Software Engineering, 42(4):322–344, April 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7294689.

Chen:2016:FEP

[CSJ^{+}16] T. H. Chen, W. Shang, Z. M. Jiang, A. E. Hassan, M. Nasser, and P. Flora. Finding and evaluating the performance impact of redundant data access for applications that are developed using object-relational mapping frameworks. IEEE Transactions on Software Engineering, 42(12):1148– 1161, December 2016. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7451264.

Crnkovic:2011:CFS

[CSVC11] I. Crnkovic, S. Sentilles, A. Vulgarakis, and M. R. V. Chaudron. A classification framework for software component models. IEEE Transactions on Software Engineering, 37(5): 593–615, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5587419.

Chen:2013:EEU

[CT13] H. Y. Chen and T. H. Tse.

Equality to equals and unequals: A revisit of the equivalence and nonequivalence criteria in class-level testing of object-oriented software. IEEE Transactions on Software Engineering, 39 (11):1549–1563, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6570471.

Chen:2011:SPI

[CTZ11] T. Y. Chen, T. H. Tse, and Z. Q. Zhou. Semi-Proving: An integrated method for program proving, testing, and debugging. IEEE Transactions on Software Engineering, 37(1): 109–125, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5406529.

Calinescu:2018:ETS

[CWG^{+}18] Radu Calinescu, Danny Weyns, Simos Gerasimou, Muhammad Usman Iftikhar, Ibrahim Habli, and Tim Kelly. Engineering trustworthy self-adaptive software with dynamic assurance cases. IEEE Transactions on Software Engineering, 44(11):1039–1069, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8008800/.

Casagrande:2014:NKS

[CWW^{+}14] E. Casagrande, S. Woldeamlak, W. L. Woon, H. H. Zeineldin, and D. Svetinovic. NLP-KAOS for systems goal elicitation: Smart metering system case study. IEEE Transactions on Software Engineering, 40(10):941–956, October 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6857327.

Cai:2019:DRS

[CXK^{+}19] Y. Cai, L. Xiao, R. Kazman, R. Mo, and Q. Feng. Design rule spaces: A new model for representing and analyzing software architecture. IEEE Transactions on Software Engineering, 45(7): 657–682, July 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Chen:2011:DSU

[CYH^{+}11] H. Chen, J. Yu, C. Hang, B. Zang, and P. C. Yew. Dynamic software updating using a relaxed consistency model. IEEE Transactions on Software Engineering, 37(5):

679–694, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5551162.

Chen:2013:ACO

[CZ13] W. N. Chen and J. Zhang.

Ant colony optimization for software project scheduling and staffing with an eventbased scheduler. IEEE Transactions on Software Engineering, 39(1):1–17, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6165315.

Cornelissen:2011:CEP

[CZvD11] B. Cornelissen, A. Zaidman, and A. van Deursen. A controlled experiment for program comprehension through trace visualization. IEEE Transactions on Software Engineering, 37(3):341–355, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5441291.

Dwyer:2016:CSS

[DBF^{+}16] M. B. Dwyer, E. Bodden, B. Fitzgerald, M. Kim, S. Kim, A. J. Ko, E. Mendes, R. Mirandola, A. Moreira, F. Shull, S. Siegel, T. Xie, and C. Zhang. Connecting and serving the software engineering community. IEEE Transactions on Software Engineering, 42(3):203–280, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7432058.

diBella:2013:PPS

[dBFP^{+}13] E. di Bella, I. Fronza, N. Phaphoom, A. Sillitti, G. Succi, and J. Vlasenko. Pair programming and software defects — a large, industrial case study. IEEE Transactions on Software Engineering, 39(7):930–953, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6331491.

Dashevskyi:2019:STD

[DBM19] S. Dashevskyi, A. D. Brucker, and F. Massacci. A screening test for disclosed vulnerabilities in FOSS components. IEEE Transactions on Software Engineering, 45(10): 945–966, October 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

deCaso:2012:AAC

[dCBGU12] G. de Caso, V. Braberman, D. Garbervetsky, and S. Uchitel. Automated abstractions for contract validation. IEEE Transactions on Software Engineering, 38(1): 141–162, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5639021.

Diaz:2014:SVN

[DCMS14] G. Díaz, M. E. Cambronero, E. Martínez, and G. Schneider. Specification and verification of normative texts using C-O diagrams. IEEE Transactions on Software Engineering, 40(8):795–817, August 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6657668.

daCosta:2017:FER

[dCMS^{+}17] D. A. da Costa, S. McIntosh, W. Shang, U. Kulesza, R. Coelho, and A. E. Hassan. A framework for evaluating the results of the SZZ approach for identifying bugintroducing changes. IEEE Transactions on Software Engineering, 43(7):641–657, July 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7588121.

Damevski:2018:PFD

[DCS^{+}18] Kostadin Damevski, Hui Chen, David C. Shepherd, Nicholas A. Kraft, and Lori Pollock. Predicting future developer behavior in the IDE using topic models. IEEE Transactions on Software Engineering, 44(11):1100–1111, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8024001/.

DeLucia:2011:ISC

[DDO11] A. De Lucia, M. Di Penta, and R. Oliveto. Improving source code lexicon via traceability and information retrieval. IEEE Transactions on Software Engineering, 37(2):205–227, March/ April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5601742.

Degano:2016:TCL

[DFG16] P. Degano, G. L. Ferrari, and L. Galletta. A two-component language for adaptation: Design, semantics and program analysis. IEEE Transactions on Software Engineering, 42(6):505–529, June 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
7314969.

Dieste:2011:SRA

[DJ11] O. Dieste and N. Juristo.

Systematic review and aggregation of empirical studies on elicitation techniques. IEEE Transactions on Software Engineering, 37(2): 283–304, March/April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5416730.

Dallmeier:2012:AGT

[DKM^{+}12] V. Dallmeier, N. Knopp, C. Mallon, G. Fraser, S. Hack, and A. Zeller. Automatically generating test cases for specification mining. IEEE Transactions on Software Engineering, 38(2):243–257, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6044587.

Duran-Limon:2015:OBP

[DLGRCBC15] H. A. Duran-Limon, C. A.

Garcia-Rios, F. E. Castillo-Barrera, and R. Capilla. An ontology-based product architecture derivation approach. IEEE Transactions on Software Engineering, 41 (12):1153–1168, December 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7134799.

Duboc:2013:SES

[DLR13] L. Duboc, E. Letier, and D. S. Rosenblum. Systematic elaboration of scalability requirements through goal-obstacle analysis. IEEE Transactions on Software Engineering, 39(1):119–140, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6152130.

delaVara:2016:ISS

[dlVBWM16] J. L. de la Vara, M. Borg, K. Wnuk, and L. Moonen. An industrial survey of safety evidence change impact analysis practice. IEEE Transactions on Software Engineering, 42 (12):1095–1117, December 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7450627.

Damas:2014:ACD

[DLvL14] C. Damas, B. Lambeau, and A. van Lamsweerde. Analyzing critical decision-based processes. IEEE Transactions on Software Engineering, 40(4):338–365, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6776491.

Dorn:2019:AET

[DLWF19] Jonathan Dorn, Jeremy Lacomis, Westley Weimer, and Stephanie Forrest. Automatically exploring tradeoffs between software output fidelity and energy costs. IEEE Transactions on Software Engineering, 45 (3):219–236, March 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8118155/.

Do:2010:ETC

[DMTR10] H. Do, S. Mirarab, L. Tahvildari, and G. Rothermel. The effects of time constraints on test case prioritization: A series of controlled experiments. IEEE Transactions on Software Engineering, 36(5): 593–617, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5482587.

Dias-Neto:2014:SCS

[DNT14] A. C. Dias-Neto and G. H.

Travassos. Supporting the combined selection of modelbased testing techniques. IEEE Transactions on Software Engineering, 40(10):

1025–1041, October 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6776497.

DiNucci:2018:DCB

[DPD^{+}18] D. Di Nucci, F. Palomba, G. De Rosa, G. Bavota, R. Oliveto, and A. De Lucia. A developer centered bug prediction model. IEEE Transactions on Software Engineering, 44(1):5–24, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7835258.

DiPietro:2012:MCS

[DPS12] I. Di Pietro, F. Pagliarecci, and L. Spalazzi. Model checking semantically annotated services. IEEE Transactions on Software Engineering, 38(3):592–608, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5680919.

Dagenais:2014:UTL

[DR14] B. Dagenais and M. P. Robillard. Using traceability links to recommend adaptive changes for documentation evolution. IEEE Transactions on Software Engineering, 40 (11):1126–1146, November 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6878435.

Dwyer:2016:EJF

[DR16] M. B. Dwyer and D. S. Rosenblum. Editorial: Journal-first publication for the software engineering community. IEEE Transactions on Software Engineering, 42(1):1, January 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7374796.

Demsky:2011:BLS

[DS11] B. Demsky and S. Sundaramurthy. Bristlecone: Language support for robust software applications. IEEE Transactions on Software Engineering, 37(1):4– 23, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5416725.

Draxler:2014:KDE

[DSB14] S. Draxler, G. Stevens, and A. Boden. Keeping the development environment up to date — a study of the situated practices of appropriating the Eclipse IDE.

IEEE Transactions on Software Engineering, 40(11): 1061–1074, November 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6898825.

Maldonado:2017:UNL

[dSMST17] E. d. S. Maldonado, E. Shihab, and N. Tsantalis. Using natural language processing to automatically detect self-admitted technical debt. IEEE Transactions on Software Engineering, 43 (11):1044–1062, November 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7820211.

Distefano:2011:UPN

[DSP11] S. Distefano, M. Scarpa, and A. Puliafito. From UML to Petri nets: The PCM-based methodology. IEEE Transactions on Software Engineering, 37(1):65– 79, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5396344.

deSouza:2011:ANW

[dSR11] C. R. B. de Souza and D. F. Redmiles. The awareness network, to whom should I display my actions? And, whose actions should I monitor? IEEE Transactions on Software Engineering, 37(3):325–340, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5710950.

Damevski:2017:MSD

[DSSP17] K. Damevski, D. C. Shepherd, J. Schneider, and L. Pollock. Mining sequences of developer interactions in visual studio for usage smells. IEEE Transactions on Software Engineering, 43(4):359–371, April 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7516714.

Dejaeger:2013:TCS

[DVB13] K. Dejaeger, T. Verbraken, and B. Baesens. Toward comprehensible software fault prediction models using Bayesian network classifiers. IEEE Transactions on Software Engineering, 39(2):237–257, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6175912.

Dejaeger:2012:DMT

[DVMB12] K. Dejaeger, W. Verbeke, D. Martens, and B. Baesens. Data mining techniques for software effort estimation: A comparative study. IEEE Transactions on Software Engineering, 38(2):375–397, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5928350.

Dwyer:2015:SJE

[Dwy15] M. B. Dwyer. State of the journal editorial. IEEE Transactions on Software Engineering, 41(1):1–2, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7004121.

Dwyer:2017:ITV

[Dwy17] M. B. Dwyer. Improving timeliness and visibility in publishing software engineering research. IEEE Transactions on Software Engineering, 43(3):205–206, March 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7876880.

Dwyer:2018:SJ

[Dwy18] M. Dwyer. State of the journal. IEEE Transactions on Software Engineering, 44(1):1–2, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
8249614.

Dou:2017:CDR

[DXCW17] W. Dou, C. Xu, S. C. Cheung, and J. Wei. CACheck: Detecting and repairing cell arrays in spreadsheets. IEEE Transactions on Software Engineering, 43(3):226–251, March 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7498607.

El-Attar:2015:EUS

[EALK^{+}15] M. El-Attar, H. Luqman, P. Kárpáti, G. Sindre, and A. L. Opdahl. Extending the UML statecharts notation to model security aspects. IEEE Transactions on Software Engineering, 41(7):661–690, July 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7042284.

Emberson:2010:SSS

[EB10] P. Emberson and I. Bate.

Stressing search with scenarios for flexible solutions to real-time task allocation problems. IEEE Transactions on Software Engineering, 36(5):704–718, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5262947.

Esfahani:2013:LBF

[EEM13] N. Esfahani, A. Elkhodary, and S. Malek. A learningbased framework for engineering feature-oriented selfadaptive software systems. IEEE Transactions on Software Engineering, 39(11): 1467–1493, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6574860.

Egyed:2011:ADT

[Egy11] A. Egyed. Automatically detecting and tracking inconsistencies in software design models. IEEE Transactions on Software Engineering, 37(2):188–204, March/ April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5432227.

El-Kassas:2016:ECC

[EKAYW16] W. S. El-Kassas, B. A. Abdullah, A. H. Yousef, and A. M. Wahba. Enhanced code conversion approach for the integrated crossplatform mobile development (icpmd). IEEE Transactions on Software Engineering, 42 (11):1036–1053, November 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7442177.

Eghbali:2016:TCP

[ET16] S. Eghbali and L. Tahvildari. Test case prioritization using lexicographical ordering. IEEE Transactions on Software Engineering, 42 (12):1178–1195, December 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7456343.

Haidry:2013:UDS

[eZHM13] S. e. Z. Haidry and T. Miller.

Using dependency structures for prioritization of functional test suites. IEEE Transactions on Software Engineering, 39(2):258–275, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
6189361.

Fraser:2013:WTS

[FA13] G. Fraser and A. Arcuri. Whole test suite generation. IEEE Transactions on Software Engineering, 39 (2):276–291, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6152257.

Fernandez-Amoros:2014:SAE

[FAHCC14] D. Fernandez-Amoros, R. Heradio, J. A. Cerrada, and C. Cerrada. A scalable approach to exact model and commonality counting for extended feature models. IEEE Transactions on Software Engineering, 40(9): 895–910, September 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6835200.

Falessi:2013:EPI

[FCC13] D. Falessi, G. Cantone, and G. Canfora. Empirical principles and an industrial case study in retrieving equivalent requirements via natural language processing techniques. IEEE Transactions on Software Engineering, 39(1):18–44, January 2013. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6112783.

Fowkes:2017:ASC

[FCR^{+}17] J. Fowkes, P. Chanthirasegaran, R. Ranca, M. Allamanis, M. Lapata, and C. Sutton. Autofolding for source code summarization. IEEE Transactions on Software Engineering, 43 (12):1095–1109, December 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7843666.

Franzago:2018:CMD

[FDMM18] Mirco Franzago, Davide Di Ruscio, Ivano Malavolta, and Henry Muccini. Collaborative model-driven software engineering: A classification framework and a research map. IEEE Transactions on Software Engineering, 44 (12):1146–1175, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8047991/.

Franca:2020:MSS

[FdS20] C. França, F. Q. B. da Silva, and H. Sharp. Motivation and satisfaction of software engineers. IEEE Transactions on Software Engineering, 46(2):118–140, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Fucci:2017:DTD

[FET^{+}17] D. Fucci, H. Erdogmus, B. Turhan, M. Oivo, and N. Juristo. A dissection of the test-driven development process: Does it really matter to test-first or to test-last? IEEE Transactions on Software Engineering, 43(7):597–614, July 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7592412.

Friedrich:2010:EHR

[FFM^{+}10] G. Friedrich, M. G. Fugini, E. Mussi, B. Pernici, and G. Tagni. Exception handling for repair in servicebased processes. IEEE Transactions on Software Engineering, 36(2):198–215, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5383376.

Feng:2011:CTE

[FPTO11] X. Feng, D. L. Parnas, T. H. Tse, and T. O’Callaghan. A comparison of tabular expression-based testing strategies. IEEE Transactions on Software Engineering, 37(5):616–634, September/October 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5975175.

Frey:2019:GBU

[FRA^{+}19] S. Frey, A. Rashid, P. Anthonysamy, M. Pinto-Albuquerque, and S. A. Naqvi. The good, the bad and the ugly: A study of security decisions in a cyber-physical systems game. IEEE Transactions on Software Engineering, 45 (5):521–536, May 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Falessi:2020:LHA

[FRGCH20] D. Falessi, J. Roll, J. L. C.

Guo, and J. Cleland-Huang. Leveraging historical associations between requirements and source code to identify impacted classes. IEEE Transactions on Software Engineering, 46(4):420–441, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Filieri:2016:SSA

[FTG16] A. Filieri, G. Tamburrelli, and C. Ghezzi. Supporting self-adaptation via quantitative verification and sensitivity analysis at run time. IEEE Transactions on Software Engineering, 42(1):75–99, January 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7083754.

Fleck:2017:MTM

[FTK^{+}17] M. Fleck, J. Troya, M. Kessentini, M. Wimmer, and B. Alkhazi. Model transformation modularization as a manyobjective optimization problem. IEEE Transactions on Software Engineering, 43 (11):1009–1032, November 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7820199.

Fraser:2012:MDG

[FZ12] G. Fraser and A. Zeller. Mutation-driven generation of unit tests and oracles. IEEE Transactions on Software Engineering, 38(2): 278–292, March/April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6019060.

Georg:2010:VTA

[GAB^{+}10] G. Georg, K. Anastasakis, B. Bordbar, S. H. Houmb, I. Ray, and M. Toahchoodee. Verification and trade-off analysis of security properties in UML system models. IEEE Transactions on Software Engineering, 36(3):338–356, May/ June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5432225.

Garousi:2010:GAB

[Gar10]  V. Garousi.  A genetic

algorithm-based stress test requirements generator tool and its empirical evaluation. IEEE Transactions on Software Engineering, 36(6):778– 797, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5383373.

Garcia-Banuelos:2018:CIC

[GBvBD^{+}18] Luciano García-Bañuelos, Nick R. T. P. van Beest, Marlon Dumas, Marcello La Rosa, and Willem Mertens. Complete and interpretable conformance checking of business processes. IEEE Transactions on Software Engineering, 44(3):262–290, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/document/7852436/.

Gao:2016:SGT

[GCZM16] Z. Gao, Z. Chen, Y. Zou, and A. M. Memon. SITAR:

GUI test script repair. IEEE Transactions on Software Engineering, 42(2): 170–186, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7214294.

Galeotti:2015:ILI

[GFM^{+}15] J. P. Galeotti, C. A. Furia, E. May, G. Fraser, and A. Zeller. Inferring loop invariants by mutation, dynamic analysis, and static checking. IEEE Transactions on Software Engineering, 41(10):1019–1037, October 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7105412.

Grundy:2013:GDS

[GHL^{+}13] J. C. Grundy, J. Hosking, K. N. Li, N. M. Ali, J. Huh, and R. L. Li. Generating domain-specific visual language tools from abstract visual specifications. IEEE Transactions on Software Engineering, 39(4):487–515, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6205768.

Giuffrida:2017:ALU

[GITT17] C. Giuffrida, C. Iorgulescu, G. Tamburrelli, and A. S. Tanenbaum. Automating live update for generic server programs. IEEE Transactions on Software Engineering, 43(3):207–225, March 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7497481.

Groce:2014:YOP

[GKZ^{+}14] A. Groce, T. Kulesza, C. Zhang, S. Shamasunder, M. Burnett, W. K. Wong, S. Stumpf, S. Das, A. Shinsel, F. Bice, and K. McIntosh. You are the only possible oracle: Effective test selection for end users of interactive machine learning systems. IEEE Transactions on Software Engineering, 40(3):307–323, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6682887.

Gazzola:2019:ASR

[GMM19] Luca Gazzola, Daniela Micucci, and Leonardo Mariani. Automatic software repair: A survey. IEEE Transactions on Software Engineering, 45(1):34–67, January 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8089448/.

Goseva-Popstojanova:2010:GEI

[GPK10] K. Goseva-Popstojanova and K. Kanoun. Guest Editors’ introduction to the special section on evaluation and improvement of software dependability. IEEE Transactions on Software Engineering, 36(3):306–308, May/ June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5473900.

Giardino:2016:SDS

[GPU^{+}16] C. Giardino, N. Paternoster, M. Unterkalmsteiner, T. Gorschek, and P. Abrahamsson. Software development in startup companies: The greenfield startup model. IEEE Transactions on Software Engineering, 42(6):585–604, June 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7360225.

Grbac:2013:SRQ

[GRH13] T. Galinac Grbac, P. Runeson, and D. Huljenić. A second replicated quantitative analysis of fault distributions in complex software systems. IEEE Transactions on Software Engineering, 39(4):462–476, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6235962.

Gay:2017:ASM

[GRH17] G. Gay, S. Rayadurgam, and M. P. E. Heimdahl. Automated steering of modelbased test oracles to admit real program behaviors. IEEE Transactions on Software Engineering, 43(6):531–555, June 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7583721.

Garcia:2010:GEI

[GRI10] A. Garcia, A. Romanovsky, and V. Issarny. Guest Editors’ introduction to the special section on exception handling: From requirements to software maintenance. IEEE Transactions on Software Engineering, 36(2):147–149, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5439569.

Galeotti:2013:TES

[GRPF13] J. P. Galeotti, N. Rosner, C. G. López Pombo, and M. F. Frias. TACO: Efficient SAT-based bounded verification using symmetry breaking and tight bounds. IEEE Transactions on Software Engineering, 39(9): 1283–1307, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6482141.

Geipel:2012:LBD

[GS12] M. M. Geipel and F. Schweitzer. The link between dependency and cochange: Empirical evidence. IEEE Transactions on Software Engineering, 38(6):1432–1444, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6363462.

Gay:2015:AOD

[GSWH15a] G. Gay, M. Staats, M. Whalen, and M. P. E. Heimdahl. Automated oracle data selection support. IEEE Transactions on Software Engineering, 41 (11):1119–1137, November 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7112189.

Gay:2015:RCD

[GSWH15b] G. Gay, M. Staats, M. Whalen, and M. P. E. Heimdahl. The risks of coverage-directed test case generation. IEEE Transactions on Software Engineering, 41(8):803–819, August 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7081779.

Gao:2019:MAT

[GW19] Ruizhi Gao and W. Eric Wong. MSeer — an advanced technique for locating multiple bugs in parallel. IEEE Transactions on Software Engineering, 45 (3):301–318, March 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8119545/.

Galster:2014:VSS

[GWT^{+}14] M. Galster, D. Weyns, D. Tofan, B. Michalik, and P. Avgeriou. Variability in software systems — a systematic literature review. IEEE Transactions on Software Engineering, 40(3):282–306, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6682901.

Guo:2015:FDT

[GZS^{+}15] X. Guo, M. Zhou, X. Song, M. Gu, and J. Sun. First, debug the test oracle. IEEE Transactions on Software Engineering, 41(10):986–1000, October 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7091939.

Harvie:2016:TSA

[HA16] D. P. Harvie and A. Agah. Targeted scrum: Applying mission command to agile software development. IEEE Transactions on Software Engineering, 42(5):476–489, May 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7296686.

Hannay:2010:EPP

[HAES10] J. E. Hannay, E. Arisholm, H. Engvik, and D. I. K. Sjoberg. Effects of personality on pair programming. IEEE Transactions on Software Engineering, 36 (1):61–80, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5089333.

Hatton:2014:CIS

[Hat14] L. Hatton. Conservation of information: Software’s hidden clockwork? IEEE Transactions on Software Engineering, 40(5):450–460, May 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6784340.

Hall:2012:SLR

[HBB^{+}12] T. Hall, S. Beecham, D. Bowes, D. Gray, and S. Counsell. A systematic literature review on fault prediction performance in software engineering. IEEE Transactions on Software Engineering, 38(6):1276–1304, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6035727.

Hayden:2013:PSE

[HBC13] R. A. Hayden, J. T. Bradley, and A. Clark. Performance specification and evaluation with unified stochastic probes and fluid analysis. IEEE Transactions on Software Engineering, 39(1):97–118, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
6133297.

Huber:2017:MBS

[HBS^{+}17] N. Huber, F. Brosig, S. Spinner, S. Kounev, and M. Bähr. Model-based self-aware performance and resource management using the Descartes modeling language. IEEE Transactions on Software Engineering, 43(5):432–452, May 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7577879.

Han:2018:TPA

[HC18] Ah-Rim Han and Sungdeok Cha. Two-phase assessment approach to improve the efficiency of refactoring identification. IEEE Transactions on Software Engineering, 44(10):1001–1023, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7990580/.

Heydarnoori:2012:TSF

[HCBB12] A. Heydarnoori, K. Czarnecki, W. Binder, and T. T. Bartolomei. Two studies of framework-usage templates extracted from dynamic traces. IEEE Transactions on Software Engineering, 38(6):1464–1487, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5975174.

Hariri:2013:SDA

[HCHM^{+}13] N. Hariri, C. Castro-Herrera, M. Mirakhorli, J. Cleland-Huang, and B. Mobasher. Supporting domain analysis through mining and recommending features from online product listings. IEEE Transactions on Software Engineering, 39(12):1736– 1752, December 2013. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6582404.

Herbold:2017:CSR

[Her17] S. Herbold. Comments on ScottKnottESD in response to “An empirical comparison of model validation techniques for defect prediction models”. IEEE Transactions on Software Engineering, 43 (11):1091–1094, November 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

8024011. See [TMHM17].

He:2014:FCE

[HHY^{+}14] Q. He, J. Han, Y. Yang, H. Jin, J. G. Schneider, and S. Versteeg. Formulating cost-effective monitoring strategies for servicebased systems. IEEE Transactions on Software Engineering, 40(5):461–482, May 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6642029.

Hierons:2012:ODT

[Hie12]  R. M. Hierons.  Ora-

cles for distributed testing. IEEE Transactions on Software Engineering, 38(3): 629–641, May/June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5750006.

Hierons:2015:GCC

[Hie15]  R. M. Hierons.  Gener-

ating complete controllable test suites for distributed testing. IEEE Transactions on Software Engineering, 41(3):279–293, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6930767.

Hierons:2017:TPF

[Hie17]  R. M. Hierons.  Testing

from partial finite state machines without harmonised traces. IEEE Transactions on Software Engineering, 43 (11):1033–1043, November 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7815407.

Hertis:2014:EAB

[HJ14] M. Hertis and M. B. Juric.

An empirical analysis of business process execution language usage. IEEE Transactions on Software Engineering, 40(8):738–757, August 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6812231.

Holzmann:2011:SVT

[HJG11] G. J. Holzmann, R. Joshi, and A. Groce. Swarm verification techniques. IEEE Transactions on Software Engineering, 37(6):845–857, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5661793.

Hebig:2017:ACE

[HKB17] R. Hebig, D. E. Khelladi, and R. Bendraou. Approaches to co-evolution of metamodels and models: A survey. IEEE Transactions on Software Engineering, 43(5):396–414, May 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7569018.

Harter:2012:DSP

[HKS12] D. E. Harter, C. F. Kemerer, and S. A. Slaughter. Does software process improvement reduce the severity of defects? A longitudinal field study. IEEE Transactions on Software Engineering, 38(4):810–827, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5928358.

Horwitz:2010:BDO

[HLP10] S. Horwitz, B. Liblit, and M. Polishchuk. Better debugging via output tracing and callstack-sensitive slicing. IEEE Transactions on Software Engineering, 36 (1):7–19, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5282499.

Harman:2010:SBS

[HM10a] M. Harman and A. Mansouri. Search based software engineering: Introduction to the special issue of the IEEE Transactions on Software Engineering. IEEE Transactions on Software Engineering, 36(6):737– 741, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5644735.

Harman:2010:TES

[HM10b] M. Harman and P. McMinn.

A theoretical and empirical study of search-based testing: Local, global, and hybrid search. IEEE Transactions on Software Engineering, 36(2):226–247, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5342440.

Harrison:2019:VTA

[HMC19] M. D. Harrison, P. Masci, and J. C. Campos. Verification templates for the analysis of user interface software design. IEEE Transactions on Software Engineering, 45(8): 802–822, August 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Habayeb:2018:UHM

[HMMB18] Mayy Habayeb, Syed Shariyar Murtaza, Andriy Miranskyy, and Ayse Basar Bener. On the use of hidden Markov model to predict the time to fix bugs. IEEE Transactions on Software Engineering, 44 (12):1224–1244, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8052546/.

Hoda:2013:SOR

[HNM13] R. Hoda, J. Noble, and S. Marshall. Self-organizing roles on agile software development teams. IEEE Transactions on Software Engineering, 39(3):422–444, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6197202.

Hoang:2019:NCM

[HOLL19] T. Hoang, R. J. Oentaryo, T. B. Le, and D. Lo. Networkclustered multi-modal bug localization. IEEE Transactions on Software Engineering, 45(10):1002–1023, October 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Henard:2014:BCE

[HPP^{+}14] C. Henard, M. Papadakis, G. Perrouin, J. Klein, P. Heymans, and Y. Le Traon. Bypassing the combinatorial explosion: Using similarity to generate and prioritize T -wise test configurations for software product lines. IEEE Transactions on Software Engineering, 40(7):650–670, July 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6823132.

Huang:2016:GTF

[HPSW16] C. H. Huang, D. A. Peled, S. Schewe, and F. Wang. A game-theoretic foundation for the maximum software resilience against dense errors. IEEE Transactions on Software Engineering, 42(7):605–622, July 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7360234.

Hayden:2012:EDS

[HSH^{+}12] C. M. Hayden, E. K. Smith, E. A. Hardisty, M. Hicks, and J. S. Foster. Evaluating dynamic software update safety using systematic testing. IEEE Transactions on Software Engineering, 38(6):1340–1354, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6035725.

Hierons:2016:PAT

[HT16] R. M. Hierons and U. C.

Türker. Parallel algorithms for testing finite state machines:generating UIO sequences. IEEE Transactions on Software Engineering, 42 (11):1077–1091, November 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7429774.

Herbold:2018:CSB

[HTG18] Steffen Herbold, Alexander Trautsch, and Jens Grabowski. A comparative study to benchmark crossproject defect prediction approaches. IEEE Transactions on Software Engineering, 44(9):811–833, September 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7972992/. See

[HTG19a].

Herbold:2019:CCS

[HTG19a] S. Herbold, A. Trautsch, and J. Grabowski. Correction of A Comparative Study to Benchmark Cross-Project Defect Prediction Approaches. IEEE Transactions on Software Engineering, 45 (6):632–636, June 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). See [HTG18].

Hosseini:2019:SLR

[HTG19b] Seyedrebvar Hosseini, Burak Turhan, and Dimuthu Gunarathna. A systematic literature review and metaanalysis on cross project defect prediction. IEEE Transactions on Software Engineering, 45(2):111–147, February 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8097045/.

Hall:2018:EIE

[HWM18] Mathew Hall, Neil Walkinshaw, and Phil McMinn. Effectively incorporating expert knowledge in automated software remodularisation. IEEE Transactions on Software Engineering, 44(7):613–630, July 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8259332/.

Henderson:2017:DPN

[HYH^{+}17] A. Henderson, L. K. Yan, X. Hu, A. Prakash, H. Yin, and S. McCamant. DECAF: A platform-neutral wholesystem dynamic binary analysis platform. IEEE Transactions on Software Engineering, 43(2):164–184, February 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7506264.

He:2014:QAS

[HYJY14] Q. He, J. Yan, H. Jin, and Y. Yang. Quality-aware service selection for servicebased systems based on iterative multi-attribute combinatorial auction. IEEE Transactions on Software Engineering, 40(2):192–215, February 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6702520.

Hao:2016:OTC

[HZZ^{+}16] D. Hao, L. Zhang, L. Zang, Y. Wang, X. Wu, and T. Xie. To be optimal or not in testcase prioritization. IEEE Transactions on Software Engineering, 42(5):490–505, May 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7314957.

He:2017:KSB

[HZZ^{+}17] Q. He, R. Zhou, X. Zhang, Y. Wang, D. Ye, F. Chen, J. C. Grundy, and Y. Yang. Keyword search for building service-based systems. IEEE Transactions on Software Engineering, 43(7):658–674, July 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
7731135.

Ikeda:2013:CEM

[IJK13] S. Ikeda, M. Jibiki, and Y. Kuno. Coverage estimation in model checking with bitstate hashing. IEEE Transactions on Software Engineering, 39(4):477–486, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6226428.

Itkonen:2013:RTK

[IML13] J. Itkonen, M. V. Mäntylä, and C. Lassenius. The role of the tester’s knowledge in exploratory software testing. IEEE Transactions on Software Engineering, 39(5):707–724, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6298893.

Itzik:2016:VAR

[IRBW16] N. Itzik, I. Reinhartz-Berger, and Y. Wand. Variability analysis of requirements: Considering behavioral differences and reflecting stakeholders’ perspectives. IEEE Transactions on Software Engineering, 42(7):687–706, July 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7366597.

Jorgensen:2011:IIM

[JG11] M. Jorgensen and S. Grimstad. The impact of irrelevant and misleading information on software development effort estimates: A randomized controlled field experiment. IEEE Transactions on Software Engineering, 37(5): 695–707, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5551161.

Jorgensen:2012:SDE

[JG12] M. Jorgensen and S. Grimstad. Software development estimation biases: The role of interdependence. IEEE Transactions on Software Engineering, 38(3):677–693, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6193066.

Janevski:2013:SRW

[JGP13] N. Janevski and K. Goseva-Popstojanova. Session reliability of Web systems under heavy-tailed workloads: An approach based on design and analysis of experiments. IEEE Transactions on Software Engineering, 39(8):1157–1178, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6409359.

Jia:2011:ASD

[JH11] Y. Jia and M. Harman. An analysis and survey of the development of mutation testing. IEEE Transactions on Software Engineering, 37(5): 649–678, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5487526.

Jiang:2015:SLT

[JH15] Z. M. Jiang and A. E. Hassan. A survey on load testing of large-scale software systems. IEEE Transactions on Software Engineering, 41 (11):1091–1118, November 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7123673.

Jhi:2015:PCU

[JJW^{+}15] Y. C. Jhi, X. Jia, X. Wang, S. Zhu, P. Liu, and D. Wu. Program characterization using runtime values and its application to software plagiarism detection. IEEE Transactions on Software Engineering, 41(9):925–943, September 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7076635.

Jovanovic:2015:IPS

[JLR15] A. Jovanović, D. Lime, and O. H. Roux. Integer parameter synthesis for realtime systems. IEEE Transactions on Software Engineering, 41(5):445–461, May 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6895298.

Jiau:2013:OOB

[JMC13] H. C. Jiau, L. W. Mar, and J. C. Chen. OBEY: Optimal batched refactoring plan execution for class responsibility redistribution. IEEE Transactions on Software Engineering, 39(9): 1245–1263, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6493333.

Jan:2019:AGT

[JPAB19] S. Jan, A. Panichella, A. Arcuri, and L. Briand. Automatic generation of tests to exploit XML injection vulnerabilities in web applications. IEEE Transactions on Software Engineering, 45(4): 335–362, April 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Jing:2017:ISB

[JWDX17] X. Y. Jing, F. Wu, X. Dong, and B. Xu. An improved SDA based defect prediction framework for both within-project and crossproject class-imbalance problems. IEEE Transactions on Software Engineering, 43(4):321–339, April 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7530877.

Kwon:2011:VEP

[KA11] Y. Kwon and G. Agha. Verifying the evolution of probability distributions governed by a DTMC. IEEE Transactions on Software Engineering, 37(1):126–141, January/February 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5557891.

Kalliamvakou:2019:WMG

[KBZ^{+}19] Eirini Kalliamvakou, Christian Bird, Thomas Zimmermann, Andrew Begel, Robert DeLine, and Daniel M. German. What makes a great manager of software engineers? IEEE Transactions on Software Engineering, 45(1):87–106, January 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8094304/.

Kastner:2014:VMC

[KDO14] C. Kästner, A. Dreiling, and K. Ostermann. Variability mining: Consistent semi-automatic detection of product-line features. IEEE Transactions on Software Engineering, 40(1):67–82, January 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6613490.

Kailasam:2013:OOT

[KGDS13] S. Kailasam, N. Gnanasambandam, J. Dharanipragada, and N. Sharma. Optimizing ordered throughput using autonomic cloud bursting schedulers. IEEE Transactions on Software Engineering, 39 (11):1564–1581, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6520852.

Kazman:2016:EEA

[KGM^{+}16] R. Kazman, D. Goldenson, I. Monarch, W. Nichols, and G. Valetto. Evaluating the effects of architectural documentation: A case study of a large scale open source project. IEEE Transactions on Software Engineering, 42(3):220–260, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7230299.

Khatchadourian:2012:PRR

[KGRX12] R. Khatchadourian, P. Greenwood, A. Rashid, and G. Xu. Pointcut rejuvenation: Recovering pointcut expressions in evolving aspect-oriented software. IEEE Transactions on Software Engineering, 38(3):642–657, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5710951.

Kim:2014:SMC

[Kim14] S. Kim. Synthesizing multithreaded code from real-time object-oriented models via schedulability-aware thread derivation. IEEE Transactions on Software Engineering, 40(4):413–426, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6617637.

Kim:2011:CSS

[KKK11] M. Kim, Y. Kim, and H. Kim. A comparative study of software model checkers as unit testing tools: An industrial case study. IEEE Transactions on Software Engineering, 37(2):146–160, March/ April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5510242.

Kimelman:2010:BAM

[KKMY10] D. Kimelman, M. Kimelman, D. Mandelin, and D. Yellin. Bayesian approaches to matching architectural diagrams. IEEE Transactions on Software Engineering, 36(2):248–274, March/April 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5232811.

Krogmann:2010:UGS

[KKR10] K. Krogmann, M. Kuperberg, and R. Reussner. Using genetic search for reverse engineering of parametric behavior models for performance prediction. IEEE Transactions on Software Engineering, 36(6):865–877, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5530323.

Kalbasi:2012:DSD

[KKRD12] A. Kalbasi, D. Krishnamurthy, J. Rolia, and S. Dawson. DEC: Service demand estimation with confidence. IEEE Transactions on Software Engineering, 38(3):561–578, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5728829.

Kessentini:2014:CPS

[KKS^{+}14] W. Kessentini, M. Kessentini, H. Sahraoui, S. Bechikh, and A. Ouni. A cooperative parallel searchbased software engineering approach for code-smells detection. IEEE Transactions on Software Engineering, 40(9):841–861, September 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6835187.

Kwan:2011:RMM

[KL11] T. W. Kwan and H. K. N.

Leung. A risk management methodology for project risk dependencies. IEEE Transactions on Software Engineering, 37(5):635–648, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5696725.

Krishna:2019:BBM

[KM19] R. Krishna and T. Menzies. Bellwethers: A baseline method for transfer learning. IEEE Transactions on Software Engineering, 45(11): 1081–1105, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Kocaguneli:2012:EEA

[KMBK12] E. Kocaguneli, T. Menzies, A. Bener, and J. W. Keung. Exploiting the essential assumptions of analogy-based effort estimation. IEEE Transactions on Software Engineering, 38(2):425–438, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5728833.

Krall:2015:GGA

[KMD15] J. Krall, T. Menzies, and M. Davies. GALE: Geometric active learning for search-based software engineering. IEEE Transactions on Software Engineering, 41(10):1001–1018, October 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7105950.

Kocaguneli:2012:VEE

[KMK12] E. Kocaguneli, T. Menzies, and J. W. Keung. On the value of ensemble effort estimation. IEEE Transactions on Software Engineering, 38(6):1403–1416, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6081882.

Kocaguneli:2013:ALE

[KMK^{+}13] E. Kocaguneli, T. Menzies, J. Keung, D. Cok, and R. Madachy. Active learning and effort estimation: Finding the essential content of software effort estimation data. IEEE Transactions on Software Engineering, 39(8):1040–1053, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6392173.

Kim:2013:ISS

[KNGW13] M. Kim, D. Notkin, D. Grossman, and G. Wilson. Identifying and summarizing systematic code changes via rule inference. IEEE Transactions on Software Engineering, 39(1):45–62, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6165314.

Krein:2016:MSJ

[KPJ^{+}16] J. L. Krein, L. Prechelt, N. Juristo, A. Nanthaamornphong, J. C. Carver, S. Vegas, C. D. Knutson, K. D. Seppi, and D. L. Eggett. A multi-site joint replication of a design patterns experiment using moderator variables to generalize across contexts. IEEE Transactions on Software Engineering, 42(4):302–321, April 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7294706.

Kintis:2018:DTM

[KPJ^{+}18] Marinos Kintis, Mike Papadakis, Yue Jia, Nicos Malevris, Yves Le Traon, and Mark Harman. Detecting trivial mutant equivalences via compiler optimisations. IEEE Transactions on Software Engineering, 44(4):308–333, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7882714.

Kramer:2010:ENE

[Kra10] J. Kramer. Editorial: A new Editor in Chief and the state of the journal. IEEE Transactions on Software Engineering, 36(1):1– 2, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5401364.

Kpodjedo:2013:MMM

[KRG^{+}13] S. Kpodjedo, F. Ricca, P. Galinier, G. Antoniol, and Y. G. Guéhéneuc. MAD-Match: Many-to-many approximate diagram matching for design comparison. IEEE Transactions on Software Engineering, 39(8): 1090–1111, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6464271.

Krishnamurthy:2011:WWA

[KRX11] D. Krishnamurthy, J. Rolia, and M. Xu. WAM

— the Weighted Average Method for predicting the performance of systems with bursts of customer sessions. IEEE Transactions on Software Engineering, 37(5):

718–735, September/October 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5953602.

Kwon:2012:ADU

[KS12] T. Kwon and Z. Su. Automatic detection of unsafe dynamic component loadings. IEEE Transactions on Software Engineering, 38(2):293–313, March/April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6065738.

Kamei:2013:LSE

[KSA^{+}13] Y. Kamei, E. Shihab, B. Adams, A. E. Hassan, A. Mockus, A. Sinha, and N. Ubayashi. A large-scale empirical study of just-in-time quality assurance. IEEE Transactions on Software Engineering, 39(6):757–773, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6341763.

Kwan:2011:DST

[KSD11] I. Kwan, A. Schroter, and D. Damian. Does sociotechnical congruence have an effect on software build success? A study of coordination in a software project. IEEE Transactions on Software Engineering, 37(3):307–324, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5740929.

Kitchenham:2013:TQH

[KSD^{+}13] B. Kitchenham, D. I. K.

Sjøberg, T. Dybå, O. P. Brereton, D. Budgen, M. Höst, and P. Runeson. Trends in the quality of human-centric software engineering experiments

— a quasi-experiment. IEEE Transactions on Software Engineering, 39(7):1002–1017, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6374196.

Kulkarni:2010:FPR

[KT10] D. Kulkarni and A. Tripathi.

A framework for programming robust context-aware applications. IEEE Transactions on Software Engineering, 36(2):184–197, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5396345.

Kim:2013:WSW

[KTKZ13] D. Kim, Y. Tao, S. Kim, and A. Zeller. Where should we fix this bug? A two-phase recommendation model. IEEE Transactions on Software Engineering, 39 (11):1597–1610, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6517844.

Kim:2011:WCS

[KWK^{+}11] D. Kim, X. Wang, S. Kim, A. Zeller, S. C. Cheung, and S. Park. Which crashes should I fix first?: Predicting top crashes at an early stage to prioritize debugging efforts. IEEE Transactions on Software Engineering, 37(3):430–447, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5711013.

Kim:2018:DSS

[KZDB18] Miryung Kim, Thomas Zimmermann, Robert DeLine, and Andrew Begel. Data scientists in software teams: State of the art and challenges. IEEE Transactions on Software Engineering, 44(11):1024–1038, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8046093/.

Kim:2014:ESR

[KZN14] M. Kim, T. Zimmermann, and N. Nagappan. An empirical study of refactoring challenges and benefits at Microsoft. IEEE Transactions on Software Engineering, 40(7):633–649, July 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6802406.

Leungwattanakit:2014:MSM

[LAH^{+}14] W. Leungwattanakit, C. Artho, M. Hagiya, Y. Tanabe, M. Yamamoto, and K. Takahashi. Modular software model checking for distributed systems. IEEE Transactions on Software Engineering, 40(5):483–501, May 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6645368.

Lin:2014:LAC

[LAL^{+}14] S. W. Lin, É. André, Y. Liu, J. Sun, and J. S. Dong. Learning assumptions for compositional verification of timed systems. IEEE Transactions on Software Engineering, 40(2):137–153, February 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?arnumber=
                                        6682903.

Latorre:2014:EDE

[Lat14] R. Latorre. Effects of developer experience on learning and applying unit test-driven development. IEEE Transactions on Software Engineering, 40(4):381–395, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6690135.

Limam:2010:ASS

[LB10] N. Limam and R. Boutaba. Assessing software service quality and trustworthiness at selection time. IEEE Transactions on Software Engineering, 36(4):559–574, July/August 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5383370.

Lawrance:2013:HPD

[LBB^{+}13] J. Lawrance, C. Bogart, M. Burnett, R. Bellamy, K. Rector, and S. D. Fleming. How programmers debug, revisited: An information foraging theory perspective. IEEE Transactions on Software Engineering, 39(2):197–215, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5674060.

LeGoues:2017:CCU

[LBFW17] C. Le Goues, Y. Brun, S. Forrest, and W. Weimer. Clarifications on the construction and use of the Many-Bugs benchmark. IEEE Transactions on Software Engineering, 43(11):1089– 1090, November 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=8048536.

Lim:2015:ICD

[LBK^{+}15] S. L. Lim, P. J. Bentley, N. Kanakam, F. Ishikawa, and S. Honiden. Investigating country differences in mobile app user behavior and challenges for software engineering. IEEE Transactions on Software Engineering, 41(1):40–64, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6913003.

Lutellier:2018:MIC

[LCG^{+}18] Thibaud Lutellier, Devin Chollak, Joshua Garcia, Lin Tan, Derek Rayside, Nenad Medvidović, and Robert Kroeger. Measuring the impact of code dependencies on software architecture recovery techniques. IEEE Transactions on Software Engineering, 44(2):159–181, February 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7859416/.

Liu:2013:VLO

[LCL^{+}13] Y. Liu, W. Chen, Y. A. Liu, J. Sun, S. J. Zhang, and J. S. Dong. Verifying linearizability via optimized refinement checking. IEEE Transactions on Software Engineering, 39(7):1018–1039, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6363443.

Liu:2012:FSB

[LCNM12] S. Liu, Y. Chen, F. Nagoya, and J. A. McDermid. Formal specification-based inspection for verification of programs. IEEE Transactions on Software Engineering, 38(5):1100–1122, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6035726.

Lerner:2010:EHP

[LCO^{+}10] B. S. Lerner, S. Christov, L. J. Osterweil, R. Bendraou, U. Kannengiesser, and A. Wise. Exception handling patterns for process modeling. IEEE Transactions on Software Engineering, 36(2):162–183, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5383369.

LaToza:2019:MP

[LDR^{+}19] T. D. LaToza, A. Di Lecce, F. Ricci, W. B. Towne, and A. van der Hoek. Microtask programming. IEEE Transactions on Software Engineering, 45(11):1106– 1124, November 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Lee:2013:TEC

[LED13] G. Lee, J. A. Espinosa, and W. H. DeLone. Task environment complexity, global team dispersion, process capabilities, and coordination in software development. IEEE Transactions on Software Engineering, 39 (12):1753–1771, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6583162.

Lim:2012:SUS

[LF12] S. L. Lim and A. Finkelstein. StakeRare: Using social networks and collaborative filtering for largescale requirements elicitation. IEEE Transactions on Software Engineering, 38 (3):707–735, May/June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5740931.

Lienhard:2012:SDA

[LGN12] A. Lienhard, T. Girba, and O. Nierstrasz. Specifying dynamic analyses by extending language semantics. IEEE Transactions on Software Engineering, 38 (3):694–706, May/June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5740933.

Liu:2013:MBI

[LGS13] H. Liu, X. Guo, and W. Shao. Monitor-based instant software refactoring. IEEE Transactions on Software Engineering, 39 (8):1112–1126, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6409360.

Lind:2012:PAS

[LH12] K. Lind and R. Heldal.

A practical approach to size estimation of embedded software components. IEEE Transactions on Software Engineering, 38(5):993– 1007, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5999672.

Lee:2014:GGD

[LH14] J. Lee and K. H. Hsu. GEA:

A goal-driven approach to discovering early aspects. IEEE Transactions on Software Engineering, 40(6):584–602, June 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6822606.

LeGoues:2015:MIB

[LHS^{+}15] C. Le Goues, N. Holtschulte, E. K. Smith, Y. Brun, P. Devanbu, S. Forrest, and W. Weimer. The ManyBugs and IntroClass benchmarks for automated repair of C programs. IEEE Transactions on Software Engineering, 41 (12):1236–1256, December 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7153570.

Li:2017:VEA

[LHWZ17] L. Li, M. Harman, F. Wu, and Y. Zhang. The value of exact analysis in requirements selection. IEEE Transactions on Software Engineering, 43(6):580–596, June 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7582553.

Li:2012:QAD

[Li12]  W. Li.  QoS assurance

for dynamic reconfiguration of component-based software systems. IEEE Transactions on Software Engineering, 38(3):658–676, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5740932.

Li:2019:MSP

[LJZ^{+}19] Z. Li, X. Jing, X. Zhu, H. Zhang, B. Xu, and S. Ying. On the multiple sources and privacy preservation issues for heterogeneous defect prediction. IEEE Transactions on Software Engineering, 45(4): 391–411, April 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Lillack:2018:TLT

[LKB18] Max Lillack, Christian Kästner, and Eric Bodden. Tracking load-time configuration options. IEEE Transactions on Software Engineering, 44 (12):1269–1291, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8049300/.

Lee:2015:IVH

[LKKS15] S. Lee, S. Kang, S. Kim, and M. Staats. The impact of view histories on edit recommendations. IEEE Transactions on Software Engineering, 41(3):314–330, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6926851.

Lano:2014:MTD

[LKR14] K. Lano and S. Kolahdouz-Rahimi. Model-transformation design patterns. IEEE Transactions on Software Engineering, 40(12):1224–1259, December 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6891324.

Liu:2010:EOS

[LKS10] Y. Liu, T. M. Khoshgoftaar, and N. Seliya. Evolutionary optimization of software quality modeling with multiple repositories. IEEE Transactions on Software Engineering, 36(6):852– 864, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5467094.

Liu:2014:HED

[LKTC14] H. Liu, F. C. Kuo, D. Towey, and T. Y. Chen. How effectively does metamorphic testing alleviate the oracle problem? IEEE Transactions on Software Engineering, 40(1):4–22, January 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6613484.

Lin:2014:RBQ

[LL14] C. T. Lin and Y. F. Li. Rate-based queueing simulation model of open source software debugging activities. IEEE Transactions on Software Engineering, 40 (11):1075–1099, November 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6891380.

Luo:2016:SRA

[LLC^{+}16] H. Luo, X. Liu, X. Chen, T. Long, and R. Jiang.

Software reliability analysis using weakest preconditions in linear assignment programs. IEEE Transactions on Software Engineering, 42(9):866–885, September 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7398131.

Li:2017:LSF

[LLHF17] F. Li, Z. Li, W. Huo, and X. Feng. Locating software faults based on minimum debugging frontier set. IEEE Transactions on Software Engineering, 43(8):760–776, August 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7755837.

Liu:2018:UDU

[LLL^{+}18] Xuanzhe Liu, Huoran Li, Xuan Lu, Tao Xie, Qiaozhu Mei, Feng Feng, and Hong Mei. Understanding diverse usage patterns from large-scale appstore-service profiles. IEEE Transactions on Software Engineering, 44(4):384–411, April 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7883939.

Liu:2015:IRO

[LLLW15] H. Liu, Q. Liu, Y. Liu, and Z. Wang. Identifying renaming opportunities by expanding conducted rename refactorings. IEEE Transactions on Software Engineering, 41(9):887–900, September 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7097720.

Liu:2016:DAF

[LLNL16] H. Liu, Q. Liu, Z. Niu, and Y. Liu. Dynamic and automatic feedback-based threshold adaptation for code smell detection. IEEE Transactions on Software Engineering, 42(6):544–558, June 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7337457.

Lallchandani:2011:DST

[LM11] J. T. Lallchandani and R. Mall. A dynamic slicing technique for UML architectural models. IEEE Transactions on Software Engineering, 37(6):737–771, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
5680909.

Leopold:2014:SPM

[LMP14] H. Leopold, J. Mendling, and A. Polyvyanyy. Supporting process model validation through natural language generation. IEEE Transactions on Software Engineering, 40(8):818–840, August 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6823180.

Liu:2012:SBS

[LMSN12] H. Liu, Z. Ma, W. Shao, and Z. Niu. Schedule of bad smell detection and resolution: A new way to save effort. IEEE Transactions on Software Engineering, 38(1): 220–235, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5680918.

Lanotte:2010:TPB

[LMST10] R. Lanotte, A. Maggiolo-Schettini, and A. Troina. Time and probability-based information flow analysis. IEEE Transactions on Software Engineering, 36(5): 719–734, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5383372.

Luo:2017:SBO

[LMW^{+}17] L. Luo, J. Ming, D. Wu, P. Liu, and S. Zhu. Semanticsbased obfuscation-resilient binary code similarity comparison with applications to software and algorithm plagiarism detection. IEEE Transactions on Software Engineering, 43(12):1157– 1177, December 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7823022.

Luo:2019:HDS

[LMZP19] Q. Luo, K. Moran, L. Zhang, and D. Poshyvanyk. How do static and dynamic test case prioritization techniques perform on modern software systems? An extensive study on GitHub projects. IEEE Transactions on Software Engineering, 45(11): 1054–1080, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

LeGoues:2012:GGM

[LNFW12] C. Le Goues, T. Nguyen, S. Forrest, and W. Weimer. GenProg: A generic method for automatic software repair. IEEE Transactions on Software Engineering, 38 (1):54–72, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6035728.

Lee:2016:DMI

[LNH^{+}16] T. Lee, J. Nam, D. Han, S. Kim, and H. Peter In. Developer micro interaction metrics for software defect prediction. IEEE Transactions on Software Engineering, 42(11):1015–1035, November 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7447797.

Li:2017:TOS

[LO17] N. Li and J. Offutt. Test oracle strategies for modelbased testing. IEEE Transactions on Software Engineering, 43(4):372–395, April 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7529115.

Littlewood:2013:CBP

[LP13a] B. Littlewood and A. Povyakalo.

Conservative bounds for the pfd of a 1-out-of-2 softwarebased system based on an assessor’s subjective probability of “Not Worse Than Independence”. IEEE Transactions on Software Engineering, 39 (12):1641–1653, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6564279.

Littlewood:2013:CRA

[LP13b] B. Littlewood and A. Povyakalo. Conservative reasoning about the probability of failure on demand of a 1-out-of-2 software-based system in which one channel is “Possibly Perfect”. IEEE Transactions on Software Engineering, 39(11):1521–1530, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6574864.

See [LR12].

Lu:2012:FAV

[LPZ12] S. Lu, S. Park, and Y. Zhou. Finding atomicity-violation bugs through unserializable interleaving testing. IEEE Transactions on Software Engineering, 38(4): 844–860, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5740930.

Littlewood:2012:RAR

[LR12] B. Littlewood and J. Rushby.

Reasoning about the reliability of diverse two-channel systems in which one channel is “Possibly Perfect”. IEEE Transactions on Software Engineering, 38(5):1178–1194, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

5975177. See comments [LP13b].

Lin:2014:AER

[LRCL14] Y. D. Lin, J. F. Rojas, E. T. H. Chu, and Y. C. Lai. On the accuracy, efficiency, and reusability of automated test oracles for Android devices. IEEE Transactions on Software Engineering, 40(10):957–970, October 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6840332.

Lee:2013:ESA

[LS13] J. Lee and I. Shin. EDZL schedulability analysis in real-time multicore scheduling. IEEE Transactions on Software Engineering, 39(7):910–916, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6374195.

Li:2018:FSV

[LSL^{+}18] Li Li, Jun Sun, Yang Liu, Meng Sun, and Jin-Song Dong. A formal specification and verification framework for timed security protocols. IEEE Transactions on Software Engineering, 44(8):725–746, August 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7939995/.

LeGoues:2012:MCQ

[LW12] C. Le Goues and W. Weimer. Measuring code quality to improve specification mining. IEEE Transactions on Software Engineering, 38(1): 175–190, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5680914.

Leue:2013:ILP

[LW13] S. Leue and W. Wei. Integer linear programmingbased property checking for asynchronous reactive systems. IEEE Transactions on Software Engineering, 39 (2):216–236, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5680910.

Lu:2019:ARO

[LWY^{+}19] Hong Lu, Shuai Wang, Tao Yue, Shaukat Ali, and Jan F. Nygård. Automated refactoring of OCL constraints with search. IEEE Transactions on Software Engineering, 45(2):148–170, February 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8114267/.

Liu:2014:GAD

[LXCL14] Y. Liu, C. Xu, S. C. Cheung, and J. Lü. GreenDroid: Automated diagnosis of energy inefficiency for Smartphone applications. IEEE Transactions on Software Engineering, 40(9):911–940, September 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6815752.

Lv:2014:ABA

[LYC14] J. Lv, B. B. Yin, and K. Y.

Cai. On the asymptotic behavior of adaptive testing strategy for software reliability assessment. IEEE Transactions on Software Engineering, 40(4):396–412, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6762895.

Liu:2011:PTV

[LYJC11] X. Liu, Y. Yang, Y. Jiang, and J. Chen. Preventing temporal violations in scientific workflows: Where and how. IEEE Transactions on Software Engineering, 37(6):805– 825, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5645643.

Liu:2012:PAA

[LZ12] P. Liu and C. Zhang. Pert: The application-aware tailoring of Java object persistence. IEEE Transactions on Software Engineering, 38(4):909–922, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5963692.

Li:2018:SSS

[LZRC18] Yi Li, Chenguang Zhu, Julia Rubin, and Marsha Chechik. Semantic slicing of software version histories. IEEE Transactions on Software Engineering, 44(2):182–201, February 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7843626/.

Lin:2010:REI

[LZX10] Z. Lin, X. Zhang, and D. Xu. Reverse engineering input syntactic structure from program execution and its applications. IEEE Transactions on Software Engineering, 36(5): 688–703, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5210120.

Mittas:2013:RCS

[MA13] N. Mittas and L. Angelis.

Ranking and clustering software cost estimation models through a multiple comparisons algorithm. IEEE Transactions on Software Engineering, 39(4):537–551, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6235961.

Mirarab:2012:SCR

[MAT12] S. Mirarab, S. Akhlaghi, and L. Tahvildari. Sizeconstrained regression test case selection using multicriteria optimization. IEEE Transactions on Software Engineering, 38(4):936–956, July/August 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5928351.

Misirli:2014:BNE

[MB14] A. T. Misirli and A. B. Bener. Bayesian networks for evidence-based decisionmaking in software engineering. IEEE Transactions on Software Engineering, 40(6):533–554, June 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6808495.

Menzies:2013:LVG

[MBC^{+}13] T. Menzies, A. Butcher, D. Cok, A. Marcus, L. Layman, F. Shull, B. Turhan, and T. Zimmermann. Local versus global lessons for defect prediction and effort estimation. IEEE Transactions on Software Engineering, 39(6):822–834, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6363444.

Moran:2020:MLB

[MBCC^{+}20] K. Moran, C. Bernal-Cárdenas, M. Curcio, R. Bonett, and D. Poshyvanyk. Machine learning-based prototyping of graphical user interfaces for mobile apps. IEEE Transactions on Software Engineering, 46(2):196–221, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Muslu:2015:RFD

[MBEN15] K. Muşlu, Y. Brun, M. D.

Ernst, and D. Notkin. Reducing feedback delay of software development tools via continuous analysis. IEEE Transactions on Software Engineering, 41(8):745–763, August 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7069257.

Moreau:2018:TSG

[MBH^{+}18] Luc Moreau, Belfrit Victor Batlajery, Trung Dong Huynh, Danius Michaelides, and Heather Packer. A templating system to generate provenance. IEEE Transactions on Software Engineering, 44(2):103–121, February 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/document/7909036/.

Menzies:2013:LPM

[MBK^{+}13] T. Menzies, A. Brady, J. Keung, J. Hihn, S. Williams, O. El-Rawas, P. Green, and B. Boehm. Learning project management decisions: A case study with case-based reasoning versus data farming. IEEE Transactions on Software Engineering, 39 (12):1698–1713, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6600685.

Mouchawrab:2011:ACC

[MBLD11] S. Mouchawrab, L. C. Briand, Y. Labiche, and M. Di Penta. Assessing, comparing, and combining state machinebased testing and structural testing: A series of experiments. IEEE Transactions on Software Engineering, 37(2):161–187, March/ April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5416729.

Meyer:2017:WLD

[MBM^{+}17] A. N. Meyer, L. E. Barton, G. C. Murphy, T. Zimmermann, and T. Fritz. The work life of developers: Activities, switches and perceived productivity. IEEE Transactions on Software Engineering, 43 (12):1178–1193, December 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7829407.

Moreno:2017:AAA

[MBP^{+}17] L. Moreno, G. Bavota, M. D.

Penta, R. Oliveto, A. Marcus, and G. Canfora. ARENA: An approach for the automated generation of release notes. IEEE Transactions on Software Engineering, 43(2):106–127, February 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7513412.

Meier:2010:EBM

[MC10] R. Meier and V. Cahill. On event-based middleware for location-aware mobile applications. IEEE Transactions on Software Engineering, 36(3):409–430, May/ June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5374426.

Mirakhorli:2016:DTM

[MCH16] M. Mirakhorli and J. Cleland-Huang. Detecting, tracing, and monitoring architectural tactics in code. IEEE Transactions on Software Engineering, 42(3):205–220, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7270338.

Moffett:2013:VPC

[MDB13] Y. Moffett, J. Dingel, and A. Beaulieu. Verifying protocol conformance using software model checking for the model-driven development of embedded systems. IEEE Transactions on Software Engineering, 39 (9):1307–13256, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6482140.

Medvidovic:2018:ENE

[Med18] N. Medvidović. Editorial from the new Editor in Chief. IEEE Transactions on Software Engineering, 44(1):3–4, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        8249710.

Medvidovic:2019:ESJ

[Med19] Nenad Medvidovic. Editorial: State of the journal. IEEE Transactions on Software Engineering, 45(1):1, January 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8605418/.

Mellegaard:2016:IID

[MFL^{+}16] N. Mellegård, A. Ferwerda, K. Lind, R. Heldal, and M. R. V. Chaudron. Impact of introducing domainspecific modelling in software maintenance: An industrial case study. IEEE Transactions on Software Engineering, 42(3):245–260, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7270333.

Martins:2020:RES

[MG20] L. E. G. Martins and T. Gorschek. Requirements engineering for safety-critical systems: An interview study with industry practitioners. IEEE Transactions on Software Engineering, 46(4):346– 361, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Mortensen:2012:AOR

[MGB12] M. Mortensen, S. Ghosh, and J. Bieman. Aspectoriented refactoring of legacy applications: An evaluation. IEEE Transactions on Software Engineering, 38(1): 118–140, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5661792.

Moha:2010:DMS

[MGDL10] N. Moha, Y. G. Guéhéneuc, L. Duchien, and A. F. Le Meur. DECOR: A method for the specification and detection of code and design smells. IEEE Transactions on Software Engineering, 36 (1):20–36, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5196681.

McMillan:2012:ESC

[MGP^{+}12] C. McMillan, M. Grechanik, D. Poshyvanyk, C. Fu, and Q. Xie. Exemplar: A source code search engine for finding highly relevant applications. IEEE Transactions on Software Engineering, 38(5):1069–1087, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5989838.

Mongiovi:2018:DOS

[MGS^{+}18] Melina Mongiovi, Rohit Gheyi, Gustavo Soares, Márcio Ribeiro, Paulo Borba, and Leopoldo Teixeira. Detecting overly strong preconditions in refactoring engines. IEEE Transactions on Software Engineering, 44(5):429–452, May 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7898404/.

Murphy-Hill:2012:PFR

[MHB12] E. Murphy-Hill and A. P.

Black. Programmer-friendly refactoring errors. IEEE Transactions on Software Engineering, 38(6):1417–1431, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6072219.

McMinn:2012:IDR

[MHL^{+}12] P. McMinn, M. Harman, K. Lakhotia, Y. Hassoun, and J. Wegener. Input domain reduction through irrelevant variable removal and its effect on local, global, and hybrid search-based structural test data generation. IEEE Transactions on Software Engineering, 38(2):453–477, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5710949.

Murphy-Hill:2012:HWR

[MHPB12] E. Murphy-Hill, C. Parnin, and A. P. Black. How we refactor, and how we know it. IEEE Transactions on Software Engineering, 38 (1):5–18, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6112738.

Monden:2013:ACE

[MHS^{+}13] A. Monden, T. Hayashi, S. Shinoda, K. Shirai, J. Yoshida, M. Barker, and K. Matsumoto. Assessing the cost effectiveness of fault prediction in acceptance testing. IEEE Transactions on Software Engineering, 39(10):1345–1357, October 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6497441.

Mei:2012:SAP

[MHZ^{+}12] H. Mei, D. Hao, L. Zhang, L. Zhang, J. Zhou, and G. Rothermel. A static approach to prioritizing JUnit test cases. IEEE Transactions on Software Engineering, 38(6):1258–1275, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6363461.

Murphy-Hill:2015:DSB

[MHZBN15] E. Murphy-Hill, T. Zimmermann, C. Bird, and N. Nagappan. The design space of bug fixes and how developers navigate it. IEEE Transactions on Software Engineering, 41(1):65–81, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6901259.

Marchetto:2016:MOT

[MIA^{+}16] A. Marchetto, M. M. Islam, W. Asghar, A. Susi, and G. Scanniello. A multiobjective technique to prioritize test cases. IEEE Transactions on Software Engineering, 42(10):918–940, October 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7362042.

Macedo:2017:FBC

[MJC17] N. Macedo, T. Jorge, and A. Cunha. A feature-based classification of model repair approaches. IEEE Transactions on Software Engineering, 43(7):615–640, July 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7605502.

McBurney:2018:TPD

[MJK^{+}18] Paul W. McBurney, Siyuan Jiang, Marouane Kessentini, Nicholas A. Kraft, Ameer Armaly, Mohamed Wiem Mkaouer, and Collin McMillan. Towards prioritizing documentation effort. IEEE Transactions on Software Engineering, 44(9):

897–913, September 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7953505/.

McIntosh:2018:FIC

[MK18] Shane McIntosh and Yasutaka Kamei. Are fixinducing changes a moving target? A longitudinal case study of just-in-time defect prediction. IEEE Transactions on Software Engineering, 44(5):412–428, May 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7898457/.

Malavolta:2013:WIN

[MLM^{+}13] I. Malavolta, P. Lago, H. Muccini, P. Pelliccione, and A. Tang. What industry needs from architectural languages: A survey. IEEE Transactions on Software Engineering, 39(6):869–891, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6374194.

Mangano:2015:HSD

[MLPvdH15] N. Mangano, T. D. LaToza, M. Petre, and A. van der Hoek. How software designers interact with sketches at the whiteboard. IEEE Transactions on Software Engineering, 41(2):135–156, February 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6922572.

Miller:2014:RES

[MLS^{+}14] T. Miller, B. Lu, L. Sterling, G. Beydoun, and K. Taveter. Requirements elicitation and specification using the agent paradigm: The case study of an aircraft turnaround simulator. IEEE Transactions on Software Engineering, 40(10):1007–1024, October 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6860260.

Milicev:2013:CBU

[MM13] D. Milicev and Z. Mijailovic.

Capsule-based user interface modeling for large-scale applications. IEEE Transactions on Software Engineering, 39(9):1190–1207, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6464270.

McBurney:2016:ASC

[MM16] P. W. McBurney and C. McMillan. Automatic source code summarization of context for Java methods. IEEE Transactions on Software Engineering, 42(2):103–119, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7181703.

Malek:2012:EFI

[MMMR12] S. Malek, N. Medvidovic, and M. Mikic-Rakic. An extensible framework for improving a distributed software system’s deployment architecture. IEEE Transactions on Software Engineering, 38(1): 73–100, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5680912.

Mirshokraie:2015:GMT

[MMP15] S. Mirshokraie, A. Mesbah, and K. Pattabiraman. Guided mutation testing for JavaScript Web applications. IEEE Transactions on Software Engineering, 41(5):429–444, May 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6960094.

Malavolta:2010:PAL

[MMPT10] I. Malavolta, H. Muccini, P. Pelliccione, and D. Tamburri. Providing architectural languages and tools interoperability through model transformation technologies. IEEE Transactions on Software Engineering, 36(1): 119–140, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5204094.

Margheri:2019:RFS

[MMPT19] Andrea Margheri, Massimiliano Masi, Rosario Pugliese, and Francesco Tiezzi. A rigorous framework for specification, analysis and enforcement of access control policies. IEEE Transactions on Software Engineering, 45(1):2–33, January 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8081817/.

Massacci:2014:EME

[MN14] F. Massacci and V. H.

Nguyen. An empirical methodology to evaluate vulnerability discovery models. IEEE Transactions on Software Engineering, 40 (12):1147–1162, December 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6891367.

Mertz:2017:QSA

[MN17] J. Mertz and I. Nunes. A qualitative study of applicationlevel caching. IEEE Transactions on Software Engineering, 43(9):798–816, September 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7762909.

Matinnejad:2019:TGT

[MNBB19] R. Matinnejad, S. Nejati, L. C. Briand, and T. Bruckmann. Test generation and test prioritization for Simulink models with dynamic behavior. IEEE Transactions on Software Engineering, 45(9):919–944, September 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Massey:2015:ELI

[MOA15] A. K. Massey, P. N. Otto, and A. I. Antón. Evaluating legal implementation readiness decision-making. IEEE Transactions on Software Engineering, 41(6):545–564, June 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6991569.

Mesgari:2019:CRR

[MOd19] M. Mesgari, C. Okoli, and A. O. de Guinea. Creating rich and representative personas by discovering affordances. IEEE Transactions on Software Engineering, 45 (10):967–983, October 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Madeyski:2014:OEM

[MOTJ14] L. Madeyski, W. Orzeszyna, R. Torkar, and M. Józala. Overcoming the equivalent mutant problem: A systematic literature review and a comparative experiment of second order mutation. IEEE Transactions on Software Engineering, 40(1):23–42, January 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6613487.

Mottola:2019:MSI

[MPO^{+}19] L. Mottola, G. P. Picco, F. J. Oppermann, J. Eriksson, N. Finne, H. Fuchs, A. Gaglione, S. Karnouskos, P. M. Montero, N. Oertel, K. Römer, P. Spieß, S. Tranquillini, and T. Voigt. make-Sense: Simplifying the integration of wireless sensor networks into business processes. IEEE Transactions on Software Engineering, 45 (6):576–596, June 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Mariani:2011:DAD

[MPP11] L. Mariani, F. Pastore, and M. Pezze. Dynamic analysis for diagnosing integration faults. IEEE Transactions on Software Engineering, 37(4):486–508, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5611554.

Mateescu:2012:ASP

[MPS12] R. Mateescu, P. Poizat, and G. Salaün. Adaptation of service protocols using process algebra and on-the-fly reduction techniques. IEEE Transactions on Software Engineering, 38(4):755–777, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5928357.

Mariani:2017:GTE

[MPS17] L. Mariani, M. Pezzè, and M. Santoro. GK-Tail+: An efficient approach to learn software models. IEEE Transactions on Software Engineering, 43(8):715–738, August 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7728088.

Maalej:2013:PKA

[MR13] W. Maalej and M. P. Robillard. Patterns of knowledge in API reference documentation. IEEE Transactions on Software Engineering, 39(9):1264–1282, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6473801.

Moser:2012:DSS

[MRD12] O. Moser, F. Rosenberg, and S. Dustdar. Domain-specific service selection for composite services. IEEE Transactions on Software Engineering, 38(4):828–843, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6231591.

Medeiros:2018:DMR

[MRG^{+}18] Flávio Medeiros, Márcio Ribeiro, Rohit Gheyi, Sven Apel, Christian Kästner, Bruno Ferreira, Luiz Carvalho, and Baldoino Fonseca. Discipline matters: Refactoring of preprocessor directives in the #ifdef hell. IEEE Transactions on Software Engineering, 44(5):453–469, May 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7888579/.

Murphy:2010:GEI

[MS10] G. C. Murphy and W. Schafer. Guest Editors’ introduction: 2008 conference on the foundations of software engineering. IEEE Transactions on Software Engineering, 36(5): 591–592, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5593044.

Margara:2018:SDR

[MS18] Alessandro Margara and Guido Salvaneschi. On the semantics of distributed reactive programming: The cost of consistency. IEEE Transactions on Software Engineering, 44(7):689–711, July 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8354906/.

Maras:2013:ICI

[MSCC13] J. Maras, M. Stula, J. Carlson, and I. Crnkovic. Identifying code of individual features in client-side Web applications. IEEE Transactions on Software Engineering, 39 (12):1680–1697, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6576749.

Martin:2017:SAS

[MSJ^{+}17] W. Martin, F. Sarro, Y. Jia, Y. Zhang, and M. Harman. A survey of app store analysis for software engineering. IEEE Transactions on Software Engineering, 43(9):817–847, September 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7765038.

Morales:2018:EEA

[MSK^{+}18] Rodrigo Morales, Rubén Saborido, Foutse Khomh, Francisco Chicano, and Giuliano Antoniol. EARMO: An energy-aware refactoring approach for mobile apps. IEEE Transactions on Software Engineering, 44 (12):1176–1206, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8052533/.

MacDonell:2010:HRS

[MSKM10] S. MacDonell, M. Shepperd, B. Kitchenham, and E. Mendes. How reliable are systematic reviews in empirical software engineering? IEEE Transactions on Software Engineering, 36(5): 676–687, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5416726.

Marchetto:2019:CCR

[MSS19] A. Marchetto, G. Scanniello, and A. Susi. Combining code and requirements coverage with execution cost for test suite reduction. IEEE Transactions on Software Engineering, 45(4):363–390, April 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Marron:2013:ARH

[MSSF13] M. Marron, C. Sanchez, Z. Su, and M. Fahndrich. Abstracting runtime heaps for program understanding. IEEE Transactions on Software Engineering, 39(6):774–786, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6331492.

Minku:2014:IEA

[MSY14] L. L. Minku, D. Sudholt, and X. Yao. Improved evolutionary algorithm design for the project scheduling problem based on runtime analysis. IEEE Transactions on Software Engineering, 40(1):83–102, January 2014. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6648326.

Maplesden:2015:PAO

[MTHG15] D. Maplesden, E. Tempero, J. Hosking, and J. C. Grundy. Performance analysis for object-oriented software: A systematic mapping. IEEE Transactions on Software Engineering, 41(7):691–710, July 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7024167.

Mateo:2013:VSO

[MUA13] P. Reales Mateo, M. Polo Usaola, and J. L. Fernández Alemán. Validating secondorder mutation at system level. IEEE Transactions on Software Engineering, 39(4):570–587, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6216382.

Mesbah:2012:IBA

[MvDR12] A. Mesbah, A. van Deursen, and D. Roest. Invariantbased automatic testing of modern Web applications. IEEE Transactions on Software Engineering, 38(1):35– 53, January/February 2012.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5728834.

Meyers:2020:FTV

[MVDS20] B. Meyers, H. Vangheluwe, J. Denil, and R. Salay. A framework for temporal verification support in domainspecific modelling. IEEE Transactions on Software Engineering, 46(4):362–404, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Manadhata:2011:ASM

[MW11] P. K. Manadhata and J. M. Wing. An attack surface metric. IEEE Transactions on Software Engineering, 37(3):371–386, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5482589.

McMinn:2019:ADR

[MWMK19] P. McMinn, C. J. Wright, C. J. McCurdy, and G. M. Kapfhammer. Automatic detection and removal of ineffective mutants for the mutation analysis of relational database schemas. IEEE Transactions on Software Engineering, 45 (5):427–463, May 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Mercan:2019:CCH

[MYK19] H. Mercan, C. Yilmaz, and K. Kaya. CHiP: A configurable hybrid parallel covering array constructor. IEEE Transactions on Software Engineering, 45(12):1270– 1291, December 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Nallur:2013:DSA

[NB13] V. Nallur and R. Bahsoon. A decentralized selfadaptation mechanism for service-based applications in the cloud. IEEE Transactions on Software Engineering, 39(5):591–612, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6249687.

Nadi:2015:WDC

[NBKC15] S. Nadi, T. Berger, C. Kästner, and K. Czarnecki. Where do configuration constraints stem from? An extraction approach and an empirical study. IEEE Transactions on Software Engineering, 41(8):820–841, August 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?arnumber=
                                        7065312.

Natella:2013:FRS

[NCDM13] R. Natella, D. Cotroneo, J. A. Duraes, and H. S. Madeira. On fault representativeness of software fault injection. IEEE Transactions on Software Engineering, 39(1):80–96, January 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6122035.

Namoun:2016:EME

[NDMX16] A. Namoun, A. Daskalopoulou, N. Mehandjiev, and Z. Xun. Exploring mobile end user development: Existing use and design factors. IEEE Transactions on Software Engineering, 42(10):960–976, October 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7416245.

Nam:2018:HDP

[NFK^{+}18] Jaechang Nam, Wei Fu, Sunghun Kim, Tim Menzies, and Lin Tan. Heterogeneous defect prediction. IEEE Transactions on Software Engineering, 44 (9):874–896, September 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7959597/.

Nguyen:2014:OME

[NM14] B. N. Nguyen and A. M.

Memon. An observemodel-exercise* paradigm to test event-driven systems with undetermined input spaces. IEEE Transactions on Software Engineering, 40(3):216–234, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6714448.

Niu:2020:IFC

[NNL^{+}20] X. Niu, C. Nie, J. Y. Lei, H. Leung, and X. Wang. Identifying failure-causing schemas in the presence of multiple faults. IEEE Transactions on Software Engineering, 46(2):141–162, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Nguyen:2012:CME

[NNP^{+}12] H. A. Nguyen, T. T. Nguyen, N. H. Pham, J. Al-Kofahi, and T. N. Nguyen. Clone management for evolving software. IEEE Transactions on Software Engineering, 38(5):1008–1026, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6007141.

Nayebi:2019:ARP

[NR19] M. Nayebi and G. Ruhe. Asymmetric release planning: Compromising satisfaction against dissatisfaction. IEEE Transactions on Software Engineering, 45(9):839– 857, September 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Nongpoh:2017:AFA

[NRDB17] B. Nongpoh, R. Ray, S. Dutta, and A. Banerjee. AutoSense: A framework for automated sensitivity analysis of program data. IEEE Transactions on Software Engineering, 43(12):1110– 1124, December 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7820185.

Nejati:2012:MMV

[NSC^{+}12] S. Nejati, M. Sabetzadeh, M. Chechik, S. Easterbrook, and P. Zave. Matching and merging of variant feature specifications. IEEE Transactions on Software Engineering, 38(6):1355–1375, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6086550.

Nuseibeh:2010:E

[Nus10a]  B. Nuseibeh.  Editorial.

IEEE Transactions on Software Engineering, 36(6):735– 736, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5644734.

Nuseibeh:2010:END

[Nus10b]  B. Nuseibeh.  Editorial:

A new decade of TSE. IEEE Transactions on Software Engineering, 36(1):3– 6, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5401365.

Nuseibeh:2010:EHS

[Nus10c]  B. Nuseibeh.  Editorial:

How special should issues be? IEEE Transactions on Software Engineering, 36(4):449–450, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5532337.

Nuseibeh:2010:ERW

[Nus10d] B. Nuseibeh. Editorial: Readers, writers, reviewers, and editors. IEEE Transactions on Software Engineering, 36(2):145–156, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5439568.

Nuseibeh:2010:MRM

[Nus10e]  B. Nuseibeh.  In memo-

riam: Robin Milner and Amir Pnueli. IEEE Transactions on Software Engineering, 36(3):305, May/June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5473899.

Nuseibeh:2011:ESJ

[Nus11a]  B. Nuseibeh.  Editorial:

State of the journal. IEEE Transactions on Software Engineering, 37(1):1–3, January/February 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5704236.

Nuseibeh:2011:EWM

[Nus11b]  B. Nuseibeh.  Editorial:

What makes a publication archival? IEEE Transactions on Software Engineering, 37(2):145, March/April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5739157.

Nuseibeh:2012:SJ

[Nus12] B. Nuseibeh. State of the journal. IEEE Transactions on Software Engineering, 38 (1):1–2, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6141070.

Nuseibeh:2013:E

[Nus13a]  B. Nuseibeh.  Editorial.

IEEE Transactions on Software Engineering, 39(9): 1187–1189, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6587460.

Nuseibeh:2013:ENA

[Nus13b] B. Nuseibeh. Editorial [new associate editors]. IEEE Transactions on Software Engineering, 39(5):588–590, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6509892.

Nuseibeh:2013:ESJ

[Nus13c] B. Nuseibeh. Editorial: State of the journal. IEEE Transactions on Software Engineering, 39(2):145–146, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6419732.

Nuseibeh:2013:MDN

[Nus13d]  B. Nuseibeh.  In memo-

riam — David Notkin (1953– 2013). IEEE Transactions on Software Engineering, 39(6):742–743, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6519247.

Nuseibeh:2013:MMJ

[Nus13e]  B. Nuseibeh.  In memo-

riam: Mary Jean Harrold (1947–2013). IEEE Transactions on Software Engineering, 39(11):1466, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6648565.

Nuseibeh:2014:SSJ

[Nus14] B. Nuseibeh. Signing off: The state of the journal. IEEE Transactions on Software Engineering, 40(1):1–3, January 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6757048.

Natella:2020:AEB

[NWCS20] R. Natella, S. Winter, D. Cotroneo, and N. Suri. Analyzing the effects of bugs on software interfaces. IEEE Transactions on Software Engineering, 46(3):280–301, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Noureddine:2016:MCS

[NZ16] M. A. Noureddine and F. A.

Zaraket. Model checking software with first order logic specifications using AIG solvers. IEEE Transactions on Software Engineering, 42(8):741–763, August 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7389426.

Ocariza:2017:SCC

[OBPM17] F. S. Ocariza, K. Bajaj, K. Pattabiraman, and A. Mesbah. A study of causes and consequences of client-side JavaScript bugs. IEEE Transactions on Software Engineering, 43(2): 128–144, February 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7501855.

Oliveira:2013:AMA

[OF13] J. N. Oliveira and M. A. Ferreira. Alloy meets the algebra of programming: A case study. IEEE Transactions on Software Engineering, 39(3):305–326, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6155724.

Octeau:2016:CCP

[OLJM16] D. Octeau, D. Luchaup, S. Jha, and P. McDaniel. Composite constant propagation and its application to Android program analysis. IEEE Transactions on Software Engineering, 42(11):999–1014, November 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7447806.

Orso:2012:GES

[OT12] A. Orso and P. Tonella. Guest editorial: Special section on the international symposium on software testing and analysis 2010. IEEE Transactions on Software Engineering, 38(2):241–242, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6173077.

Ozcelik:2016:SLO

[OY16] B. Ozcelik and C. Yilmaz.

Seer: A lightweight online failure prediction approach. IEEE Transactions on Software Engineering, 42(1):26–46, January 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7120143.

Ponzanelli:2019:AIC

[PBM^{+}19] L. Ponzanelli, G. Bavota, A. Mocci, R. Oliveto, M. D. Penta, S. Haiduc, B. Russo, and M. Lanza. Automatic identification and classification of software development video tutorial fragments. IEEE Transactions on Software Engineering, 45 (5):464–488, May 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Palomba:2015:MVH

[PBP^{+}15] F. Palomba, G. Bavota, M. D.

Penta, R. Oliveto, D. Poshyvanyk, and A. De Lucia. Mining version histories for detecting code smells. IEEE Transactions on Software Engineering, 41(5):462–489, May 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6963448.

Petersen:2018:CCO

[PBS^{+}18] Kai Petersen, Deepika Badampudi, Syed Muhammad Ali Shah, Krzysztof Wnuk, Tony Gorschek, Efi Papatheocharous, Jakob Axelsson, Séverine Sentilles, Ivica Crnkovic, and Antonio Cicchetti. Choosing component origins for software intensive systems: In-house, COTS, OSS or outsourcing?

— A case survey. IEEE Transactions on Software Engineering, 44(3):237–261, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7870688/.

Pavese:2016:PIA

[PBU16] E. Pavese, V. Braberman, and S. Uchitel. Probabilistic interface automata. IEEE Transactions on Software Engineering, 42(9):843–865, September 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7401103.

Poulding:2010:ESV

[PC10] S. Poulding and J. A. Clark. Efficient software verification: Statistical testing using automated search. IEEE Transactions on Software Engineering, 36(6):763– 777, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5406530.

Petke:2015:PCI

[PCHY15] J. Petke, M. B. Cohen, M. Harman, and S. Yoo. Practical combinatorial interaction testing: Empirical findings on efficiency and early fault detection. IEEE Transactions on Software Engineering, 41(9): 901–924, September 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7081752.

Perez:2015:ECR

[PCPS15] J. F. Pérez, G. Casale, and S. Pacheco-Sanchez. Estimating computational requirements in multi-threaded applications. IEEE Transactions on Software Engineering, 41(3):264–278, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6926798.

Pei:2014:AFP

[PFN^{+}14] Y. Pei, C. A. Furia, M. Nordio, Y. Wei, B. Meyer, and A. Zeller. Automated fixing of programs with contracts. IEEE Transactions on Software Engineering, 40(5):427–449, May 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6776507.

Pradel:2013:NBA

[PG13] M. Pradel and T. R. Gross. Name-based analysis of equally typed method arguments. IEEE Transactions on Software Engineering, 39(8): 1127–1143, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6419711.

Petke:2018:SSD

[PHLW18] Justyna Petke, Mark Harman, William B. Langdon, and Westley Weimer. Specialising software for different downstream applications using genetic improvement and code transplantation. IEEE Transactions on Software Engineering, 44(6):574–594, June 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7962212/.

Paolieri:2016:PMC

[PHV16] M. Paolieri, A. Horváth, and E. Vicario. Probabilistic model checking of regenerative concurrent systems. IEEE Transactions on Software Engineering, 42 (2):153–169, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7202875.

Praditwong:2011:SMC

[PHY11] K. Praditwong, M. Harman, and X. Yao. Software module clustering as a multi-objective search problem. IEEE Transactions on Software Engineering, 37(2):264–282, March/ April 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5406532.

Panichella:2018:ATC

[PKT18] Annibale Panichella, Fitsum Meshesha Kifetew, and Paolo Tonella. Automated test case generation as a many-objective optimisation problem with dynamic selection of the targets. IEEE Transactions on Software Engineering, 44 (2):122–158, February 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/7840029/.

Pittke:2015:ADR

[PLM15] F. Pittke, H. Leopold, and J. Mendling. Automatic detection and resolution of lexical ambiguity in process models. IEEE Transactions on Software Engineering, 41(6):526–544, June 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7027184.

Pettersson:2010:EAD

[PLN10] N. Pettersson, W. Lowe, and J. Nivre. Evaluation of accuracy in design pattern occurrence detection. IEEE Transactions on Software Engineering, 36(4):575–590, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5374428.

Palma:2019:UUS

[PMG19] F. Palma, N. Moha, and Y. Guéhéneuc. UniDoSA: The unified specification and detection of service antipatterns. IEEE Transactions on Software Engineering, 45(10): 1024–1053, October 2019.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Peters:2013:BPU

[PMGZ13] F. Peters, T. Menzies, L. Gong, and H. Zhang. Balancing privacy and utility in cross-company defect prediction. IEEE Transactions on Software Engineering, 39(8):1054–1068, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6419712.

Palyart:2018:SSI

[PMM18] Marc Palyart, Gail C. Murphy, and Vaden Masrani. A study of social interactions in open source component use. IEEE Transactions on Software Engineering, 44 (12):1132–1145, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8049385/.

Pauleve:2011:TTF

[PMR11] L. Pauleve, M. Magnin, and O. Roux. Tuning temporal features within the stochastic π-calculus. IEEE Transactions on Software Engineering, 37(6):858–871, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5611556.

Piccioni:2013:CSE

[POM13] M. Piccioni, M. Oriol, and B. Meyer. Class schema evolution for persistent objectoriented software: Model, empirical study, and automated support. IEEE Transactions on Software Engineering, 39(2):184–196, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6122034.

Panichella:2015:IMO

[POPD15] A. Panichella, R. Oliveto, M. D. Penta, and A. De Lucia. Improving multiobjective test case selection by injecting diversity in genetic algorithms. IEEE Transactions on Software Engineering, 41(4):358–383, April 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6936894.

Palomba:2018:SSE

[PPZ^{+}18] Fabio Palomba, Annibale Panichella, Andy Zaidman, Rocco Oliveto, and Andrea De Lucia. The scent of a smell: An extensive comparison between textual and structural smells. IEEE Transactions on Software Engineering, 44(10):977–1000, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8038053/.

Perepletchikov:2011:CEE

[PR11] M. Perepletchikov and C. Ryan.

A controlled experiment for evaluating the impact of coupling on the maintainability of service-oriented software. IEEE Transactions on Software Engineering, 37(4):449–465, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5482590.

Prechelt:2011:PWD

[Pre11] L. Prechelt. Plat Forms: A Web development platform comparison by an exploratory experiment searching for emergent platform properties. IEEE Transactions on Software Engineering, 37(1): 95–108, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5406528.

Pietrantuono:2010:SRT

[PRT10] R. Pietrantuono, S. Russo, and K. S. Trivedi. Software reliability and testing time allocation: An architecturebased approach. IEEE Transactions on Software Engineering, 36(3):323–337, May/ June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5383374.

Perini:2013:MLA

[PSA13] A. Perini, A. Susi, and P. Avesani. A machine learning approach to software requirements prioritization. IEEE Transactions on Software Engineering, 39(4):445–461, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6249686.

Peitek:2020:LPH

[PSA^{+}20] N. Peitek, J. Siegmund, S. Apel, C. Kästner, C. Parnin, A. Bethmann, T. Leich, G. Saake, and A. Brechmann. A look into programmers heads. IEEE Transactions on Software Engineering, 46(4):442–462, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Platenius:2017:IMR

[PSB^{+}17] M. C. Platenius, A. Shaker, M. Becker, E. Hüllermeier, and W. Schäfer. Imprecise matching of requirements specifications for software services using fuzzy logic. IEEE Transactions on Software Engineering, 43(8):739–759, August 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7755807.

Peters:2019:TFR

[PTYN19] F. Peters, T. T. Tun, Y. Yu, and B. Nuseibeh. Text filtering and ranking for security bug report prediction. IEEE Transactions on Software Engineering, 45(6): 615–631, June 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Parmer:2012:MPD

[PW12] G. Parmer and R. West.

Mutable protection domains: Adapting system fault isolation for reliability and efficiency. IEEE Transactions on Software Engineering, 38(4):875–888, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5928356.

Pan:2019:GCF

[PWSG19] Yao Pan, Jules White, Yu Sun, and Jeff Gray. Gray computing: A framework for computing with background JavaScript tasks.

IEEE Transactions on Software Engineering, 45(2): 171–193, February 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8105894/.

Palomba:2019:TSA

[PZF^{+}19] Fabio Palomba, Marco Zanoni, Francesca Arcelli Fontana, Andrea De Lucia, and Rocco Oliveto. Toward a smell-aware bug prediction model. IEEE Transactions on Software Engineering, 45(2):194–218, February 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8097044/.

Padgham:2013:MBT

[PZTM13] L. Padgham, Z. Zhang, J. Thangarajah, and T. Miller. Model-based test oracle generation for automated unit testing of agent systems. IEEE Transactions on Software Engineering, 39(9): 1230–1244, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6464272.

Qiu:2016:URE

[QSM16] J. Qiu, X. Su, and P. Ma. Using reduced execution flow graph to identify library functions in binary code. IEEE Transactions on Software Engineering, 42 (2):187–202, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7210204.

Ralph:2019:TMG

[Ral19] P. Ralph. Toward methodological guidelines for process theories and taxonomies in software engineering. IEEE Transactions on Software Engineering, 45(7):712–735, July 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Ramasubbu:2014:GSP

[Ram14]  N. Ramasubbu.  Govern-

ing software process improvements in globally distributed product development. IEEE Transactions on Software Engineering, 40(3):235–250, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6682900.

Rakha:2018:RPE

[RBH18] Mohamed Sami Rakha, Cor-Paul Bezemer, and Ahmed E. Hassan. Revisiting the performance evaluation of automated approaches for the retrieval of duplicate issue reports. IEEE Transactions on Software Engineering, 44 (12):1245–1268, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8048025/.

Robillard:2013:AAP

[RBK^{+}13] M. P. Robillard, E. Bodden, D. Kawrykow, M. Mezini, and T. Ratchford. Automated API property inference techniques. IEEE Transactions on Software Engineering, 39(5):613–637, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6311409.

Riehle:2016:ISP

[RCKH16] D. Riehle, M. Capraro, D. Kips, and L. Horn. Inner source in platformbased product engineering. IEEE Transactions on Software Engineering, 42(12): 1162–1177, December 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7452676.

Ricca:2010:HDE

[RDT^{+}10] F. Ricca, M. Di Penta, M. Torchiano, P. Tonella, and M. Ceccato. How developers’ experience and ability influence Web application comprehension tasks supported by UML stereotypes: A series of four experiments. IEEE Transactions on Software Engineering, 36(1): 96–118, January/February 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5332231.

Reder:2013:DCD

[RE13] A. Reder and A. Egyed.

Determining the cause of a design model inconsistency. IEEE Transactions on Software Engineering, 39 (11):1531–1548, November 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6560054.

Rosner:2015:BIS

[RGA^{+}15] N. Rosner, J. Geldenhuys, N. M. Aguirre, W. Visser, and M. F. Frias. BLISS: Improved symbolic execution by bounded lazy initialization with SAT support. IEEE Transactions on Software Engineering, 41(7):639–660, July 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?arnumber=
                                        7004061.

Rothlisberger:2012:EDI

[RHB^{+}12] D. Rothlisberger, M. Harry, W. Binder, P. Moret, D. Ansaloni, A. Villazon, and O. Nierstrasz. Exploiting dynamic information in IDEs improves speed and correctness of software maintenance tasks. IEEE Transactions on Software Engineering, 38(3):579–591, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6178187.

Ramasubbu:2014:MTD

[RK14] N. Ramasubbu and C. F. Kemerer. Managing technical debt in enterprise software packages. IEEE Transactions on Software Engineering, 40(8):758–772, August 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6824267.

Ramasubbu:2019:ITD

[RK19] Narayan Ramasubbu and Chris F. Kemerer. Integrating technical debt management and software quality management processes: A normative framework and field tests. IEEE Transactions on Software Engineering, 45(3):285–300, March 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8114229/.

Ramasubbu:2012:SCP

[RKH12] N. Ramasubbu, C. F. Kemerer, and J. Hong. Structural complexity and programmer team strategy: An experimental test. IEEE Transactions on Software Engineering, 38(5):1054–1068, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5999673.

Rodeghero:2015:ETS

[RLMM15] P. Rodeghero, C. Liu, P. W.

McBurney, and C. McMillan. An eye-tracking study of Java programmers and application to source code summarization. IEEE Transactions on Software Engineering, 41 (11):1038–1054, November 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7118751.

Rodriguez:2014:GTT

[RLR14] I. Rodríguez, L. Llana, and P. Rabanal. A general testability theory: Classes, properties, complexity, and testing reductions. IEEE Transactions on Software Engineering, 40(9):862–894, September 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6839051.

Rafique:2013:ETD

[RM13] Y. Rafique and V. B. Mišić. The effects of test-driven development on external quality and productivity: A meta-analysis. IEEE Transactions on Software Engineering, 39(6):835–856, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6197200.

Rempel:2017:PDI

[RM17] P. Rempel and P. Mäder. Preventing defects: The impact of requirements traceability completeness on software quality. IEEE Transactions on Software Engineering, 43(8):777–797, August 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7723818.

Rastkar:2014:ASB

[RMM14] S. Rastkar, G. C. Murphy, and G. Murray. Automatic summarization of bug reports. IEEE Transactions on Software Engineering, 40(4):366–380, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6704866.

Rodriguez-Navas:2013:UTA

[RNP13] G. Rodriguez-Navas and J. Proenza. Using timed automata for modeling distributed systems with clocks: Challenges and solutions. IEEE Transactions on Software Engineering, 39(6):857– 868, June 2013. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6374193.

Raunak:2013:RMC

[RO13] M. S. Raunak and L. J. Osterweil. Resource management for complex, dynamic environments. IEEE Transactions on Software Engineering, 39(3):384–402, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6197203.

Rosa:2013:SMA

[RRL^{+}13] L. Rosa, L. Rodrigues, A. Lopes, M. Hiltunen, and R. Schlichting. Selfmanagement of adaptable component-based applications. IEEE Transactions on Software Engineering, 39 (3):403–421, March 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6197201.

Ramirez:2019:SRI

[RRS19] A. Ramírez, J. R. Romero, and C. L. Simons. A systematic review of interaction in search-based software engineering. IEEE Transactions on Software Engineering, 45(8):760–781, August 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Raja:2012:DEM

[RT12] U. Raja and M. J. Tretter. Defining and evaluating a measure of open source project survivability. IEEE Transactions on Software Engineering, 38(1): 163–174, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6127835.

Ryder:2010:IBP

[RZ10] B. G. Ryder and A. Zeller. Introduction: The best papers of ISSTA. IEEE Transactions on Software Engineering, 36(4):451–452, July/August 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5532338.

Sadeghi:2017:TQC

[SBGM17] A. Sadeghi, H. Bagheri, J. Garcia, and S. Malek. A taxonomy and qualitative comparison of program analysis techniques for security assessment of Android software. IEEE Transactions on Software Engineering, 43(6):492–530, June 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7583740.

Shepperd:2014:RBU

[SBH14] M. Shepperd, D. Bowes, and T. Hall. Researcher bias: The use of machine learning in software defect prediction. IEEE Transactions on Software Engineering, 40(6):603–616, June 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

6824804. See [TMHM16, SHB18].

Shousha:2012:UMM

[SBL12] M. Shousha, L. Briand, and Y. Labiche. A UML/MARTE model analysis method for uncovering scenarios leading to starvation and deadlocks in concurrent systems. IEEE Transactions on Software Engineering, 38(2):354–374, March/April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5661791.

Sampath:2013:URH

[SBM13] S. Sampath, R. Bryce, and A. M. Memon. A uniform representation of hybrid criteria for regression testing. IEEE Transactions on Software Engineering, 39(10):1326–1344, October 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6484067.

Scalabrino:2019:LCR

[SBR^{+}19] Simone Scalabrino, Gabriele Bavota, Barbara Russo, Massimiliano Di Penta, and Rocco Oliveto. Listening to the crowd for the release planning of mobile apps. IEEE Transactions on Software Engineering, 45(1):68–86, January 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8057860/.

Sibay:2013:SMT

[SBUK13] G. E. Sibay, V. Braberman, S. Uchitel, and J. Kramer. Synthesizing modal transition systems from triggered scenarios. IEEE Transactions on Software Engineering, 39(7):975–1001, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311408.

Shevtsov:2018:CTS

[SBWM18] Stepan Shevtsov, Mihaly Berekmeri, Danny Weyns, and Martina Maggio. Controltheoretical software adaptation: A systematic literature review. IEEE Transactions on Software Engineering, 44(8):784–810, August 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7929422/.

Schreiber:2012:PLM

[SCF^{+}12] F. A. Schreiber, R. Camplani, M. Fortunato, M. Marelli, and G. Rota. PerLa: A language and middleware architecture for data management and integration in pervasive information systems. IEEE Transactions on Software Engineering, 38(2):478–496, March/April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5728831.

Stol:2019:CBC

[SCF19] Klaas-Jan Stol, Bora Caglayan, and Brian Fitzgerald. Competitionbased crowdsourcing software development: a multi-method study from a customer perspective. IEEE Transactions on Software Engineering, 45(3):237–260, March 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8119867/.

Song:2014:STN

[SCP14] J. Song, C. Cadar, and P. Pietzuch. SymbexNet: Testing network protocol implementations with symbolic execution and rule-based specifications. IEEE Transactions on Software Engineering, 40(7):695–709, July 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6815719.

Sharp:2016:RES

[SDdS16] H. Sharp, Y. Dittrich, and C. R. B. de Souza. The role of ethnographic studies in empirical software engineering. IEEE Transactions on Software Engineering, 42(8):786–804, August 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7387744.

Schumann:2014:MHL

[SDMW14] M. A. Schumann, D. Drusinsky, J. B. Michael, and D. Wijesekera. Modeling human-in-the-loop security analysis and decision-making processes. IEEE Transactions on Software Engineering, 40(2):154–166, February 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6727512.

Stolee:2013:IIR

[SE13] K. T. Stolee and S. Elbaum. Identification, impact, and refactoring of smells in pipe-like Web mashups. IEEE Transactions on Software Engineering, 39(12): 1654–1679, December 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6589568.

Sama:2010:CAA

[SER^{+}10] M. Sama, S. Elbaum, F. Raimondi, D. S. Rosenblum, and Z. Wang. Contextaware adaptive applications: Fault patterns and their automated identification. IEEE Transactions on Software Engineering, 36(5):644–661, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5432224.

Su:2016:APB

[SFCR16] G. Su, Y. Feng, T. Chen, and D. S. Rosenblum. Asymptotic perturbation bounds for probabilistic model checking with empirically determined probability parameters. IEEE Transactions on Software Engineering, 42(7):623–639, July 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7355393.

Sarro:2017:AMO

[SFH^{+}17] F. Sarro, F. Ferrucci, M. Harman, A. Manna, and J. Ren. Adaptive multi-objective evolutionary algorithms for overtime planning in software projects. IEEE Transactions on Software Engineering, 43(10):898–917, October 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7814340.

Segura:2016:SMT

[SFSRC16] S. Segura, G. Fraser, A. B.

Sanchez, and A. Ruiz-Cortés. A survey on metamorphic testing. IEEE Transactions on Software Engineering, 42(9):805–824, September 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7422146.

Shah:2010:UEH

[SGH10] H. Shah, C. Gorg, and M. J. Harrold. Understanding exception handling: Viewpoints of novices and experts. IEEE Transactions on Software Engineering, 36(2):150–161, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5383375.

Soares:2013:ABT

[SGM13] G. Soares, R. Gheyi, and T. Massoni. Automated behavioral testing of refactoring engines. IEEE Transactions on Software Engineering, 39(2):147–162, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6175911.

Svensson:2012:QRI

[SGR^{+}12] R. Berntsson Svensson, T. Gorschek, B. Regnell, R. Torkar, A. Shahrokni, and R. Feldt. Quality requirements in industrial practice — an extended interview study at eleven companies. IEEE Transactions on Software Engineering, 38(4):923–935, July/August 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5753901.

Song:2019:CIR

[SGS19] Q. Song, Y. Guo, and M. Shepperd. A comprehensive investigation of the role of imbalanced learning for software defect prediction. IEEE Transactions on Software Engineering, 45(12): 1253–1269, December 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Shatnawi:2010:QIA

[Sha10] R. Shatnawi. A quantitative investigation of the acceptable risk levels of objectoriented metrics in opensource systems. IEEE Transactions on Software Engineering, 36(2):216–225, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
5383377.

Shepperd:2018:ARC

[SHB18] Martin Shepperd, Tracy Hall, and David Bowes. Authors’ reply to “Comments on Researcher Bias: The Use of Machine Learning in Software Defect Prediction”. IEEE Transactions on Software Engineering, 44(11):1129–1131, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7990255/. See

[SBH14, TMHM16].

Shi:2016:VSA

[SHCX16] Q. Shi, J. Huang, Z. Chen, and B. Xu. Verifying synchronization for atomicity violation fixing. IEEE Transactions on Software Engineering, 42(3):280–296, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7254228.

Smith:2019:HDD

[SJMH^{+}19] J. Smith, B. Johnson, E. Murphy-Hill, B. Chu, and H. R. Lipford. How developers diagnose potential security vulnerabilities with a static analysis tool. IEEE Transactions on Software Engineering, 45 (9):877–897, September 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Song:2011:GSD

[SJS^{+}11] Q. Song, Z. Jia, M. Shepperd, S. Ying, and J. Liu. A general software defectproneness prediction framework. IEEE Transactions on Software Engineering, 37 (3):356–370, May/June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5611551.

Salmeron:2012:FRI

[SL12] J. L. Salmeron and C. Lopez. Forecasting risk impact on ERP maintenance with augmented fuzzy cognitive maps. IEEE Transactions on Software Engineering, 38(2): 439–452, March/April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5680917.

Shahbazi:2016:BBS

[SM16] A. Shahbazi and J. Miller. Black-box string test case generation through a multiobjective optimization. IEEE Transactions on Software Engineering, 42(4):361–378, April 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7293669.

Shen:2016:DSP

[SMBY16] X. Shen, L. L. Minku, R. Bahsoon, and X. Yao. Dynamic software project scheduling through a proactiverescheduling method. IEEE Transactions on Software Engineering, 42(7):658–686, July 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7365465.

Salleh:2011:ESP

[SMG11] N. Salleh, E. Mendes, and J. Grundy. Empirical studies of pair programming for CS/ SE teaching in higher education: A systematic literature review. IEEE Transactions on Software Engineering, 37(4):509–525, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5482588.

Song:2019:IMV

[SMJ19] W. Song, X. Ma, and H. Jacobsen. Instance migration validity for dynamic evolution of data-aware processes. IEEE Transactions on Software Engineering, 45(8):782– 801, August 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Shin:2011:ECC

[SMWO11] Y. Shin, A. Meneely, L. Williams, and J. A. Osborne. Evaluating complexity, code churn, and developer activity metrics as indicators of software vulnerabilities. IEEE Transactions on Software Engineering, 37(6):772–787, November/December 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5560680.

Syer:2015:RRE

[SNAH15] M. D. Syer, M. Nagappan, B. Adams, and A. E. Hassan. Replicating and re-evaluating the theory of relative defectproneness. IEEE Transactions on Software Engineering, 41(2):176–197, February 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6914599.

Salvaneschi:2017:PER

[SPA^{+}17] G. Salvaneschi, S. Proksch, S. Amann, S. Nadi, and M. Mezini. On the positive effect of reactive programming on software comprehension: An empirical study. IEEE Transactions on Software Engineering, 43 (12):1125–1143, December 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7827078.

Song:2014:IED

[SPF14] C. Song, A. Porter, and J. S.

Foster. iTree: Efficiently discovering high-coverage configurations using interaction trees. IEEE Transactions on Software Engineering, 40(3):251–265, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6671585.

Simons:2010:IES

[SPG10] Christopher L. Simons, Ian C.

Parmee, and Rhys Gwynllyw. Interactive, evolutionary search in upstream object-oriented class design. IEEE Transactions on Software Engineering, 36(6):798– 816, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5432223.

Sakti:2015:IGP

[SPG15] A. Sakti, G. Pesant, and Y. G. Guéhéneuc. Instance generator and problem representation to improve object oriented code coverage. IEEE Transactions on Software Engineering, 41(3):294–313, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6926828.

Segura:2018:MTR

[SPTRC18] Sergio Segura, José A. Parejo, Javier Troya, and Antonio Ruiz-Cortés. Metamorphic testing of RESTful Web APIs. IEEE Transactions on Software Engineering, 44(11):1083–1099, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8074764/.

Skene:2010:SLA

[SRE10] J. Skene, F. Raimondi, and W. Emmerich. Servicelevel agreements for electronic services. IEEE Transactions on Software Engineering, 36(2):288–304, March/ April 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5210121.

Sen:2012:MDW

[SRS12] A. Sen, K. Ramamurthy, and A. P. Sinha. A model of data warehousing process maturity. IEEE Transactions on Software Engineering, 38(2):336–353, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5680911.

Sarma:2012:PED

[SRvdH12] A. Sarma, D. F. Redmiles, and A. van der Hoek. Palantir: Early detection of development conflicts arising from parallel code changes. IEEE Transactions on Software Engineering, 38(4): 889–908, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5928359.

Schur:2015:MWM

[SRZ15] M. Schur, A. Roth, and A. Zeller. Mining workflow models from Web applications. IEEE Transactions on Software Engineering, 41 (12):1184–1201, December 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7169616.

Singh:2018:EBS

[SSP18] V. B. Singh, Meera Sharma, and Hoang Pham. Entropy based software reliability analysis of multiversion open source software. IEEE Transactions on Software Engineering, 44 (12):1207–1223, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8081836/.

Shepperd:2013:DQS

[SSSM13] M. Shepperd, Q. Song, Z. Sun, and C. Mair. Data quality: Some comments on the NASA software defect datasets. IEEE Transactions on Software Engineering, 39(9):1208–1215, September 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6464273.

Shahbazi:2013:CVT

[STM13] A. Shahbazi, A. F. Tappenden, and J. Miller. Centroidal Voronoi tessellations

— a new approach to random testing. IEEE Transactions on Software Engineering, 39(2):163–183, February 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6165316.

Schafer:2012:CAN

[STST12] M. Schäfer, A. Thies, F. Steimann, and F. Tip. A comprehensive approach to naming and accessibility in refactoring Java programs. IEEE Transactions on Software Engineering, 38(6):1233–1257, November/December 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6152131.

Shivaji:2013:RFI

[SWAK13] S. Shivaji, E. James Whitehead, R. Akella, and S. Kim. Reducing features to improve code change-based bug prediction. IEEE Transactions on Software Engineering, 39(4):552–569, April 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6226427.

Samuel:2015:CRC

[SWEK15] B. M. Samuel, L. A. Watkins III, A. Ehle, and V. Khatri. Customizing the representation capabilities of process models: Understanding the effects of perceived modeling impediments. IEEE Transactions on Software Engineering, 41(1):19–39, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6898868.

Scandariato:2014:PVS

[SWHJ14] R. Scandariato, J. Walden, A. Hovsepyan, and W. Joosen. Predicting vulnerable software components via text mining. IEEE Transactions on Software Engineering, 40(10):993–1006, October 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6860243.

Sjoberg:2013:QEC

[SYA^{+}13] D. I. K. Sjøberg, A. Yamashita, B. C. D. Anda, A. Mockus, and T. Dybå. Quantifying the effect of code smells on maintenance effort. IEEE Transactions on Software Engineering, 39(8):1144–1156, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6392174.

Shin:2018:TES

[SYB18] Donghwan Shin, Shin Yoo, and Doo-Hwan Bae. A theoretical and empirical study of diversity-aware mutation adequacy criterion. IEEE Transactions on Software Engineering, 44(10):914–931, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7994647/.

Sui:2014:DML

[SYX14] Y. Sui, D. Ye, and J. Xue.

Detecting memory leaks statically with full-sparse valueflow analysis. IEEE Transactions on Software Engineering, 40(2):107–122, February 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6720116.

Spanoudakis:2010:DSD

[SZ10] G. Spanoudakis and A. Zisman. Discovering services during service-based system design using UML. IEEE Transactions on Software Engineering, 36(3): 371–389, May/June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5374424.

Storey:2017:HSC

[SZF^{+}17] M. A. Storey, A. Zagalsky, F. F. Filho, L. Singer, and D. M. German. How social and communication channels shape and challenge a participatory culture in software development. IEEE Transactions on Software Engineering, 43(2):185–204, February 2017. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7498605.

Sumner:2012:PCC

[SZWZ12] W. N. Sumner, Y. Zheng, D. Weeratunge, and X. Zhang. Precise calling context encoding. IEEE Transactions on Software Engineering, 38(5):1160–1177, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5963696.

Tian:2014:MCM

[TDD14] C. Tian, Z. Duan, and Z. Duan. Making CE-GAR more efficient in software model checking. IEEE Transactions on Software Engineering, 40(12):1206– 1223, December 2014. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6895263.

Tribastone:2012:FRS

[TDGH12] M. Tribastone, J. Ding, S. Gilmore, and J. Hillston. Fluid rewards for a stochastic process algebra. IEEE Transactions on Software Engineering, 38(4):861–874, July/August 2012. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5975178.

Tuya:2016:CAT

[TdlRSCB16] J. Tuya, C. d. l. Riva, M. J.

Suárez-Cabal, and R. Blanco. Coverage-aware test database reduction. IEEE Transactions on Software Engineering, 42(10):941–959, October 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7384760.

Trummer:2014:MOQ

[TFB14] I. Trummer, B. Faltings, and W. Binder. Multiobjective quality-driven service selection — a fully polynomial time approximation scheme. IEEE Transactions on Software Engineering, 40(2):167–191, February 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6687160.

Tribastone:2012:SDA

[TGH12] M. Tribastone, S. Gilmore, and J. Hillston. Scalable differential analysis of process algebra models. IEEE Transactions on Software Engineering, 38(1): 205–219, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5567115.

Thimbleby:2015:SUI

[Thi15] H. Thimbleby. Safer user interfaces: A case study in improving number entry. IEEE Transactions on Software Engineering, 41(7):711–729, July 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6991548.

Taiani:2014:GUC

[TLB14] F. Taïani, S. Lin, and G. S. Blair. GossipKit: A unified component framework for Gossip. IEEE Transactions on Software Engineering, 40(2):123–136, February 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6645372.

TerBeek:2020:FQM

[TLLV20] M. H. Ter Beek, A. Legay, A. L. Lafuente, and A. Vandin. A framework for quantitative modeling and analysis of highly (re)configurable systems. IEEE Transactions on Software Engineering, 46(3):321–345, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Tian:2018:RSP

[TLZ^{+}18] Zhenzhou Tian, Ting Liu, Qinghua Zheng, Eryue Zhuang, Ming Fan, and Zijiang Yang. Reviving sequential program birthmarking for multithreaded software plagiarism detection. IEEE Transactions on Software Engineering, 44(5):491–511, May 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7888597/.

Tantithamthavorn:2016:CRB

[TMHM16] C. Tantithamthavorn, S. McIntosh, A. E. Hassan, and K. Matsumoto. Comments on “researcher bias: The use of machine learning in software defect prediction”. IEEE Transactions on Software Engineering, 42 (11):1092–1094, November 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

7450669. See [SBH14] and reply [SHB18].

Tantithamthavorn:2017:ECM

[TMHM17] C. Tantithamthavorn, S. McIntosh, A. E. Hassan, and K. Matsumoto. An empirical comparison of model validation techniques for defect prediction models. IEEE Transactions on Software Engineering, 43(1):1–18, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=

7497471. See comments

[Her17].

Tantithamthavorn:2019:IAP

[TMHM19] C. Tantithamthavorn, S. McIntosh, A. E. Hassan, and K. Matsumoto. The impact of automated parameter optimization on defect prediction models. IEEE Transactions on Software Engineering, 45 (7):683–711, July 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Tsantalis:2015:ARS

[TMK15] N. Tsantalis, D. Mazinanian, and G. P. Krishnan. Assessing the refactorability of software clones. IEEE Transactions on Software Engineering, 41 (11):1055–1090, November 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7130676.

Thomas:2013:ICC

[TNBH13] S. W. Thomas, M. Nagappan, D. Blostein, and A. E. Hassan. The impact of classifier configuration and classifier combination on bug localization. IEEE Transactions on Software Engineering, 39(10):1427–1443, October 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6520844.

Toledo:2012:AJA

[TNTN12] R. Toledo, A. Nunez, E. Tanter, and J. Noye. Aspectizing Java access control. IEEE Transactions on Software Engineering, 38(1): 101–117, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5680915.

Tufano:2017:WWY

[TPB^{+}17] M. Tufano, F. Palomba, G. Bavota, R. Oliveto, M. D. Penta, A. De Lucia, and D. Poshyvanyk. When and why your code starts to smell bad (and whether the smells go away). IEEE Transactions on Software Engineering, 43 (11):1063–1088, November 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7817894.

Treude:2015:EDT

[TRD15] C. Treude, M. P. Robillard, and B. Dagenais. Extracting development tasks to navigate software documentation. IEEE Transactions on Software Engineering, 41(6):565–581, June 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7000568.

Tribastone:2013:FML

[Tri13]  M. Tribastone.  A fluid

model for layered queueing networks. IEEE Transactions on Software Engineering, 39(6):744–756, June 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6314480.

Treude:2012:WIT

[TS12] C. Treude and M. A. Storey. Work item tagging: Communicating concerns in collaborative software development. IEEE Transactions on Software Engineering, 38 (1):19–34, January/February 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5611552.

Thome:2020:IAE

[TSBB20] J. Thomé, L. K. Shar, D. Bianculli, and L. Briand.

An integrated approach for effective injection vulnerability analysis of web applications through security slicing and hybrid constraint solving. IEEE Transactions on Software Engineering, 46(2):163– 195, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Torrado:2013:SRM

[TWL13] N. Torrado, M. P. Wiper, and R. E. Lillo. Software reliability modeling with software metrics data via Gaussian processes. IEEE Transactions on Software Engineering, 39(8):1179–1186, August 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6392172.

Tian:2015:SPD

[TZL^{+}15] Z. Tian, Q. Zheng, T. Liu, M. Fan, E. Zhuang, and Z. Yang. Software plagiarism detection with birthmarks based on dynamic key instruction sequences. IEEE Transactions on Software Engineering, 41(12): 1217–1235, December 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7153572.

Tang:2017:SNI

[TZM^{+}17] E. Tang, X. Zhang, N. T. Müller, Z. Chen, and X. Li. Software numerical instability detection and diagnosis by combining stochastic and infinite-precision testing. IEEE Transactions on Software Engineering, 43 (10):975–994, October 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7792694.

Unterkalmsteiner:2012:EMS

[UGI^{+}12] M. Unterkalmsteiner, T. Gorschek, A. K. M. M. Islam, C. K. Cheng, R. B. Permadi, and R. Feldt. Evaluation and measurement of software process improvement

— a systematic literature review. IEEE Transactions on Software Engineering, 38(2):398–424, March/ April 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5728832.

Uzuncaova:2010:ITG

[UKB10] E. Uzuncaova, S. Khurshid, and D. Batory. Incremental test generation for software product lines. IEEE Transactions on Software Engineering, 36(3):309–322, May/ June 2010. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5456077.

Upadhyaya:2018:ASC

[UR18] Ganesha Upadhyaya and Hridesh Rajan. On accelerating source code analysis at massive scale. IEEE Transactions on Software Engineering, 44(7):669–688, July 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8344505/.

Vegas:2016:CDS

[VAJ16] S. Vegas, C. Apa, and N. Juristo. Crossover designs in software engineering experiments: Benefits and perils. IEEE Transactions on Software Engineering, 42(2):120–135, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7192651.

Vara:2014:DTM

[VBJM14] J. M. Vara, V. A. Bollati, Á. Jiménez, and E. Marcos. Dealing with traceability in the MDD of model transformations. IEEE Transactions on Software Engineering, 40(6):555–583, June 2014. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6784505.

Valente:2012:SAA

[VBP12] M. T. Valente, V. Borges, and L. Passos. A semiautomatic approach for extracting software product lines. IEEE Transactions on Software Engineering, 38(4):737–754, July/August 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5928352.

Vilela:2020:SPR

[VCMG20] J. Vilela, J. Castro, L. E. G. Martins, and T. Gorschek. Safety practices in requirements engineering: The Uni-REPM safety module. IEEE Transactions on Software Engineering, 46(3):222–250, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Valero:2017:TAM

[VDC17] V. Valero, G. Díaz, and M. E. Cambronero. Timed automata modeling and verification for publish–subscribe structures using distributed resources. IEEE Transactions on Software Engineering, 43(1):76–99, January 2017. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7463051.

vanderLinden:2019:SLR

[vH19] D. van der Linden and I. Hadar. A systematic literature review of applications of the physics of notations. IEEE Transactions on Software Engineering, 45(8): 736–759, August 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

VanOorschot:2012:RUM

[VW12] P. C. Van Oorschot and G. Wurster. Reducing unauthorized modification of digital objects. IEEE Transactions on Software Engineering, 38(1):191–204, January/February 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5680916.

White:2010:AAT

[WDS10] J. White, B. Doughtery, and D. C. Schmidt. AS-CENT: An algorithmic technique for designing hardware and software in tandem. IEEE Transactions on Software Engineering, 36(6):838– 851, November/December 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5539763.

Wang:2014:PCM

[WDZ^{+}14] X. Wang, Y. Dang, L. Zhang, D. Zhang, E. Lan, and H. Mei. Predicting consistency-maintenance requirement of code clonesat copy-and-paste time. IEEE Transactions on Software Engineering, 40(8):773–794, August 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6815760.

Wnuk:2016:SST

[WGC^{+}16] K. Wnuk, T. Gorschek, D. Callele, E. A. Karlsson, E. Åhlin, and B. Regnell. Supporting scope tracking and visualization for very large-scale requirements engineering-utilizing FSC+, decision patterns, and atomic decision visualizations. IEEE Transactions on Software Engineering, 42(1):47–74, January 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7123669.

Wong:2016:SSF

[WGL^{+}16] W. E. Wong, R. Gao, Y. Li, R. Abreu, and F. Wotawa. A survey on software fault localization. IEEE Transactions on Software Engineering, 42(8):707–740, August 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7390282.

Wang:2018:FCB

[WHYY18] Yanchun Wang, Qiang He, Dayong Ye, and Yun Yang. Formulating criticality-based cost-effective fault tolerance strategies for multi-tenant service-based systems. IEEE Transactions on Software Engineering, 44(3):291–307, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/document/7876832/.

Wang:2017:DGS

[WLG^{+}17] H. Wang, T. Liu, X. Guan, C. Shen, Q. Zheng, and Z. Yang. Dependence guided symbolic execution. IEEE Transactions on Software Engineering, 43(3):252– 271, March 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7497518.

Wu:2016:LWI

[WLX^{+}16] T. Wu, J. Liu, Z. Xu, C. Guo, Y. Zhang, J. Yan, and J. Zhang. Lightweight, inter-procedural and callback-aware resource leak detection for Android apps. IEEE Transactions on Software Engineering, 42(11): 1054–1076, November 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7442579.

Weidlich:2011:ECM

[WMW11] M. Weidlich, J. Mendling, and M. Weske. Efficient consistency measurement based on behavioral profiles of process models. IEEE Transactions on Software Engineering, 37(3):410–429, May/ June 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5611557.

Wilkerson:2012:CDR

[WNM12] J. W. Wilkerson, J. F. Nunamaker, and R. Mercer. Comparing the defect reduction benefits of code inspection and test-driven development. IEEE Transactions on Software Engineering, 38(3):547–560, May/ June 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5750007.

Wu:2020:ECC

[WNP^{+}20] H. Wu, C. Nie, J. Petke, Y. Jia, and M. Harman. An empirical comparison of combinatorial testing, random testing and adaptive random testing. IEEE Transactions on Software Engineering, 46(3):302–320, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Wu:2015:EAD

[WSC^{+}15] L. Wu, K. Su, S. Cai, X. Zhang, C. Zhang, and S. Wang. An I/O efficient approach for detecting all accepting cycles. IEEE Transactions on Software Engineering, 41(8):730–744, August 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7056483.

Wang:2015:SSE

[WSW^{+}15] T. Wang, J. Sun, X. Wang, Y. Liu, Y. Si, J. S. Dong, X. Yang, and X. Li. A systematic study on explicitstate non-Zenoness checking for timed automata. IEEE Transactions on Software Engineering, 41(1):3–18, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6908008.

Wang:2017:LIC

[WSWQ17] X. Wang, J. Sun, T. Wang, and S. Qin. Language inclusion checking of timed automata with non-zenoness. IEEE Transactions on Software Engineering, 43(11): 995–1008, November 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7819478.

Wang:2017:ORP

[WWY^{+}17] H. Wang, L. Wang, Q. Yu, Z. Zheng, A. Bouguettaya, and M. R. Lyu. Online reliability prediction via motifsbased dynamic Bayesian networks for service-oriented systems. IEEE Transactions on Software Engineering, 43(6):556–579, June 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7585067.

Wang:2018:ULC

[WWY^{+}18] Zhe Wang, Chenggang Wu, Xiang Yuan, Zhenjiang Wang, Jianjun Li, Pen-Chung Yew, Jeff Huang, Xiaobing Feng, Yanyan Lan, Yunji Chen, Yuanming Lai, and Yong Guan. Using local clocks to reproduce concurrency bugs. IEEE Transactions on Software Engineering, 44 (11):1112–1128, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/8038023/.

Wang:2018:ASR

[WYZ^{+}18] Ying Wang, Hai Yu, Zhiliang Zhu, Wei Zhang, and Yuli Zhao. Automatic software refactoring via weighted clustering in method-level networks. IEEE Transactions on Software Engineering, 44(3):202–236, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/document/7874207/.

Ward:2011:DSA

[WZ11] M. P. Ward and H. Zedan.

Deriving a slicing algorithm via Fermat transformations. IEEE Transactions on Software Engineering, 37(1):24– 47, January/February 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5401170.

Wang:2013:LNE

[WZX^{+}13] X. Wang, L. Zhang, T. Xie, H. Mei, and J. Sun. Locating need-to-externalize constant strings for software internationalization with generalized string-taint analysis. IEEE Transactions on Software Engineering, 39(4):516–536, April 2013. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6216383.

Xia:2018:MPC

[XBL^{+}18] Xin Xia, Lingfeng Bao, David Lo, Zhenchang Xing, Ahmed E. Hassan, and Shanping Li. Measuring program comprehension: A largescale field study with professionals. IEEE Transactions on Software Engineering, 44(10):951–976, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7997917/.

Xie:2019:ALS

[XCZ^{+}19] X. Xie, B. Chen, L. Zou, Y. Liu, W. Le, and X. Li. Automatic loop summarization via path dependency analysis. IEEE Transactions on Software Engineering, 45(6): 537–557, June 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Xuan:2012:SLS

[XJRL12] J. Xuan, H. Jiang, Z. Ren, and Z. Luo. Solving the large scale next release problem with a backbone-based multilevel algorithm. IEEE Transactions on Software Engineering, 38(5):1195–1212, September/October 2012.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6261327.

Xia:2017:IAB

[XLD^{+}17] X. Xia, D. Lo, Y. Ding, J. M.

Al-Kofahi, T. N. Nguyen, and X. Wang. Improving automated bug triaging with specialized topic model. IEEE Transactions on Software Engineering, 43(3):272–297, March 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7484672.

Xia:2016:HMC

[XLP^{+}16] X. Xia, D. Lo, S. J.

Pan, N. Nagappan, and X. Wang. HYDRA: Massively compositional model for cross-project defect prediction. IEEE Transactions on Software Engineering, 42(10):977–998, October 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7435328.

Xuan:2017:NAR

[XMD^{+}17] J. Xuan, M. Martinez, F. De-Marco, M. Clément, S. L. Marcote, T. Durieux, D. Le Berre, and M. Monperrus.

Nopol: Automatic repair of conditional statement bugs in Java programs. IEEE Transactions on Software Engineering, 43(1):34–55, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7463060.

Xu:2015:CSD

[XXC^{+}15] C. Xu, W. Xi, S. C. Cheung, X. Ma, C. Cao, and J. Lu. Cina: Suppressing the detection of unstable context inconsistency. IEEE Transactions on Software Engineering, 41(9):842–865, September 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7078871.

Xue:2019:ASC

[XXCL19] Y. Xue, Z. Xu, M. Chandramohan, and Y. Liu. Accurate and scalable crossarchitecture cross-OS binary code search with emulation. IEEE Transactions on Software Engineering, 45(11): 1125–1149, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). See author affiliation correction [XXCL21].

Xue:2021:EAS

[XXCL21] Yinxing Xue, Zhengzi Xu, Mahinthan Chandramohan, and Yang Liu. Erratum to Accurate and Scalable Cross-Architecture Cross-OS Binary Code Search With Emulation. IEEE Transactions on Software Engineering, 47(5):1088, 2021. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). See [XXCL19].

Xiong:2015:RFI

[XZH^{+}15] Y. Xiong, H. Zhang, A. Hubaux, S. She, J. Wang, and K. Czarnecki. Range fixes: Interactive error resolution for software configuration. IEEE Transactions on Software Engineering, 41(6):603–619, June 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6991616.

Ye:2016:MBR

[YBL16] X. Ye, R. Bunescu, and C. Liu. Mapping bug reports to relevant files: A ranking model, a fine-grained benchmark, and feature evaluation. IEEE Transactions on Software Engineering, 42(4):379–402, April 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7270328.

Yuan:2011:GIT

[YCM11] X. Yuan, M. B. Cohen, and A. M. Memon. GUI interaction testing: Incorporating event context. IEEE Transactions on Software Engineering, 37(4): 559–574, July/August 2011. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5444885.

Yilmaz:2014:RME

[YDCP14] C. Yilmaz, E. Dumlu, M. B. Cohen, and A. Porter. Reducing masking effects in combinatorial interaction testing: A feedback driven adaptive approach. IEEE Transactions on Software Engineering, 40(1):43–66, January 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6654147.

Ying:2011:FLQ

[YF11] M. Ying and Y. Feng. A flowchart language for quantum programming. IEEE Transactions on Software Engineering, 37(4):466–485, July/August 2011. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5611555.

Yusop:2017:RUD

[YGV17] N. S. M. Yusop, J. Grundy, and R. Vasa. Reporting usability defects: A systematic literature review. IEEE Transactions on Software Engineering, 43(9): 848–867, September 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7779159.

Yu:2020:CDT

[YHW20] T. Yu, Z. Huang, and C. Wang. ConTesa: Directed test suite augmentation for concurrent software. IEEE Transactions on Software Engineering, 46(4):405– 419, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yilmaz:2013:TCA

[Yil13] Cemal Yilmaz. Test caseaware combinatorial interaction testing. IEEE Transactions on Software Engineering, 39(5):684–706, May 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6311411.

Ye:2013:WST

[YJ13] C. Ye and H. A. Jacobsen.

Whitening SOA testing via event exposure. IEEE Transactions on Software Engineering, 39(10):1444–1465, October 2013. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6495456.

Yu:2018:EDB

[YLQ^{+}18] Le Yu, Xiapu Luo, Chenxiong Qian, Shuai Wang, and Hareton K. N. Leung. Enhancing the description-to-behavior fidelity in Android apps with privacy policy. IEEE Transactions on Software Engineering, 44(9):834–854, September 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/7987793/.

Yuan:2010:GES

[YM10] X. Yuan and A. M. Memon. Generating event sequencebased test cases using GUI runtime state feedback. IEEE Transactions on Software Engineering, 36(1):81–95, January/February 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5306073.

Yu:2018:TCG

[YT18] L. Yu and W. T. Tsai. Test case generation for Boolean expressions by cell covering. IEEE Transactions on Software Engineering, 44(1):70–99, January 2018. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7855791.

Yu:2019:CCD

[YWHH19] T. Yu, W. Wen, X. Han, and J. H. Hayes. ConPredictor: Concurrency defect prediction in real-world applications. IEEE Transactions on Software Engineering, 45(6): 558–575, June 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yan:2019:ACL

[YXS^{+}19] M. Yan, X. Xia, E. Shihab, D. Lo, J. Yin, and X. Yang. Automating changelevel self-admitted technical debt determination. IEEE Transactions on Software Engineering, 45(12):1211– 1229, December 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yi:2018:EPR

[YYG^{+}18] Q. Yi, Z. Yang, S. Guo, C. Wang, J. Liu, and C. Zhao. Eliminating path redundancy via postconditioned symbolic execution. IEEE Transactions on Software Engineering, 44(1):25–43, January 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?arnumber=
                                        7835264.

Yang:2015:SBC

[YZL^{+}15] Y. Yang, Y. Zhou, H. Lu, L. Chen, Z. Chen, B. Xu, H. Leung, and Z. Zhang. Are slice-based cohesion metrics actually useful in effort-aware post-release fault-proneness prediction? An empirical study. IEEE Transactions on Software Engineering, 41(4):331–357, April 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6954519.

Zave:2017:RAI

[Zav17] P. Zave. Reasoning about identifier spaces: How to make chord correct. IEEE Transactions on Software Engineering, 43(12):1144– 1156, December 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7823003.

Zhang:2012:WDW

[ZB12a] C. Zhang and D. Budgen. What do we know about the effectiveness of software design patterns? IEEE Transactions on Software Engineering, 38(5):1213–1231, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5975176.

Zickert:2012:CES

[ZB12b] F. Zickert and R. Beck.

Coping with existing systems in information systems development. IEEE Transactions on Software Engineering, 38(5):1027–1039, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5999674.

Zhu:2015:CDP

[ZB15] H. Zhu and I. Bayley. On the composability of design patterns. IEEE Transactions on Software Engineering, 41 (11):1138–1152, November 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7123660.

Zhang:2015:RRV

[ZCW15] L. Zhang, A. Chattopadhyay, and C. Wang. Round-Up: Runtime verification of quasi linearizability for concurrent data structures. IEEE Transactions on Software Engineering, 41(12): 1202–1216, December 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7192659.

Zhang:2017:USA

[ZHMZ17] F. Zhang, A. E. Hassan, S. McIntosh, and Y. Zou. The use of summation to aggregate software metrics hinders the performance of defect prediction models. IEEE Transactions on Software Engineering, 43(5):476–491, May 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        7539677.

Zhang:2012:MCC

[ZJ12] C. Zhang and H. A. Jacobsen. Mining crosscutting concerns through random walks. IEEE Transactions on Software Engineering, 38(5):1123–1137, September/October 2012. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5989837.

Zanjani:2016:ARP

[ZKB16] M. B. Zanjani, H. Kagdi, and C. Bird. Automatically recommending peer reviewers in modern code review. IEEE Transactions on Software Engineering, 42(6):530–543, June 2016. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7328331.

Zhao:2010:PBD

[ZKZ10] C. Zhao, J. Kong, and K. Zhang. Program behavior discovery and verification: A graph grammar approach. IEEE Transactions on Software Engineering, 36(3):431–448, May/ June 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5383371.

Zhou:2015:WWS

[ZM15] M. Zhou and A. Mockus. Who will stay in the FLOSS community? Modeling participant’s initial behavior. IEEE Transactions on Software Engineering, 41(1):82–99, January 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6880395.

Zhong:2019:ESA

[ZM19] H. Zhong and H. Mei. An empirical study on API usages. IEEE Transactions on Software Engineering, 45(4): 319–334, April 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zhang:2018:EQC

[ZNKZ18] Feng Zhang, Haoran Niu, Iman Keivanloo, and Ying Zou. Expanding queries for code search using semantically related API classnames. IEEE Transactions on Software Engineering, 44(11):1070–1082, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8031055/.

Zimmermann:2010:WMG

[ZPB^{+}10] T. Zimmermann, R. Premraj, N. Bettenburg, S. Just, A. Schroter, and C. Weiss. What makes a good bug report? IEEE Transactions on Software Engineering, 36(5): 618–643, September/October 2010. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5487527.

Zisman:2013:PRR

[ZSDS13] A. Zisman, G. Spanoudakis, J. Dooley, and I. Siveroni. Proactive and reactive runtime service discovery: A framework and its evaluation. IEEE Transactions on Software Engineering, 39(7):954–974, July 2013. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
6375710.

Zhang:2016:LSD

[ZWL^{+}16] M. Zhang, Y. Wu, S. Lu, S. Qi, J. Ren, and W. Zheng. A lightweight system for detecting and tolerating concurrency bugs. IEEE Transactions on Software Engineering, 42(10):899–917, October 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7412768.

Zhou:2016:MTS

[ZXC16] Z. Q. Zhou, S. Xiang, and T. Y. Chen. Metamorphic testing for software quality assessment: A study of search engines. IEEE Transactions on Software Engineering, 42(3):264–284, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7254235.

Zhang:2019:PMT

[ZZH^{+}19] J. Zhang, L. Zhang, M. Harman, D. Hao, Y. Jia, and L. Zhang. Predictive mutation testing. IEEE Transactions on Software Engineering, 45(9):898–918, September 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).